### v2

Abstract
With a view on bilevel and PDE-constrained optimisation, we develop iterative estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ for composite functions $F:=J\circ S$, where $S$ is the solution mapping of the inner optimisation problem or PDE. The idea is to form a single-loop method by interweaving updates of the iterate $x^{k}$ by an outer optimisation method, with updates of the estimate by single steps of standard optimisation methods and linear system solvers. When the inner methods satisfy simple tracking inequalities, the differential estimates can almost directly be employed in standard convergence proofs for general forward-backward type methods. We adapt those proofs to a general inexact setting in normed spaces, that, besides our differential estimates, also covers mismatched adjoints and unreachable optimality conditions in measure spaces. As a side product of these efforts, we provide improved convergence results for nonconvex Primal-Dual Proximal Splitting (PDPS).
1 Introduction
First-order methods are slow. To be precise, they require a high number of iterations, but if those iterations are fast, they have the chance to practically overpower second-order methods with expensive iterations. In PDE-constrained or bilevel optimisation, the steps of basic first-order methods are very expensive, involving the solution of the inner problem or PDE and its adjoint. To make first-order methods fast, it is, therefore, imperative to reduce the cost of solving these subproblems—for instance, by employing inexact solution schemes.
Consequently, especially in the machine learning community, an interest has surfaced in single-loop methods for bilevel optimisation; see [28] and references therein. Many of these methods are very specific constructions. In [16] we started work on a more general approach to PDE-constrained optimisation: we showed that on each step of an outer primal-dual optimisation method, we can take single steps of standard linear system splitting schemes for the PDE constraint and its adjoint, and still obtain a convergent method that is computationally significantly faster than solving the PDEs exactly. In [29] we then presented an approach to bilevel optimisation that allowed general inner and adjoint algorithms that satisfy certain tracking inequalities. These were proved for standard splitting schemes for the adjoint equation, and for forward-backward splitting and the Primal-Dual Proximal Splitting (PDPS) of [4] for the inner problem. The overall analysis was still tied to bilevel optimisation in Hilbert spaces, with forward-backward splitting as the outer optimisation method.
Writing $F=J\circ S_{u}$ for a solution mapping $S_{u}:X\to U$ of an inner problem (PDE or an optimisation problem, or multiple nested levels of such), and a differentiable function $J:U\to\mathbb{R}$, on normed spaces $X$ and $U$, we are interested in the solution of composite optimisation problems of the form
(1) $\min_{x\in X}\leavevmode\nobreak\ F(x)+G(x),$
or, more generally, the solution of optimality conditions
(2) $0\in F^{\prime}(x)+\partial G(x)+\Xi x,$
for $G$ convex but possibly nonsmooth, and $\Xi\in\mathbb{L}(X;X^{*})$ skew-adjoint. If $\Xi=0$, then this optimality condition is typically necessary for (1). More generally, the operator allows the modelling of primal-dual problems, and treating the PDPS and Douglas–Rachford splitting as generalised forward-backward splitting methods [7].
Our contributions are as follows. In Section 2, which forms our inner theory,
  1. (a)
    we show in general normed spaces that we can approximate in a single-loop fashion the differentials of compositions $F=J\circ S_{u}$, given abstract inner and adjoint algorithms for $S_{u}$, satisfying certain tracking inequalities.
In contrast to [29] and, indeed, all single-loop bilevel optimisation methods that we are aware of, our approach can also work with the adjoint dimension reduction trick typically employed in PDE-constrained optimisation. We show that, subject to additive error terms with a bounded sum, the differential estimates $\widetilde{F^{\prime}}(x^{k})$ satisfy standard smoothness properties, such as Lipschitz differential and the two- and three-point descent inequalities [31, 7]. Based on this, in Section 4, which forms our outer theory,
  1. 2.
    we then prove various forms of convergence of general inexact splitting methods for (2).
In Section 4.1 we also elaborate on the aforementioned connection of primal-dual methods to Eq. 2. To facilitate the analysis of such methods, we first introduce in Section 3 operator-relative variants of the descent inequality.
Through our approach to inexactness, besides gradient estimates for multilevel problems, we can model mismatched adjoints [17], and difficult-to-solve-exactly optimality conditions in measure spaces [34]. We also adopt the approach of [34] to optimisation in normed spaces: instead of Bregman divergences, we construct an inner product structure with a self-adjoint $M\in\mathbb{L}(X;X^{*})$. Our work is related to the study of gradient oracles for smooth convex optimisation in [8], and for nonconvex composite optimisation in [10, 20], both in finite-dimensional Euclidean spaces. Based on sufficient descent and the Kurdyka–Łojasiewicz property, [22] also study inexact methods in $\mathbb{R}^{n}$. Moreover, [3] introduce approaches to control model inexactness in proximal trust region methods, and [27] in non-single-loop gradient methods for bilevel optimisation.
Not content to merely adapt existing proofs to inexact steps and normed spaces, we also present some improvements, especially for the nonconvex PDPS of [30]. We do, however, treat a slightly simplified problem. The original PDPS of [4] applies to $\min_{x}g(x)+h(K(x))$ with $K$ linear and $g$ and $h$ convex. The extension of [30] allows $K$ to be nonlinear. It is further analysed in [5, 6, 18, 11], with a simplified review of the different variants in Banach space in [32]. An alternative extension in [19] allows $g$ and $h$ to be semiconvex. We, instead, consider the special case $\min_{x}f(x)+g(x)+h(Kx)$ with $K$ linear, but $f$ nonconvex, and
  1. 3.
    we show that, for the nonconvex PDPS, the values of the convex envelope of the objective function at ergodic iterates locally converge to a minimum.
We do not include numerical results, as that has already been done in [16, 29]. Through our work, the specific algorithms presented therein can be understood through a clean and generic differential estimation approach.
Notation and basic concepts
We write $\mathbb{L}(X;Y)$ for the space of bounded linear operators between the normed spaces $X$ and $Y$, and $\operatorname{Id}$ for the identity operator. $X^{*}$ stands for the dual space of $X$. When $X$ is Hilbert, we identify $X^{*}$ with $X$. We write $\langle x,y\rangle$ for an inner product, $\langle x^{*}|x\rangle_{X^{*},X}$ for a dual product. We call $M\in\mathbb{L}(X;X^{*})$ self-adjoint if the restriction $M^{*}|X=M$, and positive semi-definite if $\langle x|Mx\rangle_{X,X^{*}}\geq 0$ for all $x\in X$. If both hold, we set $\|x\|_{M}:=\sqrt{\langle Mx|x\rangle}$, and write $\mathbb{O}_{M}(x,r)$ for the radius-$r$ open ball at $x$ in the $M$-(semi-)norm. We also write $M\geq N$ if $M-N$ is positive semi-definite. We extensively use the vectorial Young’s inequality
$\langle x^{*}|x\rangle_{X^{*},X}\leq\frac{a}{2}\|x\|_{X}^{2}+\frac{1}{2a}\|x^{*}\|_{X^{*}}^{2}\qquad\text{for all }x\in X,\,x^{*}\in X^{*}\,a>0.$
For $F:X\to\mathbb{R}$, we write $DF(x)$ for the Gâteaux and $F^{\prime}(x)\in X^{*}$ for the Fréchet derivative at $x$, if they exist. If $X$ is Hilbert, $\nabla F(x)\in X$ stands for the Riesz representation of $F^{\prime}(x)$, i.e., the gradient. For partial derivatives, we use the notation $F^{(x)}(u,x)$. We also write $\operatorname{sub}_{c}F:=\{x\in X\mid F(x)\leq c\}$ for the $c$-sublevel set. With $\overline{\mathbb{R}}:=[-\infty,\infty]$, for a convex $G:X\to\overline{\mathbb{R}}$, we write $\operatorname{dom}G$ for the effective domain, $\partial G(x)$ for the subdifferential at $x$, and $G^{*}:X^{*}\to\overline{\mathbb{R}}$ for the Fenchel conjugate. When $X$ is a Hilbert space, we write $\operatorname{prox}_{F}$ for the proximal map and, with a slight abuse of notation, identify $\partial G(x)$ with the set of Riesz representations of its elements.
2 Tracking estimate recursion
Let $J:U\to\mathbb{R}$ and $S_{u}:X\to U$ be Fréchet differentiable on normed spaces $X$ and $U$. We consider the functional
$F(x)=J(S_{u}(x)).$
As $S_{u}$ and its differential can be expensive to compute, given an iterate $x^{k}$ of an arbitrary outer algorithm for minimising an objective that involves $F$, such as (1), we estimate $S_{u}(x^{k})$ by $u^{k+1}\in U$, and $S_{u}^{\prime}(x^{k})$ by $p^{k+1}\in\mathbb{L}(X;U)$, that is, we estimate
$F^{\prime}(x^{k})=J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\quad\text{by}\quad\widetilde{F^{\prime}}(x^{k})=J^{\prime}(u^{k+1})p^{k+1}.$
When $X$ is Hilbert, we write $\widetilde{\nabla F}(x^{k})$ for the Riesz representation of $\widetilde{F^{\prime}}(x^{k})$. We do not provide a single explicit formula for $u^{k+1}$ and $p^{k+1}$, instead, we assume them to satisfy tracking estimates as in [16, 29]. We formulate these tracking estimates—that are essentially contractivity estimates with suitable penalties for parameter change—in Section 2.1, and provide examples of inner and adjoint methods that satisfy them. We follow with an estimate on the error of approximation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ in Section 2.2. Our goal is to derive, in Section 2.3, variants of standard descent inequalities and Lipschitz bounds for the estimate $\widetilde{F^{\prime}}(x^{k})$.
Although $\widetilde{F^{\prime}}(x^{k})$ will have the above structure, we want to avoid constructing $p^{k+1}\approx S_{u}^{\prime}(x^{k})\in\mathbb{L}(X;U)$ directly due to its high dimensionality. Instead, we seek to only construct the necessary projections through a lower-dimensional variable $w^{k+1}$. We illustrate this idea in the following example.
Example 2.1 (Adjoint equations).
Suppose $S_{u}(x)$ arises from the satisfaction of
(3) $0=T(S_{u}(x),x)\quad\text{for a}\quad T:U\times X\to W_{*}\quad\text{with}\quad W_{*}\text{ a normed space},$
modelling, e.g., a PDE or the first-order optimality conditions of an inner optimisation problem, both parametrised by $x$. By implicit differentiation, subject to sufficient differentiability and (3) holding in a neighbourhood of $x$, we obtain the basic adjoint
(4) $T^{(u)}(S_{u}(x),x)S_{u}^{\prime}(x)+T^{(x)}(S_{u}(x),x)=0\in\mathbb{L}(X;W),$
where $S_{u}^{\prime}(x)\in\mathbb{L}(X;U)$, $T^{(u)}(S_{u}(x),x)\in\mathbb{L}(U;W)$, and $T^{(x)}(S_{u}(x),x)\in\mathbb{L}(X;W)$. Hence, following the derivation of adjoint PDEs in, e.g., [14, §1.6.2] or [24, §1.2], assuming $T^{(u)}(S_{u}(x),x)$ to be invertible, we solve from (4) that
$[J\circ S_{u}]^{\prime}(x)=J^{\prime}(S_{u}(x))S_{u}^{\prime}(x)=w_{x}T^{(x)}(S_{u}(x),x),$
for a $w_{x}=S_{w}(x)\in W$ satisfying the reduced adjoint
(5) $w_{x}T^{(u)}(S_{u}(x),x)+J^{\prime}(S_{u}(x))=0.$
For $x=x^{k}$, we will in practise take $w^{k+1}$ as an operator splitting approximation to
(6) $w^{k+1}T^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0,$
and then set
$\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})\approx J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k}).$
2.1 Basic constructions and assumptions
To track the inexact computations of inner and adjoint variables across iterations, we introduce abstract but verifiable conditions that quantify how closely the computed values follow the outputs of the exact inner and adjoint solution mappings evaluated at the current outer iterate. These tracking assumptions ensure that the accumulated errors remain controlled and that the approximate gradient remains meaningful for descent. The following assumption formalises this idea. There, and in the rest of this section, unless otherwise explicitly mentioned, the distances $b_{X}$, $d_{X^{*}}$, $d_{U}$, and $d_{W}$ are entirely formal expressions. Typically they will be given by norms, but the theory does not require this. Likewise, the “target” $F^{\prime}(x^{k})$ will typically be the Fréchet derivative of $J\circ S_{u}$ at $x^{k}$, but this need not be the case; it is likewise an entirely formal expression. We write squared distances as $b_{X}^{2}(x,y):=b_{X}(x,y)^{2}$.
Assumption 2.1.
For spaces $X$, $X^{*}$, $U$, and $W$, a subset $\Omega\subset X$, an inner solution map $S_{u}:X\to U$ and an adjoint solution map $S_{w}:X\to W$, the following hold:
  1. (i)
    We are given an inner algorithm that, on each iteration $k\geq 1$, given $\{x^{n}\}_{n=0}^{k}\subset\Omega$ and $u^{k}\in U$, produces $u^{k+1}\in U$ satisfying
    $\kappa_{u}d_{U}(u^{k+1},S_{u}(x^{k}))\leq d_{U}(u^{k},S_{u}(x^{k-1}))+\pi_{u}b_{X}(x^{k},x^{k-1})$
    for some $\pi_{u}>0$, $\kappa_{u}>1$, and distance expressions $d_{U}$ and $b_{X}$.
  2. (ii)
    We are given an adjoint algorithm that, on each iteration $k\geq 1$, given $\{(x^{n},u^{n+1})\}_{n=0}^{k}\subset\Omega\times U$, and $w^{k}\in W$, produces $w^{k+1}\in W$ satisfying
    $\kappa_{w}d_{W}(w^{k+1},S_{w}(x^{k}))\leq d_{W}(w^{k},S_{w}(x^{k-1}))+{\mu_{u}}d_{U}(u^{k+1},S_{u}(x^{k}))+\pi_{w}b_{X}(x^{k},x^{k-1})$
    for some ${\mu_{u}},\pi_{w}>0$, $\kappa_{w}>1$, and a distance expression $d_{W}$.
  3. (iii)
    We are given a differential transformation that, on each iteration $k\in\mathbb{N}$, given $u^{k+1}\in U$ and $w^{k+1}\in W$, produces $\widetilde{F^{\prime}}(x^{k})\in X^{*}$ that, for a target $F^{\prime}(x^{k})\in X^{*}$, satisfies for some $\alpha_{u},\alpha_{w}\geq 0$ and a distance expression $d_{X^{*}}$ the bound
    $d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq\alpha_{u}d_{U}(u^{k+1},S_{u}(x^{k}))+\alpha_{w}d_{W}(w^{k+1},S_{w}(x^{k})).$
The inner and adjoint tracking conditions Items (i) and (ii) are parameter change aware contractivity conditions for the inner and adjoint algorithms: if $x^{k}=x^{k-1}$, the former reduces to a standard contractivity condition. The condition Item (iii) allows converting the construction error of $\widetilde{F^{\prime}}(x^{k})$ to the tracking errors of the inner and adjoint algorithms.
We next provide brief examples of inner and adjoint methods that satisfy the corresponding parts of Section 2.1. For the inner and adjoint tracking conditions Items (i) and (ii), detailed proofs are available in [29] for most of the examples presented below. To aid the reader, we explicitly include the proof for the inner forward-backward splitting method as an example.
Example 2.2 (Inner algorithm: forward-backward).
On a Hilbert space $U$ and a normed space $X$, consider the parametric inner problem
$S_{u}(x)=\operatorname*{arg\,min}_{u}f(u;x)+g(u;x)$
for $f$ and $g$ convex in $u$, and differentiable in $(u,x)$; $\nabla f(\,\boldsymbol{\cdot}\,;x)$ $L$-Lipschitz, and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex, both uniformly in $x$. This is an instance of (3) with
$T(u,x)=\nabla f(u;x)+\nabla g(u;x).$
If $S_{u}$ is Lipschitz and $\tau L\leq 1$ for a step length parameter $\tau>0$, then the forward-backward splitting updates
$u^{k+1}:=\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\tau\nabla f(u^{k};x^{k}))$
satisfy Section 2.1Item (i) [29, Theorem 3.4]. Indeed, since $\tau g(\,\boldsymbol{\cdot}\,;x^{k})$ is $\tau\gamma$-strongly convex, its proximal map is $1/(1+\tau\gamma)$-contractive [33, Lemma 4.4], satisfying for all $u,\tilde{u}\in U$ that
$(1+\tau\gamma)\|\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u)-\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(\tilde{u})\|\leq\|u-\tilde{u}\|.$
Likewise the gradient descent map $V(u):=u-\tau\nabla f(u;x^{k})$ is non-expansive when $\tau L\leq 1$, i.e., $\|V(u)-V(\tilde{u})\|\leq\|u-\tilde{u}\|.$ Since $\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(S_{u}(x^{k}))=S_{u}(x^{k})$ [7, Theorem 4.2 & Corollary 6.22], combining these inequalities yields
$(1+\tau\gamma)\|u^{k+1}-S_{u}(x^{k})\|\leq\|V(u^{k})-V(S_{u}(x^{k}))\|\leq\|u^{k}-S_{u}(x^{k})\|.$
Thus, by adding and substracting $S_{u}(x^{k-1})$, applying triangle inequality, and using the assumption that $S_{u}$ is Lipschitz, we conclude that Section 2.1Item (i) is satisfied with $\kappa_{u}=1+\tau\gamma$ and $\pi_{u}=L_{s}$ where $L_{s}$ denotes the Lipschitz constant of $S_{u}$.
Remark 2.3 (Lipschitz solution mapping).
The Lipschitz assumption on $S_{u}$ is guaranteed in sufficiently smooth cases by the classical implicit function theorem applied to the equation $T(u,x)=0$; see [29, Appendix B]. Nonsmooth implicit function theorems and the Aubin or pseudo-Lipschitz property of the set-valued mapping $S_{u}$ are studied in, e.g., [9, 15] as well as [7, Theorem 28.3]. For the the specific case $f(u;x)=\bar{f}(u)$ and $g(u;x)=x\bar{g}(u)$ with a scalar $x$, we refer to [7, Theorem 28.5]. If $S_{u}$ has the Aubin property, it will be Lipschitz if we assume, e.g., strict convexity to ensure the uniqueness of solutions.
Example 2.4 (Inner algorithm: primal-dual).
On a Hilbert space $Z$ and a normed space $X$, consider the inner problem
$\min_{z}f(z;x)+g^{*}(Kz;x).$
for $K\in\mathbb{L}(Z;Y^{*})$ linear and bounded to a Hilbert space $Y^{*}$, both $f$ and $g$ convex in the first parameter, differentiable in both parameters; and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex uniformly in $x$. As an instance of (3), represent the Fenchel–Rockafellar primal-dual optimality conditions of this problem as the root $u$ of the mapping
$T(u,x)=(\nabla f(z;x)+K^{*}y,\nabla g(y;x)-Kz)\quad\text{where}\quad u=(z,y)\in U=Z\times Y.$
If $S_{u}(x)=T^{-1}(\,\boldsymbol{\cdot}\,;x)(0)$ is Lipschitz in $\Omega$, then the PDPS updates [4]
$z^{k+1}=\operatorname{prox}_{\tau f(\,\boldsymbol{\cdot}\,;x^{k})}(z^{k}-\tau K^{*}y^{k})\quad\text{and}\quad y^{k+1}=\operatorname{prox}_{\sigma g(\,\boldsymbol{\cdot}\,;x^{k})}(y^{k}+\sigma K(2z^{k+1}-z^{k}))$
satisfy Section 2.1Item (i) when $\tau\sigma\|K\|\leq 1$. The proof in [29, Theorem 3.6] is fundamentally similar to the forward-backward in Example 2.2, but requires working with operator-induced norms and monotone operators.
The next example covers discretised PDE-constrained optimisation.
Example 2.5 (Inner algorithm: linear system splitting).
For $U$ a Euclidean space and $X$ a normed space, let both $A_{x}\in\mathbb{L}(U;U)$ and $b_{x}\in U$ be Lipschitz in $x\in X$. Consider the inner constraint of $u=S_{u}(x)$ satisfying
(7) $A_{x}u=b_{x}.$
This is again an instance of (3) when we set
$T(u,x)=A_{x}u-b_{x}.$
If $S_{u}(x)=A^{-1}_{x}b_{x}$ is Lipschitz in $\Omega$, then splitting $A_{x}=N_{x}+M_{x}$ per the Jacobi or Gauss–Seidel schemes, the updates $u^{k+1}=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})$ satisfy Section 2.1Item (i) subject to uniform versions of standard convergence conditions for these schemes [29, Examples 4.3 and 4.4]. These proofs in [29] are based on [16, Section 4.2], where squared versions of both Section 2.1Items (i) and 2.6 are proved.
The following two examples treat adjoint methods and the differential transformation for all of Examples 2.2, 2.4 and 2.5. As we have seen, $S_{u}$ for all the former examples arises from (3) for a corresponding $T$, so the next two examples can be combined with each of those examples for the satisfaction of Section 2.1Items (ii) and (iii).
Example 2.6 (Basic adjoint: linear system splitting).
With $U$ and $W_{*}$ Euclidean spaces, and $X$ a normed space, let $S_{u}$ and $T$ be given by (3), and define
$\widetilde{F^{\prime}}(x^{k}):=J^{\prime}(u^{k+1})p^{k+1},$
for $p^{k+1}\in\mathbb{L}(X;U)$ computed by taking (single or multiple) Jacobi or Gauss–Seidel steps on the linear equation
(8) $0=T^{(u)}(u^{k+1},x^{k})p^{k+1}+T^{(x)}(u^{k+1},x^{k}).$
Correspondingly, let $S_{w}=S_{p}=S_{u}^{\prime}$ arise from the basic adjoint (4).
If $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable, then, subject to uniformised standard convergence conditions for each of the above-mentioned linear system splitting schemes, Section 2.1Item (ii) holds with $w^{k}=p^{k}$. The proof in [29, Theorem 4.9] is analogous to that of Example 2.5, indeed, (8) has the form (7) with the parameter $x=(u^{k+1},x^{k})$.
Suppose further that
$N_{J^{\prime}}:=\sup\{\|J^{\prime}(S_{u}(x))\|_{U^{*}}\mid x\in\Omega\}<\infty,\quad N_{S_{u}^{\prime}}:=\sup\{\|S_{u}^{\prime}(x)\|_{\mathbb{L}(X;U)}\mid x\in\Omega\}<\infty,$
and that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz. Then Section 2.1Item (iii) holds due to
$\|\widetilde{F^{\prime}}(x^{k})-$ $F^{\prime}(x^{k})\|_{X^{*}}=\|J^{\prime}(u^{k+1})p^{k+1}-J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\|_{X^{*}}$
$\leq\|J^{\prime}(u^{k+1})[p^{k+1}-S_{u}^{\prime}(x^{k})]\|_{X^{*}}+\|[J^{\prime}(u^{k+1})-J^{\prime}(S_{u}(x^{k}))]S_{u}^{\prime}(x^{k})\|_{X^{*}}$
$\leq\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|J^{\prime}(u^{k+1})\|_{U^{*}}+L_{\nabla J}\|S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|u^{k+1}-S_{u}(x^{k})\|_{U}$
$\leq N_{\nabla J}\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}+L_{\nabla J}N_{\nabla S_{u}}\|u^{k+1}-S_{u}(x^{k})\|_{U}.$
Example 2.7 (Reduced adjoint: linear system splitting).
With $U$ and $W_{*}$ Euclidean spaces, and $X$ a normed space, let $S_{u}$ and $T$ be given by (3), and define
$\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})$
for $w^{k+1}\in W$ computed by taking (single or multiple) Jacobi or Gauss–Seidel steps on the linear equation
$w^{k+1}T^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0.$
Correspondingly, let $S_{w}$ arise from the reduced adjoint (5).
If $T|U\times\Omega$, $S_{u}|\Omega$, and $J$ are Lipschitz-continuously differentiable, then Section 2.1Item (ii) holds subject to uniformised standard convergence conditions for each of the above-mentioned linear system splitting schemes. Although not proved in [29], this claim follows similarly to Example 2.6, as both involve linear systems of similar form, with $J^{\prime}$ in the reduced adjoint taking the place of $T^{(x)}$ in the basic adjoint.
Suppose then that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(x)};u}$-Lipshitz for all $x\in\Omega$ with both
$M_{T^{(x)}}:=\sup\{\|T^{(x)}(u,x)\|\mid u\in U,x\in\Omega\}<\infty\enskip\text{and}\enskip N_{S_{w}}:=\sup\{\|S_{w}(x)\|\mid x\in\Omega\}<\infty.$
Then the differential transformation Section 2.1Item (iii) holds due to
$\|\widetilde{F^{\prime}}(x^{k})$ $-F^{\prime}(x^{k})\|_{X^{*}}=\|w^{k+1}T^{(x)}(u^{k+1},x^{k})-w_{x^{k}}T^{(x)}(S_{u}(x^{k}),x^{k})\|_{X^{*}}$
$=\|[w^{k+1}-w_{x^{k}}]T^{(x)}(u^{k+1},x^{k})-w_{x^{k}}[T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})]\|_{X^{*}}$
$\leq\|T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\|w^{k+1}-w_{x^{k}}\|_{W}$
$+\|w_{x^{k}}\|_{W}\|T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}$
$\leq N_{S_{w}}L_{T^{(x)};u}\|u^{k+1}-S_{u}(x^{k})\|_{U}+M_{T^{(x)}}\|w^{k+1}-S_{w}(x^{k})\|_{W}.$
2.2 An error estimate
We now study the error of approximating of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$, for all $k\in\mathbb{N}$. To this end, we first develop a technical estimate for sequences of real numbers satisfying certain recursive inequalities. These abstract estimates will later be instantiated using the inner and adjoint tracking conditions introduced earlier. The main results are core estimates that allow us to isolate the contributions of initialisation and update errors, and thereby quantify the impact of inexact inner and adjoint solutions over multiple iterations on the differential approximations.
Lemma 2.8.
For some $\kappa_{u},\kappa_{w}>1$ and ${\mu_{u}},\pi_{u},\pi_{w}>0$ suppose $b_{k},c_{k},d_{k}\geq 0$ for all $k\geq 0$ satisfy
$\kappa_{u}b_{k+1}\leq b_{k}+\pi_{u}d_{k}\quad\text{and}\quad\kappa_{w}c_{k+1}\leq c_{k}+{\mu_{u}}b_{k+1}+\pi_{w}d_{k}.$
Then, letting $\iota_{k}:=\sum_{m=1}^{k}\kappa_{u}^{-m}\kappa_{w}^{-(k+1-m)}$ (understanding that $\iota_{0}=0$), for all $k\geq 0$ and $\alpha_{u},\alpha_{w}\geq 0$, we have
(9) $R^{k+1}(\alpha_{u},\alpha_{w}):=\alpha_{u}b_{k+1}+\alpha_{w}c_{k+1}$ $\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})b_{1}+\alpha_{w}\kappa_{w}^{-k}c_{1}$
$+\sum_{j=0}^{k-1}\bigl{(}\alpha_{u}\kappa_{u}^{-(k-j)}\pi_{u}+\alpha_{w}[\iota_{k-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(k-j)}\pi_{w}]\bigr{)}d_{j+1}.$
Proof 2.9.
For $k=1$, $b_{2}\leq\kappa^{-1}_{u}b_{1}+\kappa^{-1}_{u}\pi_{u}d_{1}\text{ and }c_{2}\leq\kappa^{-1}_{w}c_{1}+\kappa^{-1}_{w}{\mu_{u}}b_{2}+\kappa^{-1}_{w}\pi_{w}d_{1}$ by assumption. Multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, then summing up, observing to cancel the two instances of $\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}b_{2}$, establishes (9).
We then take $k=n+1$, and proceed by induction, assuming (9) to hold for $k=n$. Again, $b_{n+2}\leq\kappa^{-1}_{u}b_{n+1}+\kappa^{-1}_{u}\pi_{u}d_{n+1}\text{ and }c_{n+2}\leq\kappa^{-1}_{w}c_{n+1}+\kappa^{-1}_{w}{\mu_{u}}b_{n+1}+\kappa^{-1}_{w}\pi_{w}d_{n+1}$ by assumption. As in the case $k=1$, multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, and then summing up, yields
$R^{n+2}(\alpha_{u},\alpha_{w})=\alpha_{u}b_{n+2}+\alpha_{w}\kappa^{-1}_{w}c_{n+2}$ $\leq(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})b_{n+1}+\alpha_{w}\kappa^{-1}_{w}c_{n+1}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}])d_{n+1}.$
The first two terms on the right-hand side equal $R^{n+1}(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}},\alpha_{w}\kappa^{-1}_{w})$, so using (9) for $k=n$, we continue
$R^{n+2}(\alpha_{u},\alpha_{w})$ $\leq((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-n}+\alpha_{w}\kappa^{-1}_{w}\iota_{n}{\mu_{u}})b_{1}+\alpha_{w}\kappa^{-1}_{w}\kappa_{w}^{-n}c_{1}$
$+\sum_{j=0}^{n-1}\bigl{(}(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-(n-j)}\pi_{u}+\alpha_{w}\kappa^{-1}_{w}[\iota_{n-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(n-j)}\pi_{w}]\bigr{)}d_{j+1}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}])d_{n+1}$
$=(\alpha_{u}\kappa_{u}^{-(n+1)}+\alpha_{w}\mu_{u}(\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n}))b_{1}+\alpha_{w}\kappa_{w}^{-(n+1)}c_{1}$
$+\sum_{j=0}^{n}\bigl{(}\alpha_{u}\kappa_{u}^{-(n+1-j)}\pi_{u}+\alpha_{w}[(\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}){\mu_{u}}\pi_{u}+\kappa_{w}^{-(n+1-j)}\pi_{w}]\bigr{)}d_{j+1}.$
Here $\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}=\iota_{n+1-j}$, as by the definition of $\iota_{n+1}$, for any $n\geq 0$,
(10) $\iota_{n+1}=\sum_{m=1}^{n+1}\kappa_{u}^{-m}\kappa_{w}^{-(n+2-m)}=\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\sum_{m=1}^{n}\kappa_{u}^{-m}\kappa_{w}^{-(n+2-m)}=\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n},$
Thus we obtain (9) for $k=n+1$.
The next two lemmas form our core estimates. To simplify the estimates, recalling that $\kappa_{u},\kappa_{w}>1$, we observe that
(11) $p^{k}\iota_{k}\leq p^{-1}k(\kappa/p)^{-(k+1)}\quad\text{for}\quad\kappa:=\min(\kappa_{u},\kappa_{w})>1\text{ and any }p\in(0,\kappa).$
Thus, by sum formulae for arithmetic-geometric progressions [12, formula 0.113],
(12) $\sum_{k=0}^{n-1}p^{k}\iota_{k}\leq\sum_{k=0}^{\infty}p^{k}\iota_{k}\leq p^{-1}(\kappa/p-1)^{-2}=p(\kappa-p)^{-2}\quad\text{for all }n\in\mathbb{N}.$
Lemma 2.10.
Suppose Section 2.1 holds and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$. Then for any $x\in X$ and, $p\in(0,\kappa)$, we have
(13) $d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})+e_{p,k}(x),$
where, for $\psi_{j}:=\alpha_{u}\kappa_{u}^{-j}\pi_{u}+\alpha_{w}[\iota_{j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-j}\pi_{w}]$ and $\overline{\kappa}:=\max\{\kappa_{u},\kappa_{w}\}$, we set
(14) $\varsigma_{p}$ $:=\frac{\overline{\kappa}}{p}\sum_{j=0}^{\infty}p^{j}\psi_{j}\leq\frac{(\alpha_{u}\pi_{u}+\alpha_{w}\pi_{w})\kappa\overline{\kappa}}{p(\kappa-p)}+\frac{\alpha_{w}{\mu_{u}}\pi_{u}\overline{\kappa}}{p^{2}(\kappa-p)^{2}}\quad\text{and}$
(15) $e_{p,k}(x)$ $:=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}p^{k}}d_{U}^{2}(u^{1},S_{u}(x^{0}))+\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}p^{k}}d_{W}^{2}(w^{1},S_{w}(x^{0}))$
$+\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}b_{X}^{2}(x^{j+1},x^{j})-\varsigma_{p}^{2}b_{X}^{2}(x,x^{k}).$
Proof 2.11.
We prove for all $s\in\mathbb{R}$ that
(16) $4sd_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))-4s^{2}\leq\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})+e_{p,k}(x).$
Maximising the left hand side with respect to $s$ then establishes (13).
By the differential transformation Section 2.1Item (iii), we have
$d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq\alpha_{u}d_{U}(u^{k+1},S_{u}(x^{k}))+\alpha_{w}d_{W}(w^{k+1},S_{w}(x^{k}))=:R^{k+1}.$
Since $\{x^{n}\}_{n=0}^{k}\subset\Omega$, the inner and adjoint tracking Section 2.1Items (i) and (ii) give
$\kappa_{u}d_{U}(u^{k+1},S_{u}(x^{k}))$ $\leq d_{U}(u^{k},S_{u}(x^{k-1}))+\pi_{u}b_{X}(x^{k},x^{k-1})\quad\text{and}\quad$
$\kappa_{w}d_{W}(w^{k+1},S_{w}(x^{k}))$ $\leq d_{W}(w^{k},S_{w}(x^{k-1}))+{\mu_{u}}d_{U}(u^{k+1},S_{u}(x^{k}))+\pi_{w}b_{X}(x^{k},x^{k-1}).$
Thus, invoking Lemma 2.8 with $b_{k+1}=d_{U}(u^{k+1},S_{u}(x^{k}))$, $c_{k+1}=d_{W}(w^{k+1},S_{w}(x^{k}))$, as well as $d_{k+1}=b_{X}(x^{k+1},x^{k})$, we obtain
$R^{k+1}\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})d_{U}(u^{1},S_{u}(x^{0}))+\alpha_{w}\kappa_{w}^{-k}d_{W}(w^{1},S_{w}(x^{0}))+\sum_{j=0}^{k-1}\psi_{k-j}b_{X}(x^{j+1},x^{j}).$
Using Young’s inequality several times here, and adding the productive zero,
(17) $4sR^{k+1}\leq\frac{(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})^{2}}{\theta_{k}^{u}}d_{U}^{2}(u^{1},S_{u}(x^{0}))+\frac{(\alpha_{w}\kappa_{w}^{-k})^{2}}{\theta_{k}^{w}}d_{W}^{2}(w^{1},S_{w}(x^{0}))\\ +\sum_{j=0}^{k-1}\frac{\psi_{k-j}^{2}}{\theta_{k,j}}b_{X}^{2}(x^{j+1},x^{j})+4\left(\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}\right)s^{2}+\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})-\varsigma_{p}^{2}b_{X}^{2}(x,x^{k})$
for any $\theta_{k}^{u},\theta_{k}^{w},\theta_{k,j}>0$. Take $\theta_{k}^{u}=p^{k}\varsigma_{p}^{-1}\pi_{u}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})$, $\theta_{k}^{w}=p^{k}\varsigma_{p}^{-1}\pi_{w}\alpha_{w}\kappa_{w}^{-k}$, and $\theta_{k,j}=\varsigma_{p}^{-1}p^{k-j}\psi_{k-j}$. Observe from (10) that $\iota_{k}\leq\kappa_{w}\iota_{k+1}$. Hence $p^{k}\iota_{k}\leq(\kappa_{w}/p)p^{k+1}\iota_{k+1}$, and further, $p^{k}\psi_{k}\leq(\overline{\kappa}/p)p^{k+1}\psi_{k+1}$, where $\overline{\kappa}/p>1$. Now
$\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}=\frac{1}{\varsigma_{p}}\left(p^{k}\psi_{k}+\sum_{j=1}^{k}p^{j}\psi_{j}\right)\leq\frac{\overline{\kappa}}{\varsigma_{p}p}\sum_{j=0}^{k+1}p^{j}\psi_{j}\leq 1.$
Thus rearranging (17) establishes (16). Finally, the bound in (14) on $\varsigma_{p}$ follows from Eq. 12 and $\sum_{j=0}^{\infty}(p/\kappa)^{j}=1/(1-p/\kappa)=\kappa/(\kappa-p)$.
Lemma 2.12.
Let $\{e_{p,k}(x^{k+1})\}_{k=0}^{N-1}$ be as in Lemma 2.10 with $p\geq 1$. Then
$\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})$ $\leq\frac{d_{U}^{2}(u^{1},S_{u}(x^{0}))}{\pi_{u}}\bigg{(}\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg{)}+\frac{d_{W}^{2}(w^{1},S_{w}(x^{0}))}{\pi_{w}}\bigg{(}\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg{)}.$
Proof 2.13.
We split $p^{k}e_{p,k}(x^{k+1})=:A_{k}+B_{k}+C_{k}-D_{k}$ observing (15). Now
$\sum_{k=0}^{N-1}C_{k}=\sum_{k=0}^{N-1}p^{k}\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}b_{X}^{2}(x^{j+1},x^{j})=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{k=j+1}^{N-1}\psi_{k-j}b_{X}^{2}(x^{j+1},x^{j})\\ =\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}b_{X}^{2}(x^{j+1},x^{j})\leq\sum_{j=0}^{N-2}p^{j}\varsigma_{p}^{2}b_{X}^{2}(x^{j+1},x^{j})\leq\sum_{k=0}^{N-1}D_{k}.$
Moreover, using (12) and the sum formula for geometric series, we estimate that $\sum_{k=0}^{N-1}(A_{k}+B_{k})$ is less than the right-hand side of the claim.
2.3 Smoothness of differential estimates
In this section, we derive descent- and Lipschitz-type inequalities for the approximate differential ${\widetilde{F^{\prime}}(x^{k})}$, extending these classical smoothness concepts to account for differential errors under the tracking framework.
Assume that $X$ is a normed space with $X^{*}$ its dual space, and that $F^{\prime}(x^{k})\in X^{*}$ is the Fréchet differential of $F$ at $x^{k}$. We can then produce descent inequalities and Lipschitz estimates where $\widetilde{F^{\prime}}(x^{k})$ replaces $F^{\prime}(x^{k})$. We recall that if $F^{\prime}$ is $L$-Lipschitz, it then satisfies the descent inequality
(18) $\langle F^{\prime}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{L}{2}b_{X}^{2}(x,x^{k})^{2}\quad\text{for all}\quad x,x^{k}\in X$
for the choice $b_{X}(x,x^{k})=\|x-x^{k}\|_{X}$. We will, however, allow this inequality to be satisfied by any distance expression $d_{X}$ such that that $d_{X^{*}}$ is Young to $d_{X}$. By this we mean that both expression are one-homogeneous and
$\langle\tilde{x}^{*}-x^{*}|\tilde{x}-x\rangle_{X^{*},X}\leq\frac{1}{2}d_{X^{*}}^{2}(\tilde{x}^{*},x^{*})+\frac{1}{2}d_{X}^{2}(x^{*},x)\quad\text{for all}\quad x,\tilde{x}\in X\text{ and }x^{*},\tilde{x}^{*}\in X^{*}.$
In particular, we allow $d_{X}$ and $d_{X^{*}}$ to be given by (semi-)norms generated by operators. We treat the descent inequality with respect to them in the next Section 3.
Simply taking $\bar{x}=x^{k}$ in the theorem to follow, and combining with the descent inequality (18), we obtain the inexact descent inequality
(19) $\langle\widetilde{F^{\prime}}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}b_{X}^{2}(x,x^{k})-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
In convergence proofs of optimisation methods, it is frequently convenient to use the three-point descent inequality (see [7, Corollary 7.2] for the convex case, or [31, Appendix B] for the non-convex case)
(20) $\langle F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta}{2}d_{X}^{2}(x,\bar{x})-\frac{L}{2}b_{X}^{2}(x,x^{k})\quad\text{for all}\quad x,x^{k},\bar{x}\in X.$
Likewise combining the next theorem with this inequality, we obtain the inexact version
(21) $\langle\widetilde{F^{\prime}}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta-\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x})-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}b_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
In the next two theorems, while we still assume normed spaces, $F^{\prime}(x^{k})$ again no longer need be a Fréchet derivative, and can again be the arbitrary “target” of Section 2.1.
Theorem 2.14.
Suppose Section 2.1 holds, $X$ is a normed space, $X^{*}$ its dual with $d_{X^{*}}^{2}$ Young to $d_{X}^{2}$. Also let $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$, and pick $p\in[1,\kappa)$. Then, for $e_{p,k}$ and $\varsigma_{p}$ defined in Lemma 2.10, we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})<\infty,$
and, for any $\tilde{\gamma}>0$ and $x,\bar{x}\in X$,
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x})-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}b_{X}^{2}(x,x^{k})-\frac{1}{2\tilde{\gamma}}e_{p,k}(x).$
Proof 2.15.
By the Young relationship of $d_{X}$ and $d_{X}^{*}$,
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq-\frac{1}{2\tilde{\gamma}}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))-\frac{\tilde{\gamma}}{2}d_{X}^{2}(x,\bar{x}).$
Together with (13) of Lemma 2.10, this establishes the claimed inequality. Lemma 2.12 shows the boundedness of $\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})$.
Taking $x^{*}=F^{\prime}(x)$ in the following theorem, and combining with $F^{\prime}$ being $L$-Lipschitz with respect to the distances $d_{X^{*}}$ and $d_{X}$, we can get the Lipschitz-like property with error for $\widetilde{F^{\prime}}$,
$\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x))\leq\frac{(1+\vartheta)L}{2}d_{X}^{2}(x^{k},x)+\frac{1+\vartheta^{-1}}{2}e_{\mathrm{lip},k}.$
Theorem 2.16.
Suppose Section 2.1 holds, $d_{X^{*}}$ satisfies the triangle inequality, and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\in\mathbb{N}$. Then, for any $\vartheta>0$,
(22) $\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),x^{*})\leq\frac{1+\vartheta}{2}d_{X^{*}}^{2}(F^{\prime}(x^{k}),x^{*})+\frac{1+\vartheta^{-1}}{2}e_{\mathrm{lip},k}\quad\text{for all}\quad x^{*}\in X^{*},$
where
[…]
Moreover, if $C>0$ is a constant independent of $N$, then
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty\quad\text{whenever}\quad\sum_{j=0}^{N-1}b_{X}^{2}(x^{j+1},x^{j})<C.$
Proof 2.17.
We apply Lemma 2.10 with $p=1$ and $x=x^{k}$. With these choices $e_{1,k}(x^{k})=e_{\mathrm{lip},k}$, and (13) reads
$d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))\leq e_{\mathrm{lip},k}.$
With this, (22) follows after we use the triangle and Young’s inequalities to derive
$\frac{1}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),x^{*})\leq\frac{1+\vartheta}{2}d_{X^{*}}^{2}(F^{\prime}(x^{k}),x^{*})+\frac{1+\vartheta^{-1}}{2}d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k})).$
For the sum of $e_{\mathrm{lip},k}$, we start with the first two terms of (23). Using (12) and the sum formula for geometric series, we bound their sum over $k\in\{0,\ldots,N-1\}$ by
$d_{U}^{2}(u^{1},S_{u}(x^{0}))\frac{\varsigma_{1}}{\pi_{u}}\bigg{(}\frac{\alpha_{u}\kappa}{\kappa-1}+\frac{\alpha_{w}\mu_{u}}{(\kappa-1)^{2}}\bigg{)}+d_{W}^{2}(w^{1},S_{w}(x^{0}))\frac{\varsigma_{1}}{\pi_{w}}\bigg{(}\frac{\alpha_{w}\kappa}{\kappa-1}\bigg{)}.$
We have $\varsigma_{1}<\infty$ by Lemma 2.10. We sum the third term of (23) over $k\in\{0,\ldots,N-1\}$, and change the order of summation to obtain
$\varsigma_{1}\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}\psi_{k-j}b_{X}^{2}(x^{j+1},x^{j})=\varsigma_{1}\sum_{j=0}^{N-2}\left(\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}\right)b_{X}^{2}(x^{j+1},x^{j})\leq\frac{\varsigma_{1}^{2}}{\overline{\kappa}}\sum_{j=0}^{N-2}b_{X}^{2}(x^{j+1},x^{j}).$
Thus $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty$ whenever $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}b_{X}^{2}(x^{j+1},x^{j})<C$.
3 Operator-relative regularity
To facilitate treating primal-dual methods as forward-backward methods with respect to suitable operators in the next section, we introduce operator-relative smoothness and monotonicity concepts. That is, we will not use the standard norms of the space, but (semi-)norms of the form $\|x\|_{\Lambda}:=\sqrt{\langle x|\Lambda x\rangle_{X,X^{*}}}$ for a positive semi-definite self-adjoint $\Lambda\in\mathbb{L}(X;X^{*})$. This will be useful for the treatment of primal-dual methods as forward-backward methods with respect to appropriate (semi-)norms. We start with the relevant definitions in Section 3.1, and then prove the relevant operator-relative descent inequalities and three-point monotonicity in Section 3.2.
3.1 Definitions
For a self-adjoint positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$ on a normed space $X$, we say that the Gâteaux derivative $DF$ of $F:X\to\mathbb{R}$ is $\Lambda$-firmly Lipschitz if
$\langle DF(z)-DF(x)|h\rangle_{X^{*},X}\leq\|x-z\|_{\Lambda}\|h\|_{\Lambda}\quad(x,z,h\in X).$
This implies for any $h\in X$ with $\|h\|_{X}\leq 1$ that
(24) $\|DF(z)-DF(x)\|_{X^{*}}^{2}\leq\|z-x\|_{\Lambda}^{2}\|h\|_{\Lambda}^{2}\leq\|\Lambda\|_{\mathbb{L}(X;X^{*})}\|z-x\|_{\Lambda}^{2}\leq\|\Lambda\|_{\mathbb{L}(X;X^{*})}^{2}\|z-x\|_{X}^{2}.$
In particular, $DF$ is $\|\Lambda\|$-Lipschitz.
Likewise, we call $DF$ locally $\Gamma$-monotone in $\Omega\ni\bar{x}$ for a self-adjoint $\Gamma\in\mathbb{L}(X;X^{*})$ if
$\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle\geq q_{\Gamma}(z-\bar{x})\quad\quad\text{for}\quad q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}\quad(z\in\Omega).$
We do not at this stage assume $\Gamma$ to be positive semi-definite. We, however, call $\Gamma$ Young if there exists a self-adjoint positive semi-definite $|\Gamma|\in\mathbb{L}(X;X^{*})$ such that
$2\langle\Gamma x|z\rangle_{X^{*},X}\leq\|x\|_{|\Gamma|}^{2}+\|z\|_{|\Gamma|}^{2}\quad(x,z\in X).$
Likewise, we call $G$ $\Gamma$-subdifferentiable and $\partial G$ $\Gamma$-monotone if, respectively,
$G(\tilde{x})-G(x)\geq\langle q|\tilde{x}-x\rangle+\frac{1}{2}q_{\Gamma}(\tilde{x}-x)\quad\text{or}\quad\langle\tilde{q}-q|\tilde{x}-x\rangle\geq q_{\Gamma}(\tilde{x}-x)$
for all $q\in\partial G(x)$; $\tilde{q}\in\partial\tilde{G}(x)$, and $x,\tilde{x}\in X$.
Remark 3.1.
Aside from $\|\,\boldsymbol{\cdot}\,\|_{M}$, introduced in the next section, which presently needs to satisfy the Pythagoras’ identity, our work does not strictly depend on $\|\,\boldsymbol{\cdot}\,\|_{\Lambda}$, $\|\,\boldsymbol{\cdot}\,\|_{|\Gamma|}$ or $q_{\Gamma}$ being produced by operators. They could be arbitrary semi-norms and quadratics, if we introduced the formal calculus $q_{a\Lambda+b\Gamma}:=a\|\,\boldsymbol{\cdot}\,\|_{\Lambda}^{2}+bq_{\Gamma}$, etc., for $a,b\in\mathbb{R}$, and defined the concept of $\Gamma$ being Young appropriately. For simplicity, we have chosen to work with operators.
3.2 Estimates
We first prove a $\Lambda$-firmly Lipschitz descent lemma, as a generalisation of the basic descent inequality (18).
Lemma 3.2.
On a normed space $X$, suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Gâteaux derivative for a self-adjoint positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$. Then
(25) $F(x)-F(z)-\langle DF(z)|x-z\rangle_{X^{*},X}\leq\frac{1}{2}\|z-x\|_{\Lambda}^{2}.$
Proof 3.3.
By the mean value theorem and the assumed firm Lipschitz property,
$F(x)-F(z)-\langle DF(z)|x-z\rangle_{X^{*},X}=\int_{0}^{1}\langle DF(z+t(x-z))-DF(z)|x-z\rangle\,\mathrm{d}t\leq\int_{0}^{1}t\|x-z\|_{\Lambda}^{2}\,\mathrm{d}t.$
Integrating, the claim follows.
The next result provides an operator-relative generalisation of the three-point descent inequality (20). Compared to the treatment of the latter in [31, Appendix B], it is important that $x$ ($=x^{k+1}$ in the application to forward steps at $x^{k}$) is not a priori restricted to the neighbourhood $\Omega$ of $\Gamma$-monotonicity at $\bar{x}$.
Lemma 3.4.
On a normed space $X$, let $F:X\to\mathbb{R}$ and suppose $DF$ is $\Lambda$-Lipschitz for some $\Lambda\in\mathbb{L}(X;X^{*})$, and $\Gamma$-monotone at $\bar{x}\in X$ in a convex neighbourhood $\Omega\ni\bar{x}$ and a Young $\Gamma\in\mathbb{L}(X;X^{*})$. Then, for any $\beta>0$, for all $z\in\Omega$ and $x\in X$,
$\langle DF(z)|x-\bar{x}\rangle\geq F(x)-F(\bar{x})+\frac{1}{2}q_{\Gamma-\beta|\Gamma|}(x-\bar{x})-\frac{1}{2}q_{\Lambda+\beta^{-1}|\Gamma|-\Gamma}(x-z).$
Proof 3.5.
Similarly to the proof of the descent inequality in Lemma 3.2, the mean value theorem applied to $\varphi(t):=F(\bar{x}+t(z-\bar{x}))$, followed by the assumed local $\Gamma$-monotonicity of $DF$, and the Young property of $\Gamma$, establishes
$F$ $(\bar{x})-F(z)-\langle DF(z)|\bar{x}-z\rangle_{X^{*},X}$
$=\int_{0}^{1}\langle DF(z+t(\bar{x}-z))-DF(z)|\bar{x}-z\rangle\,\mathrm{d}t\geq\int_{0}^{1}tq_{\Gamma}(\bar{x}-z)\,\mathrm{d}t=\frac{1}{2}q_{\Gamma}(\bar{x}-z)$
$=\frac{1}{2}q_{\Gamma}(x-\bar{x})+\frac{1}{2}q_{\Gamma}(x-z)-\langle\Gamma(x-\bar{x})|x-z\rangle\geq\frac{1}{2}q_{\Gamma-\beta|\Gamma|}(x-\bar{x})+\frac{1}{2}q_{\Gamma-\beta^{-1}|\Gamma|}(x-z).$
Applying Lemma 3.2 and summing this inequality with the descent inequality it provides, we obtain the claim.
Finally, we provide a three-point monotonicity inequality that can be used instead of Lemma 3.4 when function values are not desired. The proof we present here slightly improves [7, Lemma 15.1] even in the Hilbert space scalar factor case.
Lemma 3.6.
On a normed space $X$, let $F:X\to\mathbb{R}$ and suppose $DF$ is $\Lambda$-Lipschitz for some $\Lambda\in\mathbb{L}(X;X^{*})$, and $\Gamma$-monotone for a self-adjoint $\Gamma\in\mathbb{L}(X;X^{*})$ in a neighbourhood $\Omega\ni\bar{x}$ for some $\bar{x}\in X$. Then, for any $\beta,\zeta>0$, for all $z\in\Omega$ and $x\in X$, with $\tilde{\Gamma}:=\Gamma-(\zeta/2)\Lambda$, we have
$\langle DF(z)-DF(\bar{x})|x-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}-\beta|\tilde{\Gamma}|}(x-\bar{x})-q_{\Lambda/(2\zeta)+\beta^{-1}|\tilde{\Gamma}|-\tilde{\Gamma}}(x-z).$
Proof 3.7.
Using both the $\Gamma$-monotonicity and the $\Lambda$-firmly Lipschitz property, and finishing with Young’s inequality, we obtain
$\langle DF(z)-DF(\bar{x})|x-\bar{x}\rangle_{X^{*},X}$ $=\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle_{X^{*},X}+\langle DF(z)-DF(\bar{x})|x-z\rangle_{X^{*},X}$
$\geq q_{\Gamma}(z-\bar{x})-\|z-\bar{x}\|_{\Lambda}\|x-z\|_{\Lambda}\geq q_{\tilde{\Gamma}}(z-\bar{x})-\frac{1}{2\zeta}\|x-z\|_{\Lambda}^{2}.$
Arguing for $q_{\tilde{\Gamma}}$ as in the proof of Lemma 3.4, we obtain the claim.
4 Nonconvex forward-backward type methods with inexact updates
We now need to prove the convergence of outer methods for the outer/overall problem (2), given estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ by inner and adjoint methods, the latter two satisfying the tracking theory of Section 2. In this section, we do this through a convergence theory for general inexact forward backward-type methods in a normed space $X$. Our treatment encompasses primal-dual methods, seen as forward-backward methods with respect to appropriate operator-relative (semi-)norms, discussed in the previous section. We introduce such methods in Section 4.1. Then in Section 4.2 we introduce abstract growth conditions, which we will use in Sections 4.4, 4.5, 4.6 and 4.7 to prove various forms of convergence. Before that, in Section 4.3, we will verify the growth inequalities for forward-backward and primal-dual algorithms that use the tracking theory of Section 2 for (single-loop) updates of an inner problem.
4.1 General inexact forward-backward type methods
For proper $F,G:X\to\overline{\mathbb{R}}$, consider the problem
$\min_{x\in X}F(x)+G(x).$
In this subsection, and in the examples of Section 4.3, $G$ will be convex and lower semicontinuous, and $F$ Fréchet differentiable, but the general theory of Sections 4.2, 4.4, 4.5, 4.6 and 4.7 will make no such assumption.
For an initial $x^{0}$, if $X$ is Hilbert, the iterates $\{x^{k}\}_{k=1}^{\infty}$ of the basic inexact forward-backward method are generated for some step length parameter $\tau>0$ and an estimate $\widetilde{\nabla F}(x^{k})$ of $\nabla F(x^{k})$ (not necessarily the one from Section 2) by
(26) $x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\widetilde{\nabla F}(x^{k})).$
In implicit form the method reads
$-\tau^{-1}(x^{k+1}-x^{k})\in\widetilde{\nabla F}(x^{k})+\partial G(x^{k+1}).$
We generalise this problem and method by considering for a skew-adjoint $\Xi\in\mathbb{L}(X;X^{*})$, i.e., $\Xi^{*}|X=-\Xi$, the problem of finding $x\in X$ satisfying
(27) $0\in H(x):=F^{\prime}(x)+\partial G(x)+\Xi x$
with the implicit method
[…]
Algorithms of the form (28) with an exact inclusion for ~$\tilde{\partial}_{k+1}$, cover many common splitting algorithms, such as Douglas–Rachford splitting (DRS) and the primal-dual proximal splitting (PDPS) of [4]; see [7, 31]. As we will see in the following examples, with an inexact inclusion, besides the inexact gradients of Section 2, the approach also covers inexact proximal maps and mismatched adjoints [17] in primal-dual methods. In exact proximal maps were used, e.g., in [33] for point source localisation in measure spaces.
Example 4.1 (Forward-backward splitting).
For forward-backward splitting with inexact $\widetilde{F^{\prime}}(x^{k})\approx F^{\prime}(x^{k})$, we take ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})$ with $M=\tau^{-1}\operatorname{Id}$ and $\Xi=0$.
Example 4.2 (Primal-dual proximal splitting).
On normed spaces $Z$ and $Y$, let $g:Z\to\overline{\mathbb{R}}$ and $h:Y^{*}\to\overline{\mathbb{R}}$ be convex, proper, and lower semicontinuous, $f:Z\to\mathbb{R}$ possibly non-convex but Fréchet differentiable, and $K\in\mathbb{L}(Z;Y^{*})$. Suppose $h=(h_{*})^{*}$ for some $h_{*}:Y\to\overline{\mathbb{R}}$, and consider the problem
(29) $\min_{z\in Z}\leavevmode\nobreak\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\leavevmode\nobreak\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y).$
If $f$ is convex, subject to the standard condition on the existence of $x_{0}\in\operatorname{int}\operatorname{dom}[h\circ K]\cap\operatorname{dom}[f+g]\neq\emptyset$ with $Kx_{0}\in\operatorname{int}\operatorname{dom}h$, [^1] Several relaxations are possible, include using the relative interior, or the formulas of [2]. the Fenchel–Rockafellar theorem [7, Theorem 5.11] gives rise to the necessary and sufficient first-order primal-dual optimality conditions
$0\in H(z,y)=\begin{pmatrix}\partial g(z)+f^{\prime}(z)+K^{*}y\\ \partial h_{*}(y)-Kz.\end{pmatrix}=F^{\prime}(z,y)+\partial G(z,y)+\Xi(z,y),$
where $F(z,y)=f(z)$, $G(z,y)=g(z)+h_{*}(y)$, and $\Xi=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$. If $f$ is nonconvex, the necessity can be shown through, e.g., Mordukhovich subdifferentials, and their compatibility with both convex subdifferentials and Fréchet derivatives; see, e.g., [7].
Pick step length parameters $\tau,\sigma>0$. With inexact gradients for $f$, the PDPS in Hilbert spaces then reads
(30) $\left\{\begin{array}[]{l}z^{k+1}:=\operatorname{prox}_{\tau g}(z^{k}-\tau\widetilde{\nabla f}(z^{k})-\tau K^{*}y^{k}),\\ y^{k+1}:=\operatorname{prox}_{\sigma h_{*}}(y^{k}+\sigma K(2z^{k+1}-z^{k})).\end{array}\right.$
When $f=j\circ S_{u}$ for $S_{u}$ a PDE solution operator, and we compute $\widetilde{\nabla f}$ following Examples 2.5 and 2.7, (30) becomes the algorithm presented in [16].
To extend (30) to general normed spaces, we write it in $X=Z\times Y$ in implicit form as (28) with ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$, where
(31) $\widetilde{F^{\prime}}(z^{k},y^{k}):=\begin{pmatrix}\widetilde{f^{\prime}}(z^{k})\\ 0\end{pmatrix}\quad\text{and}\quad M:=\begin{pmatrix}\tau^{-1}M_{z},&-K^{*}\\ -K&\sigma^{-1}M_{y}\end{pmatrix}$
for some self-adjoint positive semi-definite $M_{z}\in\mathbb{L}(Z;Z^{*})$ and $M_{y}\in\mathbb{L}(Y;Y^{*})$. For standard proximal maps in Hilbert spaces, $M_{z}=\operatorname{Id}$ and $M_{y}=\operatorname{Id}$. In that case, $M$ is self-adjoint and positive semi-definite when $\tau\sigma\|K\|^{2}\leq 1$, while the treatment of exact forward steps with respect to $f$ requires [^2] This is the requirement for gap estimates; for iterate estimates $\lambda/2$ in place of $\lambda$ is sufficient. In [35] an overall factor $4/3$ improvement is shown through an analysis that involves historical iterates. $\tau\lambda+\tau\sigma\|K\|^{2}\leq 1$ for $\lambda$ the Lipschitz factor of $f^{\prime}$ [31, 7, 13].
In normed spaces, we extend the above step length assumptions as follows. In the standard Hilbert setting with $M_{y}=\operatorname{Id}$ and $M_{z}=\operatorname{Id}$, we can take $K_{z}=K$ and $K_{y}=\operatorname{Id}$.
Assumption 4.2 (PDPS step length condition).
In the setting of Example 4.2, $M_{z}$ is positive semi-definite, and $K=K_{y}K_{z}$ for some $K_{z}\in\mathbb{L}(Z;V)$, $K_{y}\in\mathbb{L}(V;Y^{*})$, and a normed space $V$. Given $\lambda\geq 0$, the step length parameters $\tau,\sigma>0$ satisfy
$K_{y}K_{y}^{*}\leq M_{y}\quad\text{and}\quad\tau\lambda M_{z}+\tau\sigma K_{z}^{*}K_{z}\leq M_{z}.$
Lemma 4.3 (PDPS preconditioning operator).
If Section 4.1 holds, then $M$ is positive semi-definite and for any $\gamma_{z},\gamma_{y}\geq 0$ and $\gamma:=\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}/2$, we have
$\lambda\operatorname{diag}(M_{z},\;0)\leq M\quad\text{and}\quad\gamma M\leq\operatorname{diag}(\gamma_{z}M_{z},\;\gamma_{y}M_{y}).$
Proof 4.4.
By a simple application of Young’s inequality and Section 4.1, we have
$\|(z,y)\|_{M}^{2}=\tau^{-1}\|z\|_{M_{z}}+\sigma^{-1}\|y\|_{M_{y}}-2\langle K_{z}z|K_{y}^{*}y\rangle_{Y^{*},Y}\geq\|z\|^{2}_{\tau^{-1}M_{z}-\sigma K_{z}^{*}K_{z}}\geq\lambda\|z\|^{2}_{M_{z}}$
for any $x=(z,y)\in Z\times Y$. This establishes the first claimed inequality. The second follows by using Young’s inequality and Section 4.1 to establish
$\gamma\|(z,y)\|_{M}^{2}\leq\gamma\|z\|^{2}_{\tau^{-1}M_{z}+\sigma K_{z}^{*}K_{z}}+\gamma\|y\|^{2}_{\sigma^{-1}M_{y}+\sigma^{-1}K_{y}^{*}K_{y}}\leq\frac{2\gamma}{\tau}\|z\|^{2}_{M_{z}}+\frac{2\gamma}{\sigma}\|y\|^{2}_{M_{y}}.$
Remark 4.5 (Testing operators and accelerated methods).
In [31, 7], “testing operators” $Z_{k}\in\mathbb{L}(X^{*};X^{*})$ are used to encode convergence rates, and to prove “accelerated” $O(1/N^{2})$ rates for the PDPS under mere primal strong convexity. They could be incorporated into our treatise, however, for simplicity, we have chosen not to do this.
4.2 Inexact growth inequalities
We now make precise the approximate inclusion in (28). For the relevant definitions, we first define the Lagrangian gap functional
$\mathcal{G}(x;\bar{x}):=[F+G](x)-[F+G](\bar{x})-\langle\Xi x|\bar{x}\rangle_{X^{*},X}.$
Example 4.6.
For forward-backward splitting, $\mathcal{G}(x;\bar{x})=[F+G](x)-[F+G](\bar{x})$ is simply a function value difference.
Example 4.7.
For the PDPS of Example 4.2, with $x=(y,z)$, we expand
$\mathcal{G}(x;\bar{x})=\mathcal{L}(z,\bar{y})-\mathcal{L}(\bar{z},y)\quad\text{for}\quad\mathcal{L}(z,y):=[f+g](z)+\langle Kz|y\rangle-h_{*}(y).$
This is different from the true duality gap that arises from the Fenchel–Rockafellar theorem. For the latter no convergence results exist to our knowledge. In the convex case, if $0\in H(\bar{x})$, the Lagrangian gap is non-negative, however, it may be zero even if $0\not\in H(\bar{x})$, unlike for the true duality gap.
For subdifferential convergence, we will need an inexact descent inequality, as well as bounds on sums of the gaps.
Assumption 4.7.
$M\in\mathbb{L}(X;X^{*})$ is self-adjoint and positive semi-definite. Also,
  1. (i)
    For a set $\Omega\subset X$, $\eta>0$, and $\mathbb{L}(X;X^{*})\ni\breve{\Lambda}\leq 2(1-\eta)M$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$, for some errors $\varepsilon_{\mathrm{desc},k}\in\mathbb{R}$, for any $k\in\mathbb{N}$, we have
    (32) ~$\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{\mathrm{desc},k}.$
  2. (ii)
    The errors satisfy $r_{\mathrm{desc}}:=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}<\infty$.
  3. (iii)
    We have $x^{0}\in\Omega$, and for any $N\geq 1$, $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})\leq r_{\mathrm{desc}}$ implies $x^{N}\in\Omega$.
  4. (iv)
    For some $\tilde{\eta}<\eta$, we have
    $\inf_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\left(\mathcal{G}(x^{k+1};x^{k})+\tilde{\eta}\|x^{k+1}-x^{k}\|_{M}^{2}\right)>-\infty.$
Remark 4.8.
If $\Omega=X$, convergence will be global. In the examples of Section 2.1, $\Omega\neq X$ may arise from $S_{u}$, $G$, or $J$ being only locally Lipschitz continuously differentiable.
Example 4.9 (Descent inequality and gap sum for basic forward-backward splitting).
[…]
We will also need the approximations ~$\tilde{\partial}_{k+1}$ to become better as the distance between the iterates shrinks, in the sense of
Assumption 4.9.
For $H$ defined in (27), we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty\implies\lim_{k\to\infty}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}^{2}=0.$
This can be proved through Lipschitz differential estimates, as we discuss below.
For function value and iterate convergence, we cannot work with just the iterates: we need to assume properties with respect to a base point $\bar{x}\in X$, usually a solution. For iterate convergence, we assume the three-point monotonicity type estimate
(33) ~$\langle\tilde{\partial}_{k+1}-H(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}),$
for all $k\in\mathbb{N}$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ for an open neighbourhood $\Omega_{\bar{x}}$ of $\bar{x}$, a positive semi-definite self-adjoint $\breve{\Lambda}\in\mathbb{L}(X;X^{*})$, errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$, and a $\gamma\geq 0$.
For function value convergence, we need again a descent inequality similar to (32), now instantiated at the base point $\bar{x}$ instead of $x^{k}$. That is, for all $k\in\mathbb{N}$, we assume for some errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ that
(34) ~$\langle\tilde{\partial}_{k+1}|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
We write $\varepsilon_{\mathrm{desc},k}(\bar{x}):=\varepsilon_{k}(\bar{x})$ when we need draw a distinction to (33).
Example 4.10 (Growth conditions for basic forward-backward splitting).
[…]
The errors will also need to have a finite sum:
Assumption 4.10.
Given $\bar{x}\in X$, for some $\eta,\gamma\geq 0$ and $0\leq\breve{\Lambda}\leq(1-\eta)M$, either
  1. (a)
    (33) holds, and $\bar{x}\in H^{-1}(0)$; or
  2. (b)
    (34) holds, and $\inf_{x\in\Omega_{\bar{x}}}\mathcal{G}(x;\bar{x})\geq 0.$
Moreover, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for some $\delta>0$ and $p\geq 1$ with
[…]
4.3 Specific cases
We now explicitly verify Sections 4.2, 4.2 and 4.2 for both basic forward-backward splitting and the PDPS, as well as their inexact versions based on the estimation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ formed using inner and adjoint algorithms satisfying the tracking theory of Section 2.
We start with an abstract result for “operator-relative“ inexact forward-backward. It will form the basis of verification for both forward-backward and primal-dual methods. We will take
(36) $d_{X}(x,\tilde{x})=\|x-\tilde{x}\|_{M},\enskip d_{X^{*}}(x^{*},\tilde{x}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*}-\tilde{x}^{*})]^{1/2},\enskip b_{X}(x,\tilde{x})=\|x-\tilde{x}\|_{\Lambda}.$
Then the Fenchel–Young inequality and homogeneity ensure that $d_{X}^{*}$ is Young to $d_{X}$, as defined in Section 2.3. If $M$ is invertible, this gives $d_{X^{*}}(x^{*},\tilde{x}^{*})=\|x^{*}-\tilde{x}^{*}\|_{M^{-1}}$.
Theorem 4.11.
On a normed space $X$, for a self-adjoint and positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$, suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. For all $k\in\mathbb{N}$, construct $\widetilde{F^{\prime}}(x^{k})$ obeying Section 2.1 for the distances (36) and an $\Omega\subset X$. In (28), ensure ~$\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$. Let $\varsigma_{p}$ be as defined in (14), $\kappa$ in (11), and $r_{\mathrm{desc}}$ and $r_{p}$ in Sections 4.2 and 4.2. Then:
  1. (i)
    Section 4.2 holds for any $\tilde{\gamma}>0$, $\eta>\tilde{\eta}\geq 0$, $p\in[1,\kappa)$, and $\varepsilon_{\mathrm{desc},k}=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$ provided $\Xi=0$, $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, $\inf[F+G]>-\infty$, and
    $0\leq\breve{\Lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+\tilde{\gamma}M\leq 2(1-\eta)M.$
  2. (ii)
    Section 4.2 holds if $\Lambda\leq cM$ for a $c>0$.
Suppose further that $G$ is $\Gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\Gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}_{M}(\bar{x},\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\Gamma_{F},\Gamma_{G}\in\mathbb{L}(X;X^{*})$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. Set $\varepsilon_{k}(\bar{x})=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$ and define $r_{p}$ by (35). If $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ with $r_{p}<\delta^{2}/2$, then, for any $\eta\geq 0$:
  1. 3.
    Section 4.2 option Item (a) holds if, for $\tilde{\Gamma}_{F}:=\Gamma_{F}-(\zeta/2)\Lambda$ and $\zeta,\beta>0$,
    $(\gamma+\tilde{\gamma})M$ $\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|\quad\text{for a}\quad\gamma\geq(p-1)/2\quad\text{and}$
    $0\leq\breve{\Lambda}$ $:=(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+2(\beta^{-1}|\tilde{\Gamma}_{F}|-\tilde{\Gamma}_{F})\leq(1-\eta)M.$
  2. 4.
    Section 4.2 option Item (b) holds if $\Omega$ is convex, and, for $\beta>0$,
    $(\gamma+\tilde{\gamma})M$ $\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|\quad\text{for a}\quad\gamma\geq p-1\quad\text{and}$
    $0\leq\breve{\Lambda}$ $:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}\leq(1-\eta)M.$
Proof 4.12.
Item (i): By Lemma 3.2, we have
$\langle F^{\prime}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}.$
Combining this with Theorem 2.14 for $\bar{x}=x^{k}$ and $x=x^{k+1}$ establishes (compare (19))
$\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}(x^{k+1})$
with $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})<\infty$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$. Further combining with the subdifferentiability of $G$, we verify (32). Since we assume $\breve{\Lambda}\leq 2(1-\eta)M$ and take $\varepsilon_{\mathrm{desc},k}\propto e_{p,k}(x^{k+1})$, this verifies Section 4.2Items (i) and (ii). Because $\Xi=0$, Item (iii) requires $[F+G](x^{N})\leq r_{\mathrm{desc}}+[F+G](x^{0})$ to imply $x^{N}\in\Omega$. This holds whenever $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, as we have assumed. Likewise, we prove Item (iv) with the lower bound $\inf[F+G]-[F+G](x^{0})>-\infty$.
Item (ii): Theorem 2.16 with $x^{*}=F^{\prime}(x^{k+1})$ and $\vartheta=1$ establishes for any $C>0$ that
(37) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}\leq C\implies\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty$
and, since $F$ is $\Lambda$-firmly Lipschitz, together with (24) and $\Lambda\leq cM$ it establishes that
(38) $\frac{1}{2}\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}^{2}\leq c\|\Lambda\|\|x^{k+1}-x^{k}\|_{M}^{2}+e_{\mathrm{lip},k}.$
The antecedent of (37) implies $\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}\to 0$ via (38). Since $F$ is $\Lambda$-firmly Lipschitz, taking
$x_{k+1}^{*}=F^{\prime}(x^{k+1})-F^{\prime}(x^{k})+\tilde{\partial}_{k+1}\in F^{\prime}(x^{k+1})+\partial G(x^{k+1})+\Xi x^{k+1}=H(x^{k+1}),$
we estimate
$\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}^{2}\leq\|F^{\prime}(x^{k+1})-F^{\prime}(x^{k})\|_{X^{*}}^{2}.$
Thus the antecedent of Section 4.2 implies its consequent, verifying the assumption.
For the verification of both Items 3 and 4, we observe that (35) holds because the respective lower bounds $\gamma\geq(p-1)/2$ and $\gamma\geq p-1$ guarantee $p\leq\overline{p}_{\gamma}$, and we have explicitly assumed $1\leq p$ and $\delta^{2}>2r_{p}$, where, by our choice of $\varepsilon_{k}(\bar{x})=e_{p,k}(x^{k+1})/(2\tilde{\gamma})$, the definition of $r_{p}$ in Section 4.2, and Theorem 2.14 and
$r_{p}:=\sup_{N\in\mathbb{N}}p^{-N}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})/(2\tilde{\gamma})\leq\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}(x^{k+1})/(2\tilde{\gamma})<\infty.$
We have also explicitly assumed the remaining neighbourhood conditions of Section 4.2, as well as $0\leq\breve{\Lambda}\leq(1-\eta)M$, so only need to verify the respective (33) or (34).
Item 3: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma 3.6 and our assumption $\breve{\Lambda}\geq 0$, since $x^{k}\subset\Omega_{\bar{x}}$, we have
$\langle F^{\prime}(x^{k})-F^{\prime}(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\zeta^{-1}\Lambda+2(\beta^{-1}|\tilde{\Gamma}_{F}|-\tilde{\Gamma}_{F})}^{2}.$
Combining this with Theorem 2.14 and the definition of $\breve{\Lambda}$ gives
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
We verify (33) by combining this with the $\Gamma_{G}$-monotonicity of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|$.
Item 4: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma 3.4, since $x^{k}\subset\Omega_{\bar{x}}$, we have
$\langle F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq F(x^{k+1})-F(\bar{x})+\frac{1}{2}q_{\Gamma_{F}-\beta{|\Gamma_{F}|}}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}}^{2}.$
Combining this with Theorem 2.14 gives and the definition of $\breve{\Lambda}$ gives (compare (21))
$\begin{split}\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}&\geq F(x^{k+1})-F(\bar{x})\\ &+\frac{1}{2}q_{\Gamma_{F}-\beta{|\Gamma_{F}|}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).\end{split}$
We now verify (34) by combining this with the ($\Gamma_{G}$-strong) subdifferentiability of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$.
We now interpret the result for standard “scalar” forward backward splitting, as was as outer forward backward splitting for bilevel problems with inner and adjoint methods satisfying the tracking theory of Section 2.
Example 4.13 (Standard forward-backward splitting on a Hilbert space).
Let $X$ be a Hilbert space, and suppose $\nabla F$ is $L$-Lipschitz for some $L\geq 0$, Pick a step length parameter $\tau>0$. The exact forward-backward splitting
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\nabla F(x^{k})$
trivially satisfies the assumptions of the tracking theory in Section 2.1 with $\widetilde{F^{\prime}}\equiv F^{\prime}$, $S_{u}\equiv 0$, $S_{w}\equiv 0$, $\kappa_{u}=\kappa_{w}=1$, $\pi_{u}=\pi_{w}=\mu_{u}=0$ as well sa $\alpha_{u}=\alpha_{w}=0$. It follows that $e_{p,k}\equiv 0$, hence $\varepsilon_{\mathrm{desc},k}\equiv 0$ and $\varepsilon_{k}(\bar{x})\equiv 0$.
To model this algorithm in the operator-relative framework of Theorem 4.11, we take $M=\tau^{-1}\operatorname{Id}$ and $\Lambda=L\operatorname{Id}$. Then the condition $\breve{\Lambda}\leq 2(1-\eta)M$ with $\eta>0$ in Theorem 4.11Item (i) reduces to the standard step length condition $\tau L<2$.
With $G$ $\gamma_{G}$-strongly subdifferentiable for some $\gamma_{G}\geq 0$, and $F^{\prime}$ locally $\gamma_{F}$-monotone for some $\gamma_{F}\in\mathbb{R}$, taking $\Gamma_{G}=\gamma_{G}\operatorname{Id}$, $\Gamma_{F}=\gamma_{F}\operatorname{Id}$, and $\breve{\Lambda}=\breve{\lambda}\operatorname{Id}$ for some $\breve{\lambda}\geq 0$, the conditions in Item 4 reduce to finding $\tau,\beta>0$ and $\eta\geq 0$ such that
$0\leq\tau[L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]\leq 1-\eta\quad\text{and}\quad 0\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|].$
The conditions of Item 3 are analogous.
Example 4.14 (Single-loop/inexact outer forward-backward splitting on a Hilbert space).
Let $X$ be a Hilbert space, and suppose $\nabla F$ is $L$-Lipschitz for some $L\geq 0$. Construct $\widetilde{\nabla F}$ following Section 2.1, for example, using the inner methods of Examples 2.2, 2.4 and 2.5 and the adjoint methods and differential transformation of Examples 2.6 and 2.7. For a step length parameter $\tau>0$, consider the inexact forward-backward splitting
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\widetilde{\nabla F}(x^{k}))$
In the operator-relative framework of Theorem 4.11, we again take $M=\tau^{-1}\operatorname{Id}$ and $\Lambda=L\operatorname{Id}$. Then the condition in Theorem 4.11Item (i) reduces to $20\leq\tilde{\gamma}+\tau(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L<$ 2, where we again recall that $\varsigma_{p}$ is defined in (14).
Suppose then that $G$ is $\gamma_{G}$-strongly subdifferentiable for some $\gamma_{G}\geq 0$, and $F^{\prime}$ is locally $\gamma_{F}$-monotone for some $\gamma_{F}\in\mathbb{R}$. Then the conditions in Item 4 reduce to the bounds
$0\leq\tau[(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\gamma_{F}|-\gamma_{F})]\leq 1-\eta\quad\text{and}\quad 0\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|]-\tilde{\gamma}$
holding for some $\beta,\tilde{\gamma}>0$ and $\eta\geq 0$. Similarly we can translate Item 3.
Thus, we can expect the subdifferential convergence Section 4.2 to hold for small enough $\tau>0$, but stronger forms of convergence via Section 4.2 will require the differential approximation to be good enough that $\varsigma_{p}^{2}L$ is small with respect to the available strong monotonicity. Studying (14), the main constants that control this quantity are $\pi_{u}$ and $\pi_{w}$, which can always be made small if, instead a single iteration, we take sufficiently many iterations of the inner and adjoint solvers that satisfy Section 2.1Items (i) and (ii).
We can also translate Theorem 4.11 to the outer PDPS of Example 4.2 when inner and adjoint methods for estimating the smooth component $f$ of the problem (29) satisfy the tracking theory of Section 2. This is the content of the next theorem. If $f^{\prime}$ is locally $\gamma_{f}$-strongly monotone, its conditions Eqs. 40 and 41 for the satisfaction of Section 4.2 can be ensured by taking the free parameters $\tilde{\gamma},\zeta,\beta>0$ sufficiently small (to ensure Eqs. 40a and 41a), and then $\lambda>0$ sufficiently large (to ensure Eqs. 40b and 41b). Through the basic step length conditions Section 4.1, a large $\lambda$ will force the primal step length $\tau>0$ to small, and allow the dual step length $\sigma>0$ to be correspondingly large. The initial iterate should also satisfy the locality bound (39), where both $\lambda$ and $\tau$ appear, the latter through $M$ defined in (31). For Section 4.2 only the basic step length conditions are required besides the tracking estimates on $f$. Because $\Xi$ is not cyclically monotone (see [26, Chapter 24]), we see no way in general for the PDPS to satisfy Section 4.2Items (iii) and (iv). [^1] However, we could try to enforce the conditions, monitoring for convergence failure by setting expected bounds on $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})=[F+G](x^{N})-[F+G](x^{0})-\sum_{k=0}^{N-1}\langle\Xi x^{k+1}|x^{k}\rangle.$ In fact, if $\inf F+G>-\infty$, we only need to ensure that the latter sum term sum stays within chosen bounds, without having to calculate potentially costly function values.
Theorem 4.15 (PDPS with inexact $\widetilde{f^{\prime}}$; everything else exact).
Assume the setup of Example 4.2 with $g$ and $h_{*}$ convex, and $f$ with an $LM_{z}$-firmly Lipschitz Fréchet derivative for a $L\geq 0$. Suppose that Section 2.1 holds for $f$ in $\Omega\subset Z$ with [^2] We again recall that $d_{Z^{*}}(z^{*},\tilde{z}^{*})=\|z^{*}-\tilde{z}^{*}\|_{M^{-1}_{z}}$ if $M_{z}$ is invertible, in particular with $M_{z}=\operatorname{Id}$ for $Z$ a Hilbert space.
$d_{Z}(z,\tilde{z})=\|z-\tilde{z}\|_{M_{z}},\ d_{Z^{*}}(z^{*},\tilde{z}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M_{z}}^{2})^{*}(z^{*}-\tilde{z}^{*})]^{1/2},\ \text{and}\ b_{Z}(z,\tilde{z})=\|x-\tilde{x}\|_{LM_{z}}.$
Ensure the step length Section 4.1 for some $\tau,\sigma,\lambda>0$. Then
  1. (i)
    Section 4.2 holds.
Suppose further that $g$ and $h_{*}$ are, respectively, $\gamma_{g}M_{z}$ and $\gamma_{h_{*}}M_{y}$(-strongly) subdifferentiable for some $\gamma_{g},\gamma_{h_{*}}\geq 0$, and that $f^{\prime}$ is $\gamma_{f}M_{z}$-monotone in $\Omega_{\bar{z}}:=\Omega\ni\bar{z}$ for some $\gamma_{f}\geq 0$. Let $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$ and $\Omega_{\bar{x}}:=\Omega_{\bar{z}}\times\operatorname{dom}h_{*}$. Suppose $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. Set $\varepsilon_{k}(\bar{x})=e_{p,k}(z^{k+1})/(2\lambda\tilde{\gamma})$, and define $r_{p}$ by (35). If
[…]
then
  1. 2.
    Section 4.2 option Item (a) holds if, for $\tilde{\gamma}_{f}:=\gamma_{f}-(\zeta/2)L$ and some $\beta,\zeta>0$ and $\eta\geq 0$,
    (40a) $(p-1)/2\leq\gamma$ $:=\min\{(\gamma_{g}+\tilde{\gamma}_{f}-\beta|\tilde{\gamma}_{f}|)\tau,\gamma_{h_{*}}\sigma\}/2-\tilde{\gamma}\quad\text{and}$
    (40b) $0\leq\breve{\lambda}$ $:=\zeta^{-1}L+2(\beta^{-1}|\tilde{\gamma}_{f}|-\tilde{\gamma}_{f})+2\varsigma_{p}^{2}\tilde{\gamma}^{-1}L\leq(1-\eta)\lambda.$
  2. 3.
    Section 4.2 option Item (b) holds if $\Omega_{\bar{z}}$ is convex, and, for some $\beta>0$ and $\eta\geq 0$,
    (41a) $p-1\leq\gamma$ $:=\min\{(\gamma_{g}+\gamma_{f}-\beta|\gamma_{f}|)\tau,\gamma_{h_{*}}\sigma\}/2-\tilde{\gamma}\quad\text{and}$
    (41b) $0\leq\breve{\lambda}$ $:=L+\beta^{-1}|\gamma_{f}|-\gamma_{f}+\varsigma_{p}^{2}\tilde{\gamma}^{-1}L\leq(1-\eta)\lambda.$
Proof 4.16.
$F^{\prime}$ is $\Lambda$-firmly Lipschitz and $\Gamma_{F}$-monotone, and $G$ is $\Gamma_{G}$-strongly convex for
[…]
Adopting the distances (36) for $X$ and $X^{*}$, Lemma 4.3 shows that
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}((z^{*},0))=\sup_{(z,y)}\langle z^{*}|z\rangle-\frac{1}{2}\|(z,y)\|_{M}^{2}\leq\sup_{z}\langle z^{*}|z\rangle-\frac{\lambda}{2}\|z\|_{M_{z}}^{2}=\frac{1}{\lambda}(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M_{z}}^{2})^{*}(z^{*}).$
Combining this with Section 2.1 for $f$ and $\widetilde{f^{\prime}}$ in $\Omega\subset Z$, we see that Section 2.1 holds in $\Omega\times\operatorname{dom}h_{*}\subset X$ for $F$ and $\widetilde{F^{\prime}}$ defined in Example 4.2 with $\alpha_{u}$ and $\alpha_{w}$ divided by $\lambda$ compared to the case of $f$ and $\widetilde{f^{\prime}}$. This has the effect of dividing $e_{p,k}$ by $\lambda$. Our claims thus follow if we prove the remaining assumptions of Theorem 4.11.
Item (i): Lemma 4.3 proves $\Lambda\leq(L/\lambda)M$. Clearly $\Lambda\geq 0$. Now we use Theorem 4.11Item (ii).
[…]
$=\breve{\Lambda}:=\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}+\varsigma_{p}^{2}\tilde{\gamma}^{-1}\Lambda\quad\text{and}$
$\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$ $=\operatorname{diag}(\gamma_{z}M_{z},\;\gamma_{y}M_{y})\geq(1/2)\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}M\geq(\gamma+\tilde{\gamma})M.$
Taking $\delta:=\lambda\delta_{z}$, (39) implies, as required, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $2r_{p}<\delta^{2}$. By (42), we have $\mathbb{O}_{M}(\bar{x},\delta)\subset\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\times\operatorname{dom}h_{*}\subset\Omega_{\bar{z}}\times\operatorname{dom}h_{*}=\Omega_{\bar{x}}$. By construction and assumption, we have $\breve{\Lambda}\geq 0$. The claim now follows from Theorem 4.11Item 4.
Item 2: completely analogous to Item 3, observing that $\tilde{\Gamma}_{F}=\operatorname{diag}(\tilde{\gamma}_{f}M_{z},0)$.
We finally consider adjoint mismatch as in [17], keeping everything else exact.
Theorem 4.17 (PDPS with adjoint mismatch).
Assume the setup of Example 4.2 with $\tau\sigma\|K\|^{2}\leq 1$ and, for simplicity, $f=0$ and Hilbert $Z$ and $Y$. Suppose $\operatorname{dom}h_{*}$ is bounded, and that $g$ and $h_{*}$ are, respectively, $\gamma_{g}$- and $\gamma_{h_{*}}$-strongly convex for some $\gamma_{g}>0$ and $\gamma_{h_{*}}\geq 0$. Let $\gamma:=\min\{\gamma_{g}\tau/4,\gamma_{h_{*}}\sigma/2\}$. In the PDPS (30), not able to compute $K^{*}$, replace it with a “mismatched” adjoint $K^{*\approx}$. Then, for any $\bar{x}\in Z\times Y$ and $p\in(1,1+2\gamma]$, Section 4.2Item (a) holds with $\breve{\Lambda}=0$, $\Omega_{\bar{x}}=Z\times Y$, $\delta=\infty$, $r_{p}\leq\varepsilon/(1-p)$, and
$\varepsilon_{k}(\bar{x})=\frac{1}{2\gamma_{g}}\|(K^{*\approx}-K^{*})y^{k}\|_{Z}^{2}\leq\varepsilon:=\frac{1}{2\gamma_{g}}(\|K^{*\approx}-K^{*}\|\operatorname{diam}\operatorname{dom}h_{*})^{2}.$
Proof 4.18.
With $M$, $G$, and $F$ given by Example 4.2, the abstract algorithm (28) reads
$-M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}=x_{k+1}^{*}+((K^{*\approx}-K^{*})y^{k},0)\quad\text{for a}\quad x_{k+1}^{*}\in H(x^{k+1}).$
Here $H$ is defined in (27). Using Lemma 4.3 in the final step, we estimate
~$\langle\tilde{\partial}_{k+1}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ =~$=\langle\tilde{\partial}_{k+1}-x_{k+1}^{*}|x^{k+1}-\bar{x}\rangle_{X^{*},X}+\langle x_{k+1}^{*}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$
$\geq\langle(K^{*\approx}-K^{*})y^{k},z^{k+1}-\bar{z}\rangle+\gamma_{g}\|z^{k+1}-\bar{z}\|_{Z}^{2}+\gamma_{h_{*}}\|y^{k+1}-\bar{y}\|_{Y}^{2}$
$\geq\frac{\gamma_{g}}{2}\|z^{k+1}-\bar{z}\|_{Z}^{2}+\gamma_{h_{*}}\|y^{k+1}-\bar{y}\|_{Y}^{2}-\frac{1}{2\gamma_{g}}\|(K^{*\approx}-K^{*})y^{k}\|_{Z}^{2}$
$\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Therefore, (33) holds with the stated choices. Moreover, we have $\sum_{k=0}^{N-1}p^{k-N}\leq 1/(p-1)$ for any $p\in(1,1+2\gamma]$, verifying (35) and consequently Section 4.2Item (a).
Remark 4.19 (Stochastic optimisation methods).
These can be approached through lifting: we take $X$ as a space of random variables on a space $\mathscr{X}$, set $F(x)=\mathbb{E}[\mathscr{F}\circ x]$ for a function $\mathscr{F}$ on $\mathscr{X}$, $Mx:=\mathbb{E}^{*}\mathscr{M}\mathbb{E}[x]$ for $\mathscr{M}\in\mathbb{L}(\mathscr{X};\mathscr{X}^{*})$, etc., where $\mathbb{E}$ is the expectation. Without inexactness, this produces parallel copies of the very same optimisation method for every random event $\omega$. We then model the random choices made on every step through inexactness, subject to Sections 4.2, 4.2 and 4.2.
4.4 Convergence of subdifferentials and quasi-monotonicity of values
We first show the potentially global convergence of subdifferentials; see Remark 4.8. When $\Xi=0$, this could be followed by the Kurdyka–Łojasiewicz property to show function value convergence, and, afterwards, either by a growth condition or, in finite dimensions, a finite-length argument based on (43) and [1, proof of Lemma 2.6] to show iterate convergence. As the property can easily be verified only in finite dimensions (for semi-algebraic functions), we prefer a more direct approach.
Theorem 4.20.
If Section 4.2 holds, then $x^{k}\in\Omega$ and
(43) $\mathcal{G}(x^{k+1};x^{k})+\eta\|x^{k+1}-x^{k}\|_{M}^{2}\leq\varepsilon_{\mathrm{desc},k}\quad\text{for all}\quad k\in\mathbb{N}.$
If, moreover, Section 4.2 holds, then also $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$
Proof 4.21.
By the implicit algorithm (28), the properties of Fenchel conjugates (e.g., [7, Lemma 5.7]) and $-M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)(x^{k+1}-x^{k})$, we have
(44) $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})=2\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)^{*}(\tilde{\partial}_{k+1})=\|x^{k+1}-x^{k}\|_{M}^{2}=-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}.$
If $\{x^{j}\}_{j=0}^{N-1}\subset\Omega$, Section 4.2Item (i) thus yields for all $k=0,\ldots,N-1$ that
(45) $\mathcal{G}(x^{k+1};x^{k})$ $=\mathcal{G}(x^{k+1};x^{k})-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}-\|x^{k+1}-x^{k}\|_{M}^{2}$
$\leq\varepsilon_{\mathrm{desc},k}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{2M-\breve{\Lambda}}^{2}\leq\varepsilon_{\mathrm{desc},k}-\eta\|x^{k+1}-x^{k}\|_{M}^{2}.$
Summing over all such $k$, and using Section 4.2Item (ii), it follows
(46) $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})+\sum_{k=0}^{N-1}\eta(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})=\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})+\sum_{k=0}^{N-1}\eta\|x^{k+1}-x^{k}\|_{M}^{2}\leq r_{\mathrm{desc}}.$
From Section 4.2Item (iii), it now follows that $x^{N}\in\Omega$. Since, by the same assumption, $x^{0}\in\Omega$, induction establishes (43) and $x^{k}\in\Omega$ for all $k\in\mathbb{N}$. Using Section 4.2Item (iv) in (46), we, moreover, deduce $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ and $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})\to 0$. Let $c\geq\|M\|_{\mathbb{L}(X;X^{*})}$. By $\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\leq c\|\,\boldsymbol{\cdot}\,\|_{X}^{2}$ and the properties of conjugates (e.g., [7, Lemmas 5.4 and 5.7]),
$\frac{4}{c}\|\tilde{\partial}_{k+1}\|_{X^{*}}^{2}=c\|2\tilde{\partial}_{k+1}/c\|_{X^{*}}^{2}=(c\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(2\tilde{\partial}_{k+1})\leq(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1}).$
Thus also ~$\|\tilde{\partial}_{k+1}\|_{X^{*}}\to 0$. Section 4.2 proves that $\inf_{x^{*}\in H(x^{k+1})}\|\tilde{\partial}_{k+1}-x^{*}\|_{X^{*}}\to 0$. Hence an application of the triangle inequality establishes $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$.
Example 4.22 (Forward-backward splitting).
For the (inexact) forward-backward splitting of Examples 4.1, 4.13 and 4.14, the theorem establishes the monotonicity of function values, as well as the convergence of subdifferentials to zero, $\inf_{x^{*}\in\partial G(x^{k+1})}\|F^{\prime}(x^{k+1})+x^{*}\|\to 0$.
Remark 4.23 (Bregman divergences).
The argument of Theorem 4.20 extends to algorithms where the $M$-seminorm for $M\in\mathbb{L}(X;X^{*})$ is replaced by a Bregman divergence $B_{M}$ generated by some convex $M:X\to\overline{\mathbb{R}}$. Recalling the definition
(47) $B_{M}^{\omega}(x,z):=M(z)-M(x)-\langle\omega|z-x\rangle\quad(\omega\in\partial M(x);x,z\in X),$
in the algorithm (28), we would impose ~$\tilde{\partial}_{k+1}=\omega^{k+1}-\omega^{k}\in\partial_{2}B_{M}^{\omega^{k}}(x^{k},x^{k+1})$ for a given $\omega^{k}\in\partial M(x^{k})$ and some $\omega^{k+1}\in\partial M(x^{k+1})$. Using the Fenchel–Young identity, we could then replace (44) by
$B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})$ $=B_{M}^{\omega^{k}}(x^{k},x^{k+1})=M(x^{k+1})-M(x^{k})-\langle\omega^{k}|x^{k+1}-x^{k}\rangle$
$\leq\langle\omega^{k+1}-\omega^{k}|x^{k+1}-x^{k}\rangle=-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}.$
Thus repeating the arguments of the theorem would establish both $B_{M}^{\omega^{k}}(x^{k},x^{k+1})\to 0$ as well as $B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})=B_{M^{*}}^{x^{k+1}}(\omega^{k}+\tilde{\partial}_{k+1},\omega^{k})\to 0$. A variant of Section 4.2 could then establish a form of convergence for $H(x^{k+1})$.
4.5 Non-escape, quasi-Féjer monotonicity, linear convergence
The next lemma is essential for all our strong convergence results. The proof is standard; see, e.g., [7, Chapter 15] for the case $\varepsilon_{k}(\bar{x})=0$ and $\Xi=0$. Observe that (48) with the triangle inequality may be used to again prove Section 2.1Item (i) for multilevel methods.
Lemma 4.24.
Suppose Section 4.2 holds at $\bar{x}\in X$. Then $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, and the sequence is ($p$-strongly) quasi-Féjer, i.e.,
(48) $\frac{p}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}\leq\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}+\varepsilon_{k}(\bar{x}).$
Moreover, $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ if $\eta>0$.
Proof 4.25.
We first treat Section 4.2 option Item (a). Fix $N\in\mathbb{N}$ and suppose $\{x^{j}\}_{j=0}^{N-1}\subset\Omega_{\bar{x}}$. Observe that $\langle\Xi x|x\rangle=0$ for all $x\in X$ by the skew-adjointness of $\Xi$. Since $0\in H(\bar{x})$, using (33) in the implicit algorithm (28), we thus get
$-\langle M(x^{k+1}-x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ $\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x})$
for all $k\in\{0,\ldots,N-1\}$. By $\breve{\Lambda}\leq(1-\eta)M$ and the Pythagoras’ identity (see [34, (2.3)])
$\langle M(x-z)|x-\bar{x}\rangle=\frac{1}{2}\|x-z\|_{M}^{2}+\frac{1}{2}\|x-\bar{x}\|_{M}^{2}-\frac{1}{2}\|z-\bar{x}\|_{M}^{2}\quad(x,z,\bar{x}\in X),$
we obtain
(49) $\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}\geq\frac{\eta}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1+2\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Using $1+2\gamma\geq p$, multiplying by $p^{k}$, and summing over $k=0,\ldots,N-1$ yields
(50) $\frac{1}{2}\|x^{0}-\bar{x}\|_{M}^{2}+\sum_{k=0}^{N-1}p^{k}\varepsilon_{k}(\bar{x})\geq\sum_{k=0}^{N-1}\frac{\eta p^{k}}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{p^{N}}{2}\|x^{N}-\bar{x}\|_{M}^{2}.$
Multiplying by $p^{-N}\leq 1$ and using $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and (35), it follows
(51) $\frac{\delta^{2}}{2}=\frac{\delta^{2}-2r_{p}}{2}+r_{p}>\sum_{k=0}^{N-1}\frac{\eta p^{k-N}}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1}{2}\|x^{N}-\bar{x}\|_{M}^{2}.$
Hence $x^{N}\in\mathbb{O}_{M}(\bar{x},\delta)$. Since $x^{0}\in\Omega_{\bar{x}}$ by Section 4.2, an inductive argument shows that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, justifying the above steps. Finally, (49) shows (48), while $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ follows from (51) and $\eta>0$.
Regarding option Section 4.2Item (b), arguing as above with (34) in place of (33), we get in place of (49) the estimate
(52) $\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\eta}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1+\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Using $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$, we proceed (with $1+\gamma\geq p$) as in option Item (a) to establish (51), and from there onwards.
A closer look at (50) immediately yields linear convergence if $p>1$ and we remove $p^{-N}$ from (35).
Corollary 4.26.
Suppose Section 4.2 holds at $\bar{x}\in X$ with $p>1$ and the inequality in (35) strengthened to
(53) $\frac{1}{2}\delta^{2}>\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}\varepsilon_{k}(\bar{x})<\infty.$
Then $\|x^{N}-\bar{x}\|_{M}^{2}\to 0$ at the rate $O(p^{-N})$.
Example 4.27.
From the proof of Theorem 4.11, it is clear that (53) holds for the exact and inexact forward backwards methods of Examples 4.13 and 4.14. Since Theorem 4.15 is based on Theorem 4.11, it also holds for exact and inexact primal-dual methods. For $p>1$, sufficient second-order growth (strong convexity), is still required.
4.6 Local convergence of function values
We now proceed to function values and duality gaps. The idea of possibly assuming both Section 4.2Item (a) and a relaxed version of Item (b), as an alternative to just the latter, is to be able to study descent at non-minimising critical points. For simplicity, we only treat sublinear convergence.
Theorem 4.28.
Suppose Section 4.2 holds at $\bar{x}\in X$ and, for a non-empty set $\hat{X}\subset X$, Eq. 34 holds for all $\hat{x}\in\hat{X}$ with $\breve{\Lambda}=\breve{\Lambda}_{\hat{x}}\leq M$, $\gamma=\gamma_{\hat{x}}\geq 0$, and $\Omega_{\hat{x}}\supset\mathbb{O}_{M}(\bar{x},\delta)$. Then
(54) $\sup_{\hat{x}\in\hat{X}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\leq\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}(\hat{x})\right)\quad\text{for all}\quad N\in\mathbb{N}.$
If $\Xi=0$ and Section 4.2 holds [^3] Since the proof of the present Theorem 4.28 shows that $x^{k}\in\mathbb{O}_{M}(\hat{x},\delta)$ for all $k\in\mathbb{N}$, to prove the required (43), it would be enough to assume that just Section 4.2Item (i) holds with $\Omega\supset\mathbb{O}_{M}(\hat{x},\delta)$., then, for all $N\in\mathbb{N}$,
(55) $[F+G](x^{N})\leq\inf_{\hat{x}\in\hat{X}}[F+G](\hat{x})+\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2N}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\left(\frac{1}{N}\varepsilon_{\mathrm{desc},k}(\hat{x})+\frac{k+1}{N}\varepsilon_{\mathrm{desc},k}\right)\right).$
Proof 4.29.
Lemma 4.24 shows for all $k\in\mathbb{N}$ that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\bigcap_{\hat{x}\in\hat{X}}\Omega_{\hat{x}}$. Hence, for any $\hat{x}\in\hat{X}$, we may follow the proof of the lemma for case Item (b) of Section 4.2 to establish (52) for $\bar{x}=\hat{x}$. To reach this point, the assumption $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$ was not yet needed. Now, summing (52) over $k=0,\ldots,N-1$, we obtain
(56) $\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}(\hat{x})\geq\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})+\frac{1}{2}\|x^{N}-\hat{x}\|_{M}^{2}.$
Taking the supremum over $\hat{x}\in\hat{X}$, this establishes (54).
Suppose then that $\Xi=0$ and Section 4.2 holds. Theorem 4.20 now establishes (43), i.e., the quasi-monotonicity $[F+G](x^{k+1})\leq[F+G](x^{k})+\varepsilon_{\mathrm{desc},k}.$ Repeatedly using this and $\mathcal{G}(x^{k+1};\hat{x})=[F+G](x^{k+1})-[F+G](\hat{x})$ in (56), and dividing by $N$, we obtain (55).
We next specialise the result to the PDPS of Example 4.2. Besides inexactness, as a novelty compared to [5, 6, 18, 11], subject to $h_{*}$ having a bounded domain, we get an estimate on the convex envelope of the objective, i.e., the Fenchel biconjugate. In non-reflexive spaces, we define the latter as a function in $X$ instead of $X^{**}$ by taking first the conjugate and then the equivalently defined preconjugate: $h^{**}:=(h_{*})_{*}$.
Corollary 4.30.
Assume the setup of Examples 4.2 and 4.1 for some $\tau,\sigma,\lambda>0$, as well as that Theorem 4.15Items 2 and 3 hold for $p=1$ at some $\bar{z}\in Z$ with $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Also suppose that that $\operatorname{dom}h_{*}$ is bounded, $0\in H(\bar{x})$ for some $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$, and that the condition (39) on the initial iterate holds. Then, for the ergodic iterates $\tilde{z}^{N}:=\frac{1}{N}\sum_{k=0}^{N-1}z^{k}$, for all $N\in\mathbb{N}$, we have
$[f+g+h\circ K]^{**}(\tilde{z}^{N})\leq[f+g+h\circ K](\bar{z})+\sup_{\hat{y}\in\operatorname{dom}h_{*}}\frac{1}{2N}\|(z^{0},y^{0})-(\bar{z},\hat{y})\|_{M}^{2}+\frac{\sum_{k=0}^{N-1}e_{1,k}(z^{k+1})}{2\tilde{\gamma}\lambda N}.$
Here $[f+g+h\circ K](\bar{z})=[f+g+h\circ K]^{**}(\bar{z})$ if $\bar{z}$ is a global minimiser of $f+g+h\circ K$.
Proof 4.31.
Theorem 4.15Item 2 proves Section 4.2 option Item (a) at $\bar{x}$. Likewise, Theorem 4.15Item 3 shows (34) and $\Omega_{\hat{x}}:=\Omega_{\bar{z}}\times\operatorname{dom}h_{*}\supset\mathbb{O}_{M}(\bar{x},\delta)$ at any $\hat{x}\in\hat{X}:=\{\bar{z}\}\times\operatorname{dom}h_{*}$. Theorem 4.28 now establishes (54), whose left-hand-side we still have to estimate.
With the expression of Example 4.7 for the gap, we expand and estimate using the definition of the Fenchel (bi)conjugate and $h^{**}=h$ as well as $[f+g]^{**}\leq f+g$ that
$\mathcal{G}(x^{k+1};\bar{x})$ $=([f+g](z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y}))-([f+g](\bar{z})+\langle K\bar{z}|y^{k+1}\rangle-h_{*}(y^{k+1}))$
$\geq\bigl{(}[f+g]^{**}(z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y})\bigr{)}-N[f+g+h\circ K](\bar{z}).$
Summing over $k\in\{0,\ldots,N-1\}$, taking the supremum over $\hat{y}\in\operatorname{dom}h_{*}$, and using Jensen’s inequality, therefore
$\sup_{\hat{y}\in\operatorname{dom}h_{*}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\bar{x})\geq N[(f+g)^{**}+h\circ K](\tilde{z}^{N})-N[f+g+h\circ K](\bar{z}).$
[…]
$f+g+h\circ K\geq[f+g+h\circ K]^{**}=((f+g)^{*}\mathop{\Box}[h\circ K]^{*})^{*}=(f+g)^{**}+h\circ K.$
Moreover, the inequality is an equality at a global minimiser (or if $f$ is convex). Now the claim follows from (54).
Remark 4.32.
Taking $p>1$ in the proof of Corollary 4.30, linear convergence rates could be obtained as in Corollary 4.26 for the iterates.
4.7 Weak convergence
We next prove the weak convergence of the iterates. We call the self-adjoint and positive semi-definite preconditioner $M\in\mathbb{L}(X;X^{*})$ admissible for weak convergence if $\|x^{k}\|_{M}\to 0$ implies $Mx^{k}\to 0$.
Example 4.33.
Suppose $M=A^{*}A$ for some $A\in\mathbb{L}(X;V)$ for a Hilbert space $V$. Then the $M$-semi-norm convergence clearly implies $Ax^{k}\to 0$, and consequently $Mx^{k}\to 0$. Thus $M$ is weak-* admissibility. In Hilbert spaces, every positive semi-definite self-adjoint operator has such a square root $A$ with $V=X$. For a convolution-based construction in the space of Radon measures, see [34, Theorem 2.4].
Theorem 4.34.
Suppose Sections 4.2 and 4.2 hold with $p=1$ and $\eta>0$ at some $\bar{x}=\hat{x}\in H^{-1}(0)$, and that either Section 4.2Item (a) or Item (b) (only the item, not the entire assumption) holds with $\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$ and $\sum_{k=0}^{\infty}\varepsilon_{k}(\hat{x})<\infty$ at all $\hat{x}\in\hat{X}:=H^{-1}(0)\cap\mathbb{O}_{M}(\bar{x},\delta)$. Also suppose that the preconditioner $M$ is admissible for weak convergence, and $F$ is either convex or $F^{\prime}$ is weak-to-strong continuous. Then $x^{k}\mathrel{\rightharpoonup}\hat{x}$ weakly for some $\hat{x}\in\hat{X}$.
Proof 4.35.
Lemma 4.24 proves that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, as well as that $\sup_{N\in N}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$. The latter establishes $\|x^{k+1}-x^{k}\|_{M}\to 0$, and through admissibility for weak convergence, and (28), that ~$\tilde{\partial}_{k+1}=-M(x^{k+1}-x^{k})\to 0$ strongly in $X^{*}$. Moreover, Section 4.2 yields ~$\|\tilde{\partial}_{k+1}-x^{*}_{k+1}\|_{X^{*}}\to 0$ for some $x^{*}_{k+1}\in H(x^{k+1})$. Consequently $x^{*}_{k+1}\to 0$. Since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$, as in the proof of Lemma 4.24, we show the quasi-Féjer monotonicity (48) for all $\hat{x}\in\hat{X}$ and $k\in\mathbb{N}$.
Suppose then that $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$ for a subsequence $\{k_{j}\}_{j\in\mathbb{N}}\subset\mathbb{N}$ and a $\hat{x}\in X$. We want to show that $\hat{x}\in\hat{X}$. We consider two cases:
  1. 1.
    If $F$ is convex, $H$ is maximally monotone [^4] That the additive skew-adjoint term $\Xi$ does not destroy maximal monotonicity, can be proved completely analogously to the Hilbert space case in [7, Lemma 9.9]., hence weak-to-strong outer semicontinuous [7, Lemma 6.10]. Now $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$ and $H(x^{k_{j}+1})\ni x^{*}_{k_{j}+1}\to 0$ obliges $0\in H(\hat{x})$.
  2. 2.
    Suppose then that $F^{\prime}$ is weak-to-strong continuous. Now still $P:x\mapsto\partial G(x)+\Xi x$ is maximally monotone4, hence weak-to-strong outer semicontinuous. We have $P(x^{k_{j}+1})\ni x^{*}_{k_{j}+1}-F^{\prime}(x^{k_{j}+1})\to-F^{\prime}(\hat{x})$ strongly in $X^{*}$, as well as $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$, so we must have $-F^{\prime}(\hat{x})\in P(\hat{x})$. But this again says $0\in H(\hat{x})$.
Thus every weak limiting point $\hat{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$ satisfies $0\in H(\hat{x})$. But, since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, also $\hat{x}\in\mathbb{O}_{M}(\bar{x},\delta)$. This proves that $\hat{x}\in\hat{X}$. Since, by assumption, $\sum_{k=0}^{\infty}\varepsilon_{k}(\hat{x})<\infty$ for all $\hat{x}\in\hat{X}$, the quasi-Féjer monotonicity (48) with the quasi-Opial’s Lemma A.2 finishes the proof.
Example 4.36.
In the setting of Section 2 and Theorem 2.16, the weak-*-to-strong continuity of $F^{\prime}$ can be achieved, for example, when $F(x)=\frac{1}{2}\|S(x)-b\|^{2}$ for a Lipschitz and bounded $S$ with finite-dimensional range.
Remark 4.37.
All of our theory also applies when $X$ is the dual space of a separable normed space $X_{*}$, and we replace in our definitions $X^{*}$ by the predual space $X_{*}$, that is, subdifferentials are subsets of $X_{*}$, and $M,\Lambda\in\mathbb{L}(X;X_{*})$, etc. With this change the theory applies, for example, to $X$ a space of Radon measures, as in [34]. Then Theorem 4.34 proves the weak-* convergence.
Appendix A Opial’s lemma for quasi-Féjer monotonicity
Here we prove a generalisation of Opial’s lemma [23] for quasi-Féjer monotonicity, i.e, Féjer monotonicity with an additive error term. We prove it in normed spaces for Bregman divergences (47), as they add no extra difficulties. In an even more general variable-metric framework, a similar result is also proved in [21, Proposition 2.7]. Our simplified proof follows the outline of that in [7], and is nearly identical to the one in [34], where the errors took a more specific form.
For the proof, we recall the following deterministic version of the results of [25]:
Lemma A.1.
Let $\{a_{k}\}_{k\in\mathbb{N}}$, $\{b_{k}\}_{k\in\mathbb{N}}$, $\{c_{k}\}_{k\in\mathbb{N}}$, and $\{d_{k}\}_{k\in\mathbb{N}}$ be non-negative and $a_{k+1}\leq a_{k}(1+b_{k})+c_{k}-d_{k}$ for all $k\in\mathbb{N}$. If $\sum_{k=0}^{\infty}b_{k}<\infty$ and $\sum_{k=0}^{\infty}c_{k}<\infty$, then (i) $\lim_{k\to\infty}a_{k}$exists and is finite; and (ii) $\sum_{k=0}^{\infty}d_{k}<\infty$.
Lemma A.2.
Let either $X$ be the dual space of a corresponding separable normed space $X_{*}$, or, alternatively, let $X$ be reflexive. Also let $M:X\to\mathbb{R}$ be convex, proper, and Gâteaux differentiable with $M^{\prime}:X\to X_{*}$ weak-*-to-weak continuous. Finally, let $\hat{X}\subset X$ be non-empty and $\{e_{k}(\bar{x})\}_{k\in\mathbb{N}}\in\mathbb{R}$ for all $\bar{x}\in\hat{X}$. If
  1. (i)
    all weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$ belong $\hat{X}$;
  2. (ii)
    $B_{M}(x^{k+1},\bar{x})\leq B_{M}(x^{k},\bar{x})+e_{k}(\bar{x})$ for some $e_{k}(\bar{x})\geq 0$ for all $\bar{x}\in\hat{X}$ and $k\in\mathbb{N}$; and
  3. (iii)
    $\sum_{k=0}^{\infty}e_{k}(\bar{x})<\infty$ for all $\bar{x}\in\hat{X}$;
then all weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$ satisfy $\hat{x},\bar{x}\in\hat{X}$ and
(57) $\langle M^{\prime}(\hat{x})-M^{\prime}(\bar{x})|\hat{x}-\bar{x}\rangle=0.$
[…]
Proof A.3.
Let $\bar{x}$ and $\hat{x}$ be weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$. Since Bregman divergences $B_{M}\geq 0$ for convex $M$, the conditions Items (iii) and (ii) establish the assumptions of Lemma A.1 for $a_{k}=B_{M}(x^{k};\bar{x})$, $b_{k}=0$, $c_{k}=e_{k}(\bar{x})$, and $d_{k}=0$. It follows that $\{B_{M}(x^{k};\bar{x})\}_{k\in\mathbb{N}}$ is convergent. Likewise we establish that $\{B_{M}(x^{k};\hat{x})\}_{k\in\mathbb{N}}$ is convergent. Therefore, by the obvious three-point identity for Bregman divergences (see, e.g., [32]),
$\langle M^{\prime}(x^{k})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=B_{M}(x^{k};\hat{x})-B_{M}(x^{k};\bar{x})+B_{M}(\hat{x};\bar{x})\to c\in\mathbb{R}.$
Since $\bar{x}$ and $\hat{x}$ are a weak-* limit point, there exist subsequences $\{x^{k_{n}}\}_{n\in\mathbb{N}}$ and $\{x^{k_{m}}\}_{m\in\mathbb{N}}$ with $x^{k_{n}}\mathrel{\rightharpoonup}\bar{x}$ and $x^{k_{m}}\mathrel{\rightharpoonup}\hat{x}$. By the weak-*-to-weak continuity of $M^{\prime}:X\to X_{*}$, (57) follows from
$\langle M^{\prime}(\bar{x})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=\lim_{n\to\infty}\langle M^{\prime}(x^{k_{n}})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=c=\lim_{m\to\infty}\langle M^{\prime}(x^{k_{m}})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=0.$
If $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, and $X$ is the dual space of some separable normed space $X_{*}$, it contains a weakly-* convergent subsequence by the Banach–Alaoglu theorem, so a limit point exists as claimed. If $X$ is reflexive, the Eberlein–S̆mulyan theorem establishes the same result. Hence, if (57) implies $\bar{x}=\hat{x}$, then every convergent subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ has the same weak limit. It lies in $\hat{X}$ by Item (i). The final claim now follows from a standard subsequence–subsequence argument: Assume to the contrary that there exists a subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ not convergent to $\hat{x}$. Then the above argument provides a further subsequence converging to $\hat{x}$. This contradicts the fact that any subsequence of a convergent sequence converges to the same limit.