### v3

Abstract
With a view on bilevel and PDE-constrained optimisation, we develop iterative estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ for composite functions $F:=J\circ S$, where $S$ is the solution mapping of the inner optimisation problem or PDE, the latter seen in this work as a special case of the former.. The idea is to form a single-loop method by interweaving updates of the iterate $x^{k}$ by an outer optimisation method, with updates of the estimate by single steps of standard optimisation methods and linear system solvers. When the inner methods satisfy simple tracking inequalities, the differential estimates can almost directly be employed in standard convergence proofs for general forward-backward type methods. We adapt those proofs to a general inexact setting in normed spaces, that, besides our differential estimates, also covers mismatched adjoints and unreachable optimality conditions in measure spaces. As a side product of these efforts, we provide improved convergence results for nonconvex Primal-Dual Proximal Splitting (PDPS).
1 Introduction
First-order methods are slow. To be precise, they require a high number of iterations, but if those iterations are fast, they have the chance to practically overpower second-order methods with expensive iterations. In bilevel or PDE-constrained optimisation—the latter seen in this work as a special case of the former—the steps of basic first-order methods are very expensive, involving the solution of the inner problem or PDE and its adjoint. To make first-order methods fast, it is, therefore, imperative to reduce the cost of solving these subproblems—for instance, by employing inexact solution schemes.
Consequently, especially in the machine learning community, an interest has surfaced in single-loop methods for bilevel optimisation; see [29] and references therein. Many of these methods are very specific constructions. In [18] we started work on a more general approach to PDE-constrained optimisation: we showed that on each step of an outer primal-dual optimisation method, we can take single steps of standard linear system splitting schemes for the PDE constraint and its adjoint, and still obtain a convergent method that is computationally significantly faster than solving the PDEs exactly. The adjoint equation is used to compute differentials of the PDE or inner problem solutions with respect to its parameters. In [30] we then presented an approach to bilevel optimisation that allowed general inner and adjoint algorithms that satisfy certain tracking inequalities. These were proved for standard splitting schemes for the adjoint equation, and for forward-backward splitting and the Primal-Dual Proximal Splitting (PDPS) of [4] for the inner problem. The overall analysis was still tied to bilevel optimisation in Hilbert spaces, with forward-backward splitting as the outer optimisation method.
Our goal, in this work, is to develop a general theory of optimisation methods for bilevel and, by extension, multilevel problems. This theory is based on the idea of algorithmic single-loop differential estimation, with a flexible choice of outer, inner, and adjoint algorithms. The inner and adjoint methods only need to satisfy the aforementioned tracking inequalities. The outer method only has to tolerate controlled inexactness of differentials involving the inner solution mapping.
Problem setup
Write
$F=J\circ S_{u}$
for a mapping $S_{u}:X\to U$ and a differentiable function $J:U\to\mathbb{R}$, on normed spaces $X$ and $U$. Typically, but not necessarily, $S_{u}(x)$ is an inner solution mapping that arises from the satisfaction of
(1) $0=T(S_{u}(x),x)\quad\text{for a}\quad T:U\times X\to W_{*}\quad\text{with}\quad W_{*}\text{ a normed space}.$
The idea is that $T$ encodes the optimality conditions of an inner problem, parametrised by $x$, or a PDE, likewise parametrised by $x$. We are then interested in the solution of composite optimisation problems of the form
(2) $\min_{x\in X}\penalty 10000\ F(x)+G(x),$
or, more generally, the solution of optimality conditions
(3) $0\in F^{\prime}(x)+\partial G(x)+\Xi x,$
for $G$ convex but possibly nonsmooth, and $\Xi\in\mathbb{L}(X;X^{*})$ skew-adjoint. If $\Xi=0$, then this optimality condition is typically necessary for (2). More generally, the operator allows the modelling of primal-dual problems, and treating the PDPS and Douglas–Rachford splitting as generalised forward-backward splitting methods [7, 32]. We will discuss such formulations in more detail in Section˜4.
Contributions
Our contributions are as follows. In Sections˜2 and 3, which form our inner theory,
  1. (a)
    we show in general normed spaces that we can approximate in a single-loop fashion the differentials of compositions $F=J\circ S_{u}$, given abstract inner and adjoint algorithms for $S_{u}$, satisfying certain tracking inequalities. We then illustrate how standard algorithms satisfy these inequalities.
The corresponding results on sequences of scalars, on which these results are based on, are relegated to Appendix˜A. In contrast to [30] and, indeed, all single-loop bilevel optimisation methods that we are aware of, our approach can also work with the adjoint dimension reduction trick typically employed in PDE-constrained optimisation. We show that, subject to additive error terms with a bounded sum, the differential estimates $\widetilde{F^{\prime}}(x^{k})$ satisfy standard smoothness properties, such as Lipschitz differential and the two- and three-point descent inequalities [32, 7]. Based on this, in Section˜4, which forms our outer theory,
  1. 2.
    we prove various forms of convergence of general inexact splitting methods for (3).
In Section˜4.1 we also elaborate on the aforementioned connection of primal-dual methods to Eq.˜3. To facilitate the analysis of such methods, we introduce in Section˜5 operator-relative variants of the descent inequality. We then interpret the conditions of Section˜4 for outer forward-backward splitting and outer PDPS in Section˜6.
Not content to merely adapt existing proofs to inexact steps and normed spaces, we also present some improvements, especially for the nonconvex PDPS of [31]. We do, however, treat a slightly simplified problem. The original PDPS of [4] applies to $\min_{x}g(x)+h(K(x))$ with $K$ linear and $g$ and $h$ convex. The extension of [31] allows $K$ to be nonlinear. It is further analysed in [5, 6, 20, 12], with a simplified review of the different variants in Banach space in [33]. An alternative extension in [21] allows $g$ and $h$ to be semiconvex. We, instead, consider the special case $\min_{x}f(x)+g(x)+h(Kx)$ with $K$ linear, but $f$ nonconvex, and, in Section˜6,
  1. 3.
    we show that, for the nonconvex PDPS, the values of the convex envelope of the objective function at ergodic iterates locally converge to a minimum.
We do not include numerical results, as that has already been done in [18, 30]. Through our work, the specific algorithms presented therein can be understood through a clean and generic differential estimation approach.
Examples
The following examples illustrate the kind of algorithms that can be constructed, and whose convergence can be proved using our general theory. Our overall theory is, however, much more general than these two specific algorithms, as it allows the overall method to be composed of arbitrary inner, adjoint, and outer algorithms. Instead of gradients in Hilbert spaces, we will, moreover, work with differentials in general normed spaces.
Example 1.1 (Forward-backward–linear system splitting–forward-backward).
Consider the bilevel optimisation problem
$\min_{x,u}J(u)+G(x)\quad\text{subject to}\quad u\in\operatorname*{arg\,min}_{x}f(u;x)+g(u;x).$
Assume for simplicity that $f$ and $g$ are twice differentiable, convex in the first variable, and all the spaces are Hilbert. Then we can rewrite the problem as Eqs.˜2 and 1 by setting
$T(u,x)=\nabla_{u}f(u;x)+\nabla_{u}g(u;x).$
A standard forward backward method with step length parameter $\tau>0$ would update
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\nabla F(x^{k})),$
however the computation of $\nabla F(x)=\nabla S_{u}(x)^{*}\nabla J(S_{u}(x))$ is generally expensive. We therefore propose to form the iteratively updated single-loop differential estimate $\widetilde{\nabla F}(x^{k})$, as in the overall algorithm
(4) $\left\{\begin{aligned} u^{k+1}&:=\operatorname{prox}_{\sigma\nabla g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\sigma\nabla f(u^{k};x^{k})),&\text{(inner step)}\\ w^{k+1}&:=-N^{-1}(Mw^{k}+\nabla J(u^{k+1}))\quad\text{where}\quad N+M=\nabla_{u}T(u^{k+1},x^{k}),&\text{(adjoint step)}\\ \widetilde{\nabla F}(x^{k})&:=\nabla_{x}T(u^{k+1},x^{k})w^{k+1},&\text{(diff.\penalty 10000\ transform)}\\ x^{k+1}&:=\operatorname{prox}_{\tau G}(x^{k}-\widetilde{\nabla F}(x^{k})).&\text{(outer step)}\end{aligned}\right$.
The first line is a standard forward-backward step for the inner variable. The second line applies, e.g., Gauss–Seidel or block-Gauss–Seidel splitting (depending on the choice of the operators $N$ and $M$) to the linear equation
$\nabla_{u}T(u^{k+1},x^{k})w^{k+1}+\nabla J(u^{k+1})=0.$
Of course, the choice $M=0$ and $N=\nabla_{u}T(u^{k+1},x^{k})$ would solve this equation exactly. The third line of (4) transforms the inner and adjoint iterates $u^{k+1}$ and $w^{k+1}$ into the differential estimate $\widetilde{\nabla F}(x^{k})$. As we later elaborate in Example˜2.1, this construction is motivated by the reduced adjoint equation
$\nabla_{u}T(S_{u}(x),x)w_{x}+\nabla J(S_{u}(x))=0\quad\text{whose solution satisfies}\quad\nabla F(x)=\nabla_{x}T(S_{u}(x),x)w_{x}.$
We treat specific inner and adjoint methods and differential transformations in Section˜3 after the general theory in Section˜2. The fourth line of (4) is an outer inexact forward-backward step. We present a general theory of outer/inexact methods in Section˜4, and provide examples in Section˜6.
Example 1.2 (A simple PDE-constrained optimisation problem).
With $U\subset H^{1}(\Omega)$ and $X\subset L^{2}(\Omega)$, define for admissible $x\in X$ the linear operator $A_{x}\in\mathbb{L}(U;U^{*})$ by $A_{x}u:=\nabla^{*}[x\cdot\nabla u]$, and consider the PDE-constrained optimisation problem
$\min_{x,u}J(u)+G(x)\quad\text{subject to}\quad A_{x}u=b,$
where we use $G$ to restrict $x$ to a domain where the PDE is well-posed. Setting $T(u,x)=A_{x}u-b$, we are again the setup of Eqs.˜2 and 1. Observing that $\nabla_{u}T(u^{k+1},x^{k})=A_{x}$ and $\nabla_{x}T(u^{k+1},x^{k})w^{k+1}=\nabla u^{k+1}\cdot\nabla w^{k+1}$, similarly to Example˜1.1, but this time also applying a linear system splitting scheme to the PDE itself, we arrive at the algorithm
$\left\{\begin{aligned} \text{split }&N+M:=A_{x^{k}},&\text{(operator splitting)}\\ u^{k+1}&:=-N^{-1}(Mu^{k}-b),&\text{(PDE step)}\\ w^{k+1}&:=-N^{-1}(Mw^{k}+\nabla J(u^{k+1})),&\text{(adjoint step)}\\ \widetilde{\nabla F}(x^{k})&:=\nabla u^{k+1}\cdot\nabla w^{k+1},&\text{(diff.\penalty 10000\ transform)}\\ x^{k+1}&:=\operatorname{prox}_{\tau G}(x^{k}-\widetilde{\nabla F}(x^{k})).&\text{(outer step)}\end{aligned}\right$.
Note that, to use standard schemes such as Gauss–Seidel, we need to work in finite-dimensional subspaces. However, in infinite dimensions, the abstract splitting $A_{x}=N+M$ can applied on the block structure of $A_{x}$ in more complex problems.
Further related work
Through our general approach to inexactness in Section˜4, independent of the differential estimation theory of Section˜2, besides differential estimates for multilevel problems, we can model mismatched adjoints [19], and difficult-to-solve-exactly optimality conditions in measure spaces [35]. We also adopt the approach of [35] to optimisation in normed spaces: instead of Bregman divergences, we construct an inner product structure with a self-adjoint $M\in\mathbb{L}(X;X^{*})$. Our work is related to the study of gradient oracles for smooth convex optimisation in [9], and for nonconvex composite optimisation in [11, 22], both in finite-dimensional Euclidean spaces. Based on sufficient descent and the Kurdyka–Łojasiewicz property, [24] also study inexact methods in $\mathbb{R}^{n}$. Moreover, [3] introduce approaches to control model inexactness in proximal trust region methods, and [28] in non-single-loop gradient methods for bilevel optimisation.
Notation and basic concepts
We write $\mathbb{L}(X;Y)$ for the space of bounded linear operators between the normed spaces $X$ and $Y$, and $\operatorname{Id}$ for the identity operator. $X^{*}$ stands for the dual space of $X$. When $X$ is Hilbert, we identify $X^{*}$ with $X$. We write $\langle x,y\rangle$ for an inner product, $\langle x^{*}|x\rangle_{X^{*},X}$ for a dual product. We call $M\in\mathbb{L}(X;X^{*})$ self-adjoint if the restriction $M^{*}|X=M$, and positive semi-definite if $\langle x|Mx\rangle_{X,X^{*}}\geq 0$ for all $x\in X$. If both hold, we set $\|x\|_{M}:=\sqrt{\langle Mx|x\rangle}$, and write $\mathbb{O}_{M}(x,r)$ for the radius-$r$ open ball at $x$ in the MM-(semi-)norm. The open ball in the standard norm of the space is denoted by $\mathbb{O}(x,r)$. We also write $M\geq N$ if MNM-N is positive semi-definite. We extensively use the vectorial Young’s inequality
$\langle x^{*}|x\rangle_{X^{*},X}\leq\frac{a}{2}\|x\|_{X}^{2}+\frac{1}{2a}\|x^{*}\|_{X^{*}}^{2}\qquad\text{for all }x\in X,\,x^{*}\in X^{*}\,a>0.$
For sets $A$ and $B$, we often write $\langle A-B|x\rangle\geq 0$, which means that $\langle y-z|x\rangle\geq 0$ for all $y\in A$ and $z\in B$.
For $F:X\to\mathbb{R}$, we write $DF(x)$ for the Gâteaux and $F^{\prime}(x)\in X^{*}$ for the Fréchet derivative at $x$, if they exist. If $X$ is Hilbert, $\nabla F(x)\in X$ stands for the Riesz representation of $F^{\prime}(x)$, i.e., the gradient. For partial derivatives, we use the notation $F^{(x)}(u,x)$. We also write $\operatorname{sub}_{c}F:=\{x\in X\mid F(x)\leq c\}$ for the $c$-sublevel set. With $\overline{\mathbb{R}}:=[-\infty,\infty]$, for a convex $G:X\to\overline{\mathbb{R}}$, we write $\operatorname{dom}G$ for the effective domain, $\partial G(x)$ for the subdifferential at $x$, and $G^{*}:X^{*}\to\overline{\mathbb{R}}$ for the Fenchel conjugate. When $X$ is a Hilbert space, we write $\operatorname{prox}_{F}$ for the proximal map and, with a slight abuse of notation, identify $\partial G(x)$ with the set of Riesz representations of its elements.
2 Differential estimation from tracking inequalities
Let $J:U\to\mathbb{R}$ and $S_{u}:X\to U$ be Fréchet differentiable on normed spaces $X$ and $U$. We consider
$F(x)=J(S_{u}(x)).$
Typically, but not necessarily in the overall theory, $S_{u}(x)$ arises from the satisfaction of (1).
As $S_{u}$ and its differential can be expensive to compute, given an iterate $x^{k}$ of an arbitrary outer algorithm for minimising an objective that involves $F$, such as (2), we estimate $S_{u}(x^{k})$ by an inner iterate $u^{k+1}\in U$, and $S_{u}^{\prime}(x^{k})$ by an adjoint iterate $p^{k+1}\in\mathbb{L}(X;U)$, that is, we estimate
$F^{\prime}(x^{k})=J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\quad\text{by}\quad\widetilde{F^{\prime}}(x^{k})=J^{\prime}(u^{k+1})p^{k+1}.$
When $X$ is a Hilbert space, we write $\widetilde{\nabla F}(x^{k})$ for the Riesz representation of $\widetilde{F^{\prime}}(x^{k})$. We do not provide a single explicit formula for $u^{k+1}$ and $p^{k+1}$. Instead, we assume them to satisfy tracking estimates as in [18, 30]. We formulate these tracking estimates—that are essentially contractivity estimates with suitable penalties for parameter change—in Section˜2.1. Our goal is to derive, in Section˜2.2, variants of standard descent inequalities and Lipschitz bounds for the estimate $\widetilde{F^{\prime}}(x^{k})$. In Section˜3 we provide examples of inner and adjoint methods that satisfy the tracking inequalities.
Although $\widetilde{F^{\prime}}(x^{k})$ will have the above structure, we want to avoid constructing $p^{k+1}\approx S_{u}^{\prime}(x^{k})\in\mathbb{L}(X;U)$ directly due to its high dimensionality. Instead, we seek to only construct the necessary projections through a lower-dimensional variable $w^{k+1}$. We illustrate this idea in the following example.
Example 2.1 (Adjoint equations).
Suppose $S_{u}(x)$ arises from the satisfaction of (1). By implicit differentiation, subject to sufficient differentiability and (1) holding in a neighbourhood of $x$, we obtain the basic adjoint
(5) $T^{(u)}(S_{u}(x),x)S_{u}^{\prime}(x)+T^{(x)}(S_{u}(x),x)=0\in\mathbb{L}(X;W),$
where $S_{u}^{\prime}(x)\in\mathbb{L}(X;U)$, $T^{(u)}(S_{u}(x),x)\in\mathbb{L}(U;W)$, and $T^{(x)}(S_{u}(x),x)\in\mathbb{L}(X;W)$. Hence, following the derivation of adjoint PDEs in, e.g., [16, §1.6.2] or [8, §1.2], assuming $T^{(u)}(S_{u}(x),x)$ to be invertible, we solve from (5) that
$[J\circ S_{u}]^{\prime}(x)=J^{\prime}(S_{u}(x))S_{u}^{\prime}(x)=S_{w}(x)T^{(x)}(S_{u}(x),x),$
for a $S_{w}(x)\in W$ satisfying the reduced adjoint
(6) $S_{w}(x)T^{(u)}(S_{u}(x),x)+J^{\prime}(S_{u}(x))=0.$
For $x=x^{k}$, we will in practise take $w^{k+1}$ as an operator splitting approximation to
(7) $w^{k+1}T^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0,$
and then set
$\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})\approx J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k}).$
2.1 The tracking assumption
To track the inexact computations of inner and adjoint variables across iterations, we introduce verifiable conditions that quantify how closely the computed values follow the outputs of the exact inner and adjoint solution mappings evaluated at the current outer iterate. These tracking assumptions ensure that the accumulated errors remain controlled and that the approximate gradient remains meaningful for descent. The following assumption formalises this idea.
Assumption 2.1.
Let the normed spaces $X$ and $X^{*}$ and abstract spaces $U$, and $W$, be equipped with the distance functions $b_{X}:X\times X\to[0,\infty]$, $d_{X^{*}}:X^{*}\times X^{*}\to[0,\infty]$, $d_{U}:U\times U\to[0,\infty]$, $d_{W}:W\times W\to[0,\infty]$. Then for a subset $\Omega\subset X$; an inner solution map $S_{u}:X\to U$ and an adjoint solution map $S_{w}:X\to W$; the following hold:
  1. (i)
    We are given an inner algorithm that, on each iteration $k\geq 1$, given $\{x^{n}\}_{n=0}^{k}\subset\Omega$ and $u^{k}\in U$, produces $u^{k+1}\in U$ satisfying for some $\pi_{u}>0$, $\kappa_{u}>1$ the inner trackign inequality
    $\kappa_{u}{d_{U}(u^{k+1},S_{u}(x^{k}))}\leq{d_{U}(u^{k},S_{u}(x^{k-1}))}+\pi_{u}{b_{X}(x^{k},x^{k-1})}.$
  2. (ii)
    We are given an adjoint algorithm that, on each iteration $k\geq 1$, given $\{(x^{n},u^{n+1})\}_{n=0}^{k}\subset\Omega\times U$, and $w^{k}\in W$, produces $w^{k+1}\in W$ satisfying for some ${\mu_{u}},\pi_{w}>0$, $\kappa_{w}>1$ the adjoint tracking inequality
    $\kappa_{w}{d_{W}(w^{k+1},S_{w}(x^{k}))}\leq{d_{W}(w^{k},S_{w}(x^{k-1}))}+{\mu_{u}}{d_{U}(u^{k+1},S_{u}(x^{k}))}+\pi_{w}{b_{X}(x^{k},x^{k-1})}.$
  3. (iii)
    We are given a differential transformation that, on each iteration $k\in\mathbb{N}$, given $u^{k+1}\in U$ and $w^{k+1}\in W$, produces $\widetilde{F^{\prime}}(x^{k})\in X^{*}$ that satisfies for some $\alpha_{u},\alpha_{w}\geq 0$ the bound
    ${d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq\alpha_{u}{d_{U}(u^{k+1},S_{u}(x^{k}))}+\alpha_{w}{d_{W}(w^{k+1},S_{w}(x^{k}))}.$
Remark 2.2 (Meaning of the conditions).
The inner and adjoint tracking conditions Items˜(i) and (ii) are parameter change aware contractivity conditions for the inner and adjoint algorithms: if $x^{k}=x^{k-1}$, the former reduces to a standard contractivity condition. For examples of such algorithms, recall Examples˜1.1 and 1.2. The condition Item˜(iii) allows converting the construction error of $\widetilde{F^{\prime}}(x^{k})$ to the tracking errors of the inner and adjoint algorithms. We provide more detailed examples of algorithms that satisfy these conditions in Section˜3.
Remark 2.3 (Distance functions).
The distance functions $d_{U}:U\times U\to\overline{\mathbb{R}}$, etc., will in the simplest cases be given by norms or semi-norms, but we do not require this, especially from $d_{U}$ and $d_{W}$. Often in Section˜6, even norms will be adapted to the algorithms under study: they will not be the standard norms of the spaces. Only $d_{X^{*}}$ will later on be required to observe a Young relationship. Even then, it will in some of our examples in Section˜6 arise from the conjugation of a semi-norm, and hence have infinite values outside a subspace. We write squared distances as $d_{U}^{2}(u,\tilde{u}):=d_{U}(u,\tilde{u})^{2}$.
Remark 2.4 (Initial iterates and first-step errors).
The initial iterates are $x^{0}$, $u^{0}$, and $w^{0}$. For the inner and adjoint variables, the initial iterates do not appear in the tracking inequalities, which start from $x^{0}$, $u^{1}$, and $w^{1}$: the idea is that each $u^{k+1}$, which is computed using $x^{k}$, is compared against the exact inner solution $S_{u}(x^{k})$, and likewise for the adjoint variable. The first-step errors $d_{U}(u^{1},S_{u}(x^{0}))$} and $d_{W}(w^{1},S_{w}(x^{0}))$}, distinct from the initialisation errors ${d_{U}}(u^{0},S_{u}(x^{0}))$ and ${d_{W}}(w^{0},S_{w}(x^{0}))$, will appear in our final estimates. They can be made small by solving the first step to a high precision. Contractive algorithms guarantee them to be smaller than the initialisation errors.
2.2 Smoothness of differential estimates
We now derive descent- and Lipschitz-type inequalities for the approximate differential $\widetilde{F^{\prime}}(x^{k})$}, extending these classical smoothness concepts to account for differential errors under the tracking framework. Most of these result are straightforward consequences of Section˜2.1 and the scalar recurrence results of Appendix˜A.
We recall that if $F^{\prime}$ is $L$-Lipschitz, it then satisfies the descent inequality [7, Theorem 7.1]
(8) $\langle F^{\prime}(x^{k})|x-x^{k}\rangle_{X^{*},X}\geq F(x)-F(x^{k})-\frac{L}{2}{d_{X}^{2}(x,x^{k})}\quad\text{for all}\quad x,x^{k}\in X$
for the choice ${d_{X}^{2}(x,x^{k})}=\|x-x^{k}\|_{X}$. We allow here any distance expression ${d_{X}}:X\times X\to\overline{\mathbb{R}}$ such that $d_{X^{*}}$} is Young to $d_{X}$}. By this we mean that both expressions are one-homogeneous and
$\langle\tilde{x}^{*}-x^{*}|\tilde{x}-x\rangle_{X^{*},X}\leq\frac{1}{2}{d_{X^{*}}^{2}(\tilde{x}^{*},x^{*})}+\frac{1}{2}{d_{X}^{2}(x^{*},x)}\quad\text{for all}\quad x,\tilde{x}\in X\text{ and }x^{*},\tilde{x}^{*}\in X^{*}.$
In particular, we allow $d_{X}$} and $d_{X^{*}}$} to be given by (semi-)norms generated by operators. We treat the descent inequality with respect to them in Section˜5.
If ${b_{X}}={d_{X}}$, simply taking $\bar{x}=x^{k}$ and $\tilde{\gamma}=\varsigma_{p}$ in the theorem to follow, and combining with the descent inequality (8), we obtain the inexact descent inequality at $x=x^{k+1}$,
(9) $\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{L+2\varsigma_{p}}{2}{d_{X}^{2}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}.$
According to the theorem, the final error term has a bounded sum, which we will be able to manage in convergence proofs of optimisation methods (in Section˜4).
In such convergence proofs, it is frequently convenient to directly use the three-point descent inequality (see [7, Corollary 7.2] for the convex case, or [32, Appendix B] for the non-convex case)
(10) $\langle F^{\prime}(x^{k})|x-\bar{x}\rangle_{X^{*},X}\geq F(x)-F(\bar{x})+\frac{\beta}{2}{d_{X}^{2}(x,\bar{x})}-\frac{L}{2}{d_{X}^{2}(x,x^{k})}\quad\text{for all}\quad x,x^{k},\bar{x}\in X.$
If, again, ${b_{X}}={d_{X}}$, combining the next theorem with this inequality, we obtain the inexact version at $x=x^{k+1}$,
(11) $\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq F(x^{k+1})-F(\bar{x})+\frac{\beta-\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}-\frac{\varsigma_{p}^{2}\tilde{\gamma}^{-1}+L}{2}{d_{X}^{2}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}.$
We write
(12) $\kappa:=\min(\kappa_{u},\kappa_{w})\quad\text{and}\quad\overline{\kappa}:=\max(\kappa_{u},\kappa_{w}).$
Theorem 2.5.
Suppose Section˜2.1 holds, $X$ is a normed space, $X^{*}$ its dual with $d_{X^{*}}$} Young to $d_{X}$}. Also let $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\geq 1$, and pick $p\in[1,\kappa)$. Then, for any $\tilde{\gamma}>0$ and $\bar{x}\in X$, we have
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq-\frac{\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}{b_{X}(x^{k+1},x^{k})}-\frac{1}{2\tilde{\gamma}}e_{p,k}$
for some $\varsigma_{p}\geq 0$ and $e_{p,k}\in\mathbb{R}$ that satisfies $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}<\infty$.
Proof 2.6.
By the Young relationship of $d_{X}$} and $d_{X^{*}}$},
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq-\frac{1}{2\tilde{\gamma}}{d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}-\frac{\tilde{\gamma}}{2}{d_{X}^{2}(x^{k+1},\bar{x})}.$
Let ${\varrho_{k}}:={b_{X}(x^{k},x^{k-1})}$, ${d^{u}_{k}}:={d_{U}(u^{k},S_{u}(x^{k-1}))}$, ${d^{w}_{k}}:={d_{W}(w^{k},S_{w}(x^{k-1}))}$, and ${\widetilde{d}_{k}}={d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}$. Then Section˜2.1 and $\{x^{n}\}_{n=0}^{k}\subset\Omega$ imply Appendix˜A for the index $k$. Now an application of Lemma˜A.5 establishes
${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq\varsigma_{p}^{2}{b_{X}(x^{k+1},x^{k})}+e_{p,k},$
where $\varsigma_{p}$ and $e_{p,k}$ defined in Eqs.˜67 and 72 satisfy the claimed bounds by the lemma. Combining these two inequalities establishes our claim.
Remark 2.7.
From Eqs.˜67 and 72 we can obtain the more precise bounds
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}\leq\frac{{d_{U}^{2}(u^{1},S_{u}(x^{0}))}}{\pi_{u}}\bigg(\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg)+\frac{{d_{W}^{2}(w^{1},S_{w}(x^{0}))}}{\pi_{w}}\bigg(\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg)$
and
(13) $\varsigma_{p}\leq\frac{(\alpha_{u}\pi_{u}+\alpha_{w}\pi_{w})\kappa\overline{\kappa}}{p(\kappa-p)}+\frac{\alpha_{w}{\mu_{u}}\pi_{u}\overline{\kappa}}{p^{2}(\kappa-p)^{2}}.$
The sum of the error terms $e_{p,k}$ can, therefore, be made arbitrarily small by taking good-quality first inner and adjoint steps.
It will also be useful to have a Lipschitz-type property. If $d_{X^{*}}$ satisfies the triangle inequality, combining the next theorem with $F^{\prime}$ being $L$-Lipschitz with respect to the distances $d_{X^{*}}$} and $d_{X}$}, we can get the Lipschitz property with error for $\widetilde{F^{\prime}}$,
${d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x))}\leq L{d_{X}(x,x^{k})}+\sqrt{e_{\mathrm{lip},k}}\qquad(x\in X).$
According to the theorem, the error terms again have bounded sum, if the squares of residual terms $b_{X}(x^{k+1},x^{k})$} from the tracking inequalities have a bounded sum. Many standard convergence proofs, include the ones of Section˜4 will generally ensure the latter.
Theorem 2.8.
Suppose Section˜2.1 holds, and that $\{x^{n}\}_{n=0}^{k}\subset\Omega$ for a $k\geq 1$. Then
(14) ${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq e_{\mathrm{lip},k},$
where $e_{\mathrm{lip},k}\geq 0$ is such that
(15) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty\quad\text{whenever}\quad\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{b_{X}^{2}(x^{k+1},x^{k})}<\infty.$
Proof 2.9.
Let ${\varrho_{k}}:={b_{X}(x^{k},x^{k-1})}$, ${d^{u}_{k}}:={d_{U}(u^{k},S_{u}(x^{k-1}))}$, ${d^{w}_{k}}:={d_{W}(w^{k},S_{w}(x^{k-1}))}$, as well as ${\widetilde{d}_{k}}={d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}$. Then Section˜2.1 and $\{x^{n}\}_{n=0}^{k}\subset\Omega$ imply Appendix˜A for the index $k$. Now Lemma˜A.3 with $p=1$ gives (14) for $e_{\mathrm{lip},k}:=\breve{e}_{1,k}$ as defined in (68). The limiting property (15) is a consequence of Lemma˜A.7.
3 Inner and adjoint algorithms
We next provide brief examples of inner and adjoint methods that satisfy the corresponding parts of Section˜2.1; hence iterative ways to construct differential estimates $\widetilde{F^{\prime}}(x^{k})$ that satisfy the inexact smoothness-type properties of Section˜2.2. This section is largely based on [30], with some streamlined proofs, and some omitted proofs. The framing of the proofs in our differential estimation framework, is obviously new.
3.1 Inner algorithms and the inner tracking inequality
Forward-backward splitting
On a Hilbert space $U$ and a normed space $X$, consider the parametric inner problem
$S_{u}(x)=\operatorname*{arg\,min}_{u}f(u;x)+g(u;x)$
for $f$ and $g$ convex in $u$, and differentiable in $(u,x)$. This is an instance of (1) with
$T(u,x)=\nabla f(u;x)+\nabla g(u;x).$
Theorem 3.1.
Suppose $\nabla f(\,\boldsymbol{\cdot}\,;x)$ $L$-Lipschitz, and $g(\,\boldsymbol{\cdot}\,;x)$ $\gamma$-strongly convex, both uniformly in $x\in\Omega$ for an $\Omega\subset X$ If $S_{u}$, as defined above, is Lipschitz in $\Omega$, and $\tau L\leq 1$ for a step length parameter $\tau>0$, then Section˜2.1Item˜(i) holds with $\kappa_{u}=1+\tau\gamma$ and $\pi_{u}=L_{S_{u}}$ for the inner forward-backward splitting updates
$u^{k+1}:=\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u^{k}-\tau\nabla f(u^{k};x^{k})).$
Proof 3.2.
This is essentially [30, Theorem 3.4], but we replicate the proof here as an example. With $k\geq 1$, let $x^{k},x^{k-1}\in\Omega$. Since $\tau g(\,\boldsymbol{\cdot}\,;x^{k})$ is $\tau\gamma$-strongly convex, its proximal map is $1/(1+\tau\gamma)$-contractive [34, Lemma 4.4], satisfying for all $u,\tilde{u}\in U$ that
$(1+\tau\gamma)\|\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(u)-\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(\tilde{u})\|\leq\|u-\tilde{u}\|$.
Likewise the gradient descent map $V(u):=u-\tau\nabla f(u;x^{k})$ is non-expansive when $\tau L\leq 1$, i.e., $\|V(u)-V(\tilde{u})\|\leq\|u-\tilde{u}\|$. Since $\operatorname{prox}_{\tau g(\,\boldsymbol{\cdot}\,;x^{k})}(V(S_{u}(x^{k})))=S_{u}(x^{k})$ [7, Theorem 4.2 & Corollary 6.22], combining these inequalities yields
$\begin{split}(1+\tau\gamma)\|u^{k+1}-S_{u}(x^{k})\|&\leq\|V(u^{k})-V(S_{u}(x^{k}))\|\leq\|u^{k}-S_{u}(x^{k})\|\\ &\leq\|u^{k}-S_{u}(x^{k-1})\|+\|S_{u}(x^{k})-S_{u}(x^{k-1})\|\\ &\leq\|u^{k}-S_{u}(x^{k-1})\|+L_{S_{u}}\|x^{k}-x^{k-1}\|.\end{split}$
Remark 3.3 (Lipschitz solution mapping).
The Lipschitz assumption on $S_{u}$ is guaranteed in sufficiently smooth cases by the classical implicit function theorem applied to the equation $T(u,x)=0$; see [30, Appendix B]. Nonsmooth implicit function theorems and the Aubin or pseudo-Lipschitz property of the set-valued mapping $S_{u}$ are studied in, e.g., [10, 17] as well as [7, Theorem 28.3]. If $S_{u}$ has the Aubin property, it will be Lipschitz if we assume, e.g., strict convexity to ensure the uniqueness of solutions. For the specific case $f(u;x)=\bar{f}(u)$ and $g(u;x)=x\bar{g}(u)$ with a scalar $x$, we refer to [7, Theorem 28.5].
Primal-dual proximal splitting
On a Hilbert space $Z$ and a normed space $X$, consider the inner problem
$\min_{z}f(z;x)+g^{*}(Kz;x).$
for $K\in\mathbb{L}(Z;Y^{*})$ linear and bounded to a Hilbert space $Y^{*}$, both $f$ and $g$ convex in the first parameter, and differentiable in both parameters. As an instance of (1), we represent the Fenchel–Rockafellar primal-dual optimality conditions of this problem as a root $u$ of the mapping
$T(u,x)=(\nabla f(z;x)+K^{*}y,\nabla g(y;x)-Kz)\quad\text{where}\quad u=(z,y)\in U=Z\times Y.$
Theorem 3.4.
Suppose $g(\,\boldsymbol{\cdot}\,;x)$ is $\gamma$-strongly convex uniformly in $x\in\Omega$ for a $\Omega\subset X$. If $S_{u}(x)=T^{-1}(\,\boldsymbol{\cdot}\,;x)(0)$ is Lipschitz in $\Omega$, and the step length parameters $\tau,\sigma>0$ satisfy $\tau\sigma\|K\|\leq 1$, then Section˜2.1Item˜(i) holds for the inner PDPS updates
$z^{k+1}=\operatorname{prox}_{\tau f(\,\boldsymbol{\cdot}\,;x^{k})}(z^{k}-\tau K^{*}y^{k})\quad\text{and}\quad y^{k+1}=\operatorname{prox}_{\sigma g(\,\boldsymbol{\cdot}\,;x^{k})}(y^{k}+\sigma K(2z^{k+1}-z^{k})).$
The proof in [30, Theorem 3.6] is fundamentally similar to the forward-backward splitting in Theorem˜3.1, but requires working with operator-induced norms and monotone operators.
Linear system splitting
We now cover algorithms for PDEs as the inner problems. For $U$ a Euclidean space and $X$ a normed space, let both $A_{x}\in\mathbb{L}(U;U)$, modelling a linear PDE parametrised $x$, and the right hand side $b_{x}\in U$ be Lipschitz in $x\in X$. Consider the inner constraint of $u=S_{u}(x)$ satisfying
(16) $A_{x}u=b_{x}.$
This is again an instance of (1) when we set
$T(u,x)=A_{x}u-b_{x}.$
Two solve (16) inexactly and efficiently, we split $A_{x}$ into two components, as in standard Gauss–Seidel or Jacobi splittings.
Assumption 3.4 (Admissible splitting).
Let $X$ be a normed space and $U$ a Hilbert space, and $A_{x}\in\mathbb{L}(U;U)$ for all $x$ in a set $\Omega\subset X$. We assume to be given splittings $A_{x}=N_{x}+M_{x}$ with $N_{x}$ invertible, and such that $\|N^{-1}_{x}M_{x}\|\leq\zeta$ for some $\zeta\in[0,1)$.
Theorem 3.5.
Suppose Section˜3.1 holds and that $S_{u}(x)=A^{-1}_{x}b_{x}$ is $L_{S_{u}}$-Lipschitz in $\Omega$. Then the updates $u^{k+1}=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})$ satisfy Section˜2.1Item˜(i) in $\Omega$ with $\kappa_{u}=\zeta^{-1}$ and $\pi_{u}=L_{S_{u}}$.
Proof 3.6.
We have
$u^{k+1}-S_{u}(x^{k})=N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}u^{k})-N^{-1}_{x^{k}}(b_{x^{k}}-M_{x^{k}}S_{u}(x^{k}))=-N^{-1}_{x^{k}}M_{x^{k}}(u^{k}-S_{u}(x^{k})).$
Hence
$\zeta^{-1}\|u^{k+1}-S_{u}(x^{k})\|\leq\|u^{k}-S_{u}(x^{k})\|\leq\|u^{k}-S_{u}(x^{k-1})\|+L_{S_{u}}\|x^{k+1}-x^{k}\|.$
To provide examples of Section˜3.1, we extend the condition to splittings that are admissible for the adjoint equations that we treat next.
Assumption 3.6 (Adjoint admissible splitting).
Let $X$ be a normed space and $U$ a Hilbert space with $A_{x}\in\mathbb{L}(U;U)$ for all $x$ in a set $\Omega\subset X$. We assume to be given splittings $A_{x}=N_{x}+M_{x}$ with $N_{x}$ invertible, and such that
(17) $\|N^{-1}_{x}M_{x}\|\leq\zeta\quad\text{and}\quad\gamma_{N}\|N^{-1}_{x}\|\leq 1$
for some $\zeta\in[0,1)$ and $\gamma_{N}>0$.
Example 3.7 (Jacobi splitting [30, Example 4.3]).
Let $A_{x}\in\mathbb{R}^{n\times n}$, and take $N_{x}$ as the diagonal of $A_{x}$. Regarding Section˜3.1, the first part of (17) reduces to strict diagonal dominance, see [13, §10.1]. The second part always holds and $N_{x}$ is invertible when the main diagonal of $A_{x}$ has only positive entries. Then $\gamma_{N}$ is the minimum of the diagonal values.
Example 3.8 (Gauss–Seidel splitting [30, Example 4.4]).
Let $A_{x}\in\mathbb{R}^{n\times n}$, and take $N_{x}$ as the upper triangle and diagonal of $A_{x}$. Regarding Section˜3.1, the first inequality of (17) holds for some $\zeta\in[0,1)$ when $A_{x}$ is symmetric and positive definite. This follows similarly to [13, proof of Theorem 10.1.2]. The second inequality holds for some $\gamma_{N}$ when $N_{x}$ is invertible, i.e., has no zeros on the main diagonal.
3.2 Adjoint algorithms and the adjoint tracking inequality; the differential transformation condition
We now treat adjoint methods and the differential transformation, i.e., Appendix˜AItems˜(i) and ii. Here, $S_{u}$ arises from (1) for a corresponding $T$, so the next two approaches to the adjoint equation and differential transformation can be combined with each of inner algorithms of the previous subsection.
Basic adjoint: linear system splitting
With $U$, $W_{*}$, and $X$ normed spaces, let $S_{u}$ and $T$ be given by (1), and define
$\widetilde{F^{\prime}}(x^{k}):=J^{\prime}(u^{k+1})p^{k+1},$
for $p^{k+1}\in\mathbb{L}(X;U)$ computed by taking (single or multiple) admissible splitting steps (see Sections˜3.1, 3.8 and 3.7) on the linear equation
(18) $0=T^{(u)}(u^{k+1},x^{k})p+T^{(x)}(u^{k+1},x^{k})$
with unknown $p=p^{k+1}$. Correspondingly, let $S_{w}=S_{p}=S_{u}^{\prime}$ arise from the basic adjoint (5). The next theorem shows that the scheme satisfies Section˜2.1 in case of single steps; multiple steps follow immediately with larger $\kappa_{w}$ by chaining Section˜2.1Item˜(ii).
Theorem 3.9.
Suppose that $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable for an $\Omega\subset X$, and that the adjoint admissible splitting Section˜3.1 holds in $U\times\Omega$ for the operators $A_{(u,x)}=T^{(u)}(u,x)$. Suppose, moreover, that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ and $T^{(u)}(\,\boldsymbol{\cdot}\,,x)$ are, respectively, $L_{T^{(u)};u}$- and $L_{T^{(x)};u}$-Lipschitz for all $x\in\Omega$; that $S_{u}^{\prime}$ is $L_{S_{u}^{\prime}}$-Lipschitz in $\Omega$; and that
$N_{S_{u}^{\prime}}:=\sup\{\|S_{u}^{\prime}(x)\|_{\mathbb{L}(X;U)}\mid x\in\Omega\}<\infty.$
Let
$p^{k+1}:=N^{-1}_{(u^{k+1},x^{k})}(T^{(x)}(u^{k+1},x^{k})-M_{(u^{k+1},x^{k})}p^{k}).$
Then Section˜2.1Item˜(ii) holds in $\Omega$ for $w^{k}=p^{k}$ with $\kappa_{w}=\zeta^{-1}$, $\mu_{u}=\zeta^{-1}\gamma^{-1}_{N}(L_{T^{(u)};u}N_{S_{u}^{\prime}}+L_{T^{(x)};u})$, and $\pi_{w}=L_{S_{u}^{\prime}}$.
Suppose further that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz and
$N_{J^{\prime}}:=\sup\{\|J^{\prime}(S_{u}(x))\|_{U^{*}}\mid x\in\Omega\}<\infty.$
Then Section˜2.1Item˜(iii) holds for $\widetilde{F^{\prime}}(x^{k}):=J^{\prime}(u^{k+1})p^{k+1}$ with $\alpha_{u}=L_{J^{\prime}}N_{S_{u}^{\prime}}$ and $\alpha_{w}=N_{J^{\prime}}$.
Proof 3.10.
Section˜2.1Item˜(ii) is proved in [30, Theorem 4.9]. That theorem is written entirely in Hilbert spaces, but the proof does not require the setting. Instead of reproducing the proof here, we provide the analogous proof for reduced adjoints in the next Theorem˜3.11. Section˜2.1Item˜(iii) follows from
$\begin{split}\|\widetilde{F^{\prime}}(x^{k})-&F^{\prime}(x^{k})\|_{X^{*}}=\|J^{\prime}(u^{k+1})p^{k+1}-J^{\prime}(S_{u}(x^{k}))S_{u}^{\prime}(x^{k})\|_{X^{*}}\\ &\leq\|J^{\prime}(u^{k+1})[p^{k+1}-S_{u}^{\prime}(x^{k})]\|_{X^{*}}+\|[J^{\prime}(u^{k+1})-J^{\prime}(S_{u}(x^{k}))]S_{u}^{\prime}(x^{k})\|_{X^{*}}\\ &\leq\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|J^{\prime}(u^{k+1})\|_{U^{*}}+L_{J^{\prime}}\|S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}\|u^{k+1}-S_{u}(x^{k})\|_{U}\\ &\leq N_{J^{\prime}}\|p^{k+1}-S_{u}^{\prime}(x^{k})\|_{\mathbb{L}(X;U)}+L_{J^{\prime}}N_{S_{u}^{\prime}}\|u^{k+1}-S_{u}(x^{k})\|_{U}.\end{split}$
Reduced adjoint: linear system splitting
With $U$, $W_{*}$, and $X$ normed spaces, let $S_{u}$ and $T$ be given by (1), and define
$\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})$
for $w^{k+1}\in W$ computed by taking (single or multiple) admissible splitting steps (see Sections˜3.1, 3.8 and 3.7) on the linear equation
$wT^{(u)}(u^{k+1},x^{k})+J^{\prime}(u^{k+1})=0$
with unknown $w=w^{k+1}$. Correspondingly, let $S_{w}$ arise from the reduced adjoint (6). The next theorem shows that the scheme satisfies Section˜2.1 in case of single steps; multiple steps follow immediately with large $\kappa_{w}$ by chaining Section˜2.1Item˜(ii).
Theorem 3.11.
Suppose that $T|U\times\Omega$ and $S_{u}|\Omega$ are Lipschitz-continuously differentiable for an $\Omega\subset X$, and that the adjoint admissible splitting Section˜3.1 holds in $U\times\Omega$ for the linear operators $A_{(u,x)}:w\mapsto wT^{(u)}(u,x)$. Suppose, moreover, that $T^{(u)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(u)};u}$-Lipschitz for all $x\in\Omega$; that $J^{\prime}$ is $L_{J^{\prime}}$-Lipschitz; that $S_{w}$ is $L_{S_{w}}$-Lipschitz in $\Omega$; and that
$N_{S_{w}}:=\sup\{\|S_{w}(x)\|\mid x\in\Omega\}<\infty.$
Let
$w^{k+1}:=-N^{-1}_{(u^{k+1},x^{k})}(J^{\prime}(u^{k+1})+M_{(u^{k+1},x^{k})}w^{k}).$
Then Section˜2.1Item˜(ii) holds in $\Omega$ with $\kappa_{w}=\zeta^{-1}$, $\mu_{u}=\zeta^{-1}\gamma^{-1}_{N}(L_{T^{(u)};u}N_{S_{w}}+L_{J})$, and $\pi_{w}=L_{S_{w}}$.
Suppose further that $T^{(x)}(\,\boldsymbol{\cdot}\,,x)$ is $L_{T^{(x)};u}$-Lipshitz for all $x\in\Omega$ and that
$M_{T^{(x)}}:=\sup\{\|T^{(x)}(u,x)\|\mid u\in U,x\in\Omega\}<\infty$
Then the differential transformation Section˜2.1Item˜(iii) holds for $\widetilde{F^{\prime}}(x^{k}):=w^{k+1}T^{(x)}(u^{k+1},x^{k})$ with $\alpha_{u}=N_{S_{w}}L_{T^{(x)};u}$ and $\alpha_{w}=M_{T^{(x)}}$.
Proof 3.12.
To prove Section˜2.1Item˜(ii), for brevity, let $v=(u^{k+1},x^{k})$, $A_{v}w=wT^{(u)}(v)$, and $b_{v}=-J^{\prime}(u^{k+1})$. Likewise let $\bar{v}=(S_{u}(x^{k}),x^{k})$, $A_{\bar{v}}w=wT^{(u)}(\bar{v})$, and $b_{\bar{v}}=-J^{\prime}(S_{u}(x^{k}))$. Assuming that $x^{k}\in\Omega$, we proceed as in [30, Lemma 4.1]: Observing that $A_{\bar{v}}S_{w}(x^{k})=b_{\bar{v}}$ and $A_{v}S_{w}(x^{k})=(N_{v}+M_{v})S_{w}(x^{k})$, we rearrange
$\begin{split}w^{k+1}-S_{w}(x^{k})&=N^{-1}_{v}(b_{v}-M_{v}w^{k})-S_{w}(x^{k})\\ &=N^{-1}_{v}(b_{v}-b_{\bar{v}})-N^{-1}_{v}(A_{v}-A_{\bar{v}})S_{w}(x^{k})-N^{-1}_{v}M_{v}(w^{k}-S_{w}(x^{k})).\end{split}$
From this it follows, using our assumptions, that
$\begin{split}\zeta^{-1}\|w^{k+1}-S_{w}(x^{k})\|&\leq\zeta^{-1}\gamma^{-1}_{N}(L_{J^{\prime}}+L_{T^{(u)};u}\|S_{w}(x^{k})\|)\|v-\bar{v}\|+\|w^{k}-S_{w}(x^{k})\|\\ &\leq\zeta^{-1}\gamma^{-1}_{N}(L_{J^{\prime}}+L_{T^{(u)};u}N_{S_{w}})\|u^{k+1}-S_{u}(x^{k})\|\\ &+\|w^{k}-S_{w}(x^{k-1})\|+L_{S_{w}}\|x^{k}-x^{k-1}\|.\end{split}$
Section˜2.1Item˜(iii) follows from
$\begin{split}\|\widetilde{F^{\prime}}(x^{k})&-F^{\prime}(x^{k})\|_{X^{*}}=\|w^{k+1}T^{(x)}(u^{k+1},x^{k})-S_{w}(x^{k})T^{(x)}(S_{u}(x^{k}),x^{k})\|_{X^{*}}\\ &=\|[w^{k+1}-S_{w}(x^{k})]T^{(x)}(u^{k+1},x^{k})-S_{w}(x^{k})[T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})]\|_{X^{*}}\\ &\leq\|T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\|w^{k+1}-S_{w}(x^{k})\|_{W}\\ &+\|S_{w}(x^{k})\|_{W}\|T^{(x)}(S_{u}(x^{k}),x^{k})-T^{(x)}(u^{k+1},x^{k})\|_{\mathbb{L}(X;W^{*})}\\ &\leq N_{S_{w}}L_{T^{(x)};u}\|u^{k+1}-S_{u}(x^{k})\|_{U}+M_{T^{(x)}}\|w^{k+1}-S_{w}(x^{k})\|_{W}.\end{split}$
4 Nonconvex forward-backward type methods with inexact updates
We now need to prove the convergence of outer methods for the overall problem (3), given estimates $\widetilde{F^{\prime}}(x^{k})$ of $F^{\prime}(x^{k})$ by inner and adjoint methods, the latter two satisfying the tracking theory of Section˜2. In this section, we do this through a convergence theory for general inexact forward backward-type methods in a normed space $X$. Our treatment encompasses primal-dual methods, seen as forward-backward methods with respect to appropriate operator-relative (semi-)norms, discussed in the previous section. We introduce such methods in Section˜4.1. Then in Section˜4.2 we introduce abstract growth conditions, which we will use in Sections˜4.3, 4.4, 4.5 and 4.6 to prove various forms of convergence.
Afterwards, in Section˜6, we will verify the growth inequalities for forward-backward and primal-dual algorithms that use the tracking theory of Section˜2 for (single-loop) updates of an inner problem.
4.1 General inexact forward-backward type methods
For proper $F,G:X\to\overline{\mathbb{R}}$, consider the problem
$\min_{x\in X}F(x)+G(x).$
In this subsection, and in the examples of Section˜6, $G$ will be convex and lower semicontinuous, and $F$ Fréchet differentiable, but the general theory of Sections˜4.2, 4.3, 4.4, 4.5 and 4.6 will make no such assumption.
For an initial $x^{0}$, if $X$ is Hilbert, the iterates $\{x^{k}\}_{k=1}^{\infty}$ of the basic inexact forward-backward method are generated for some step length parameter $\tau>0$ and an estimate $\widetilde{\nabla F}(x^{k})$ of $\nabla F(x^{k})$ (not necessarily the one from Section˜2) by
(19) $x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\widetilde{\nabla F}(x^{k})).$
In implicit form the method reads
(20) $\tau^{-1}(x^{k+1}-x^{k})\in\widetilde{\nabla F}(x^{k})+\partial G(x^{k+1}).$
We generalise this problem and method by considering for a skew-adjoint $\Xi\in\mathbb{L}(X;X^{*})$, i.e., $\Xi^{*}|X=-\Xi$, the problem of finding $x\in X$ satisfying
(21) $0\in H(x):=F^{\prime}(x)+\partial G(x)+\Xi x$
with an implicit method of the rough form
[…]
Algorithms of the form (22) with an exact inclusion for $\tilde{\partial}_{k+1}$, cover many common splitting algorithms, such as Douglas–Rachford splitting (DRS) and the primal-dual proximal splitting (PDPS) of [4]; see [7, 32]. As we will see in the following examples, with an inexact inclusion, besides the inexact gradients of Section˜2, the approach also covers inexact proximal maps and mismatched adjoints [19] in primal-dual methods. Inexact proximal maps were used, e.g., in [34] for point source localisation in measure spaces.
Example 4.1 (Primal-dual proximal splitting).
On normed spaces $Z$ and $Y$, let $g:Z\to\overline{\mathbb{R}}$ and $h:Y^{*}\to\overline{\mathbb{R}}$ be convex, proper, and lower semicontinuous, $f:Z\to\mathbb{R}$ possibly non-convex but Fréchet differentiable, and $K\in\mathbb{L}(Z;Y^{*})$. Suppose $h=(h_{*})^{*}$ for some $h_{*}:Y\to\overline{\mathbb{R}}$, and consider the problem
(23) $\min_{z\in Z}\penalty 10000\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\penalty 10000\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y).$
If $f$ is convex, subject to the standard condition on the existence of $x_{0}\in\operatorname{int}\operatorname{dom}[h\circ K]\cap\operatorname{dom}[f+g]\neq\emptyset$ with $Kx_{0}\in\operatorname{int}\operatorname{dom}h$, [^1] Several relaxations are possible, include using the relative interior, or the formulas of [2]. the Fenchel–Rockafellar theorem [7, Theorem 5.11] gives rise to the necessary and sufficient first-order primal-dual optimality conditions of the form (21),
$0\in H(z,y)=\begin{pmatrix}\partial g(z)+f^{\prime}(z)+K^{*}y\\ \partial h_{*}(y)-Kz.\end{pmatrix}=F^{\prime}(z,y)+\partial G(z,y)+\Xi(z,y),$
where
$F(z,y)=f(z),\quad G(z,y)=g(z)+h_{*}(y),\quad\text{and}\quad\Xi=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$.
If $f$ is nonconvex, the necessity can be shown through, e.g., Mordukhovich subdifferentials, and their compatibility with both convex subdifferentials and Fréchet derivatives; see, e.g., [7].
Pick step length parameters $\tau,\sigma>0$. With inexact gradients for $f$, the PDPS in Hilbert spaces reads
(24) $\left\{\begin{array}[]{l}z^{k+1}:=\operatorname{prox}_{\tau g}(z^{k}-\tau\widetilde{\nabla f}(z^{k})-\tau K^{*}y^{k}),\\ y^{k+1}:=\operatorname{prox}_{\sigma h_{*}}(y^{k}+\sigma K(2z^{k+1}-z^{k})).\end{array}\right$.
When $f=j\circ S_{u}$ for $S_{u}$ a PDE solution operator, and we compute $\widetilde{\nabla f}$ following Theorems˜3.5 and 3.11, (24) becomes the algorithm presented in [18].
To extend (24) to general normed spaces, we write it in $X=Z\times Y$ in the implicit form (22) as
(25a) $0\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}+M(x^{k+1}-x^{k})\quad\text{for}$
(25b) $\widetilde{F^{\prime}}(z^{k},y^{k}):=\begin{pmatrix}\widetilde{f^{\prime}}(z^{k})\\ 0\end{pmatrix}\quad\text{and}\quad M:=\begin{pmatrix}\tau^{-1}M_{z},&-K^{*}\\ -K&\sigma^{-1}M_{y}\end{pmatrix}$
for some self-adjoint positive semi-definite $M_{z}\in\mathbb{L}(Z;Z^{*})$ and $M_{y}\in\mathbb{L}(Y;Y^{*})$. For standard proximal maps in Hilbert spaces, $M_{z}=\operatorname{Id}$ and $M_{y}=\operatorname{Id}$. In that case, $M$ is self-adjoint and positive semi-definite when $\tau\sigma\|K\|^{2}\leq 1$, while the treatment of exact forward steps with respect to $f$ requires [^2] This is the requirement for gap estimates; for iterate estimates $\lambda/2$ in place of $\lambda$ is sufficient. In [36] an overall factor $4/3$ improvement is shown through an analysis that involves historical iterates. $\tau\lambda+\tau\sigma\|K\|^{2}\leq 1$ for $\lambda$ the Lipschitz factor of $f^{\prime}$ [32, 7, 15].
4.2 Inexact growth inequalities
We now provide several alternative assumptions that give a precise meaning to the approximate inclusion in (22). For this, we first define the Lagrangian gap functional
$\mathcal{G}(x;\bar{x}):=[F+G](x)-[F+G](\bar{x})-\langle\Xi x|\bar{x}\rangle_{X^{*},X}.$
Example 4.2.
For forward-backward splitting, $\mathcal{G}(x;\bar{x})=[F+G](x)-[F+G](\bar{x})$ is simply a function value difference.
Example 4.3.
For the PDPS of Example˜4.1, with $x=(y,z)$, we obtain the Lagrangian duality gap
$\mathcal{G}(x;\bar{x})=\mathcal{L}(z,\bar{y})-\mathcal{L}(\bar{z},y)\quad\text{for}\quad\mathcal{L}(z,y):=[f+g](z)+\langle Kz|y\rangle-h_{*}(y).$
This is different from the true duality gap that arises from the Fenchel–Rockafellar theorem. For the latter no convergence results exist to our knowledge. In the convex case, if $0\in H(\bar{x})$, the Lagrangian gap is non-negative, however, it may be zero even if $0\not\in H(\bar{x})$, unlike for the true duality gap.
For subdifferential convergence, we will need an inexact descent inequality, as well as bounds on sums of the gaps.
Assumption 4.3.
$M\in\mathbb{L}(X;X^{*})$ is self-adjoint and positive semi-definite. Also,
  1. (i)
    For a set $\Omega\subset X$, $\eta>0$, and $\mathbb{L}(X;X^{*})\ni\breve{\Lambda}\leq 2(1-\eta)M$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$, for some errors $\varepsilon_{\mathrm{desc},k}\in\mathbb{R}$, for any $k\in\mathbb{N}$, we have
    (26) $\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{\mathrm{desc},k}.$
  2. (ii)
    The errors satisfy $r_{\mathrm{desc}}:=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}<\infty$.
  3. (iii)
    We have $x^{0}\in\Omega$, and for any $N\geq 1$, $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})\leq r_{\mathrm{desc}}$ implies $x^{N}\in\Omega$.
  4. (iv)
    For some $0\leq\tilde{\eta}<\eta$, we have
    $\inf_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\left(\mathcal{G}(x^{k+1};x^{k})+\tilde{\eta}\|x^{k+1}-x^{k}\|_{M}^{2}\right)>-\infty.$
Remark 4.4.
If $\Omega=X$, convergence will be global. In the examples of Section˜3, $\Omega\neq X$ may arise from $S_{u}$, $G$, or $J$ being only locally Lipschitz continuously differentiable.
We will also need the approximate subdifferentials $\tilde{\partial}_{k+1}$ to become better as the distance between the iterates shrinks, in the sense of
Assumption 4.4.
For $H$ defined in (21), we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty\implies\lim_{k\to\infty}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}^{2}=0.$
This can be proved through Lipschitz differential estimates, as we discuss below.
For function value and iterate convergence, we cannot work with just the iterates: we need to assume properties with respect to a base point $\bar{x}\in X$, usually a solution. For iterate convergence, we assume the three-point monotonicity type estimate
(27) $\langle\tilde{\partial}_{k+1}-H(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}),$
for all $k\in\mathbb{N}$, whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ for an open neighbourhood $\Omega_{\bar{x}}$ of $\bar{x}$, a positive semi-definite self-adjoint $\breve{\Lambda}\in\mathbb{L}(X;X^{*})$, errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$, and a $\gamma\geq 0$. We recall here that $H(\bar{x})$ is a set, and the notation (27) means that the inequality holds for all elements of this set.
For function value convergence, we need again a descent inequality similar to (26), now instantiated at the base point $\bar{x}$ instead of $x^{k}$. That is, for all $k\in\mathbb{N}$, we assume for some errors $\varepsilon_{k}(\bar{x})\in\mathbb{R}$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$ that
(28) $\langle\tilde{\partial}_{k+1}|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
We write $\varepsilon_{\mathrm{desc},k}(\bar{x}):=\varepsilon_{k}(\bar{x})$ when we need draw a distinction to (27).
Example 4.5 (Growth conditions for basic forward-backward splitting).
Let $G$ be convex and $F$ possibly nonconvex but with $L$-Lipschitz Fréchet derivative. To model the basic exact forward-backward splitting in our general framework, we take $\Xi=0$ and $M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial G(x^{k+1})+F^{\prime}(x^{k})$. Then Eq.˜28 with $\breve{\Lambda}=L\operatorname{Id}$, $\gamma=0$, and $\varepsilon_{k}(\bar{x})=0$ is an immediate consequence of the definition of the convex subdifferential for $G$, and of the three-point descent inequality (10) for $F$. The monotonicity (27) likewise follows from the monotonicity of $\partial G$ and a three-point monotonicity property of $\nabla F$; see [7, Chapter 7].
The errors will also need to have a finite sum:
Assumption 4.5.
Given $\bar{x}\in X$, for some $\eta,\gamma\geq 0$ and $0\leq\breve{\Lambda}\leq(1-\eta)M$, either
  1. (a)
    (27) holds, and $\bar{x}\in H^{-1}(0)$; or
  2. (b)
    (28) holds, and $\inf_{x\in\Omega_{\bar{x}}}\mathcal{G}(x;\bar{x})\geq 0.$
Moreover, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for some $\delta>0$ and $p\geq 1$ with
[…]
4.3 Convergence of subdifferentials and quasi-monotonicity of values
We first show the potentially global convergence of subdifferentials; see Remark˜4.4. When $\Xi=0$, this could be followed by the Kurdyka–Łojasiewicz property to show function value convergence, and, afterwards, either by a growth condition or, in finite dimensions, a finite-length argument based on (30) and [1, proof of Lemma 2.6] to show iterate convergence. As the property can easily be verified only in finite dimensions (for semi-algebraic functions), we prefer a more direct approach.
Theorem 4.6.
If Section˜4.2 holds, then $x^{k}\in\Omega$ and
(30) $\mathcal{G}(x^{k+1};x^{k})+\eta\|x^{k+1}-x^{k}\|_{M}^{2}\leq\varepsilon_{\mathrm{desc},k}\quad\text{for all}\quad k\in\mathbb{N}.$
If, moreover, Section˜4.2 holds, then also $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$
Proof 4.7.
By the implicit algorithm (22), the properties of Fenchel conjugates (e.g., [7, Lemma 5.7]) and $M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}\in\partial\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)(x^{k+1}-x^{k})$, we have
(31) $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})=2\left(\frac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\right)^{*}(\tilde{\partial}_{k+1})=\|x^{k+1}-x^{k}\|_{M}^{2}=-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}.$
If $\{x^{j}\}_{j=0}^{N-1}\subset\Omega$, Section˜4.2Item˜(i) thus yields for all $k=0,\ldots,N-1$ that
(32) $\mathcal{G}(x^{k+1};x^{k})$ $=\mathcal{G}(x^{k+1};x^{k})-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}-\|x^{k+1}-x^{k}\|_{M}^{2}$
$\leq\varepsilon_{\mathrm{desc},k}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{2M-\breve{\Lambda}}^{2}\leq\varepsilon_{\mathrm{desc},k}-\eta\|x^{k+1}-x^{k}\|_{M}^{2}.$
Summing over all such $k$, and using Section˜4.2Item˜(ii), it follows
(33) $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})+\sum_{k=0}^{N-1}\eta(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})=\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})+\sum_{k=0}^{N-1}\eta\|x^{k+1}-x^{k}\|_{M}^{2}\leq r_{\mathrm{desc}}.$
From Section˜4.2Item˜(iii), it now follows that $x^{N}\in\Omega$. Since, by the same assumption, $x^{0}\in\Omega$, induction establishes (30) and $x^{k}\in\Omega$ for all $k\in\mathbb{N}$. Using Section˜4.2Item˜(iv) in (33), we, moreover, deduce $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ and $(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1})\to 0$. Let $c\geq\|M\|_{\mathbb{L}(X;X^{*})}$. By $\|\,\boldsymbol{\cdot}\,\|_{M}^{2}\leq c\|\,\boldsymbol{\cdot}\,\|_{X}^{2}$ and the properties of conjugates (e.g., [7, Lemmas 5.4 and 5.7]),
$\frac{4}{c}\|\tilde{\partial}_{k+1}\|_{X^{*}}^{2}=c\|2\tilde{\partial}_{k+1}/c\|_{X^{*}}^{2}=(c\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(2\tilde{\partial}_{k+1})\leq(\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(2\tilde{\partial}_{k+1}).$
Thus also $\|\tilde{\partial}_{k+1}\|_{X^{*}}\to 0$. Section˜4.2 proves that $\inf_{x^{*}\in H(x^{k+1})}\|\tilde{\partial}_{k+1}-x^{*}\|_{X^{*}}\to 0$. Hence an application of the triangle inequality establishes $\inf_{x^{*}\in H(x^{k+1})}\|x^{*}\|_{X^{*}}\to 0$.
Remark 4.8 (Forward-backward splitting).
For the (inexact) forward-backward splitting of (20), once we verify the relevant assumptions in Corollary˜6.7, the previous theorem establishes the monotonicity of function values, as well as the convergence of subdifferentials to zero, $\inf_{x^{*}\in\partial G(x^{k+1})}\|F^{\prime}(x^{k+1})+x^{*}\|\to 0$.
Remark 4.9 (Bregman divergences).
The argument of Theorem˜4.6 extends to algorithms where the $M$-seminorm for $M\in\mathbb{L}(X;X^{*})$ is replaced by a Bregman divergence $B_{M}$ generated by some convex $M:X\to\overline{\mathbb{R}}$. Recalling the definition
(34) $B_{M}^{\omega}(x,z):=M(z)-M(x)-\langle\omega|z-x\rangle\quad(\omega\in\partial M(x);x,z\in X),$
in the algorithm (22), we would impose $\tilde{\partial}_{k+1}=\omega^{k+1}-\omega^{k}\in\partial_{2}B_{M}^{\omega^{k}}(x^{k},x^{k+1})$ for a given $\omega^{k}\in\partial M(x^{k})$ and some $\omega^{k+1}\in\partial M(x^{k+1})$. Using the Fenchel–Young identity, we could then replace (31) by
$B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})$ $=B_{M}^{\omega^{k}}(x^{k},x^{k+1})=M(x^{k+1})-M(x^{k})-\langle\omega^{k}|x^{k+1}-x^{k}\rangle$
$\leq\langle\omega^{k+1}-\omega^{k}|x^{k+1}-x^{k}\rangle=-\langle\tilde{\partial}_{k+1}|x^{k+1}-x^{k}\rangle_{X^{*},X}.$
Thus repeating the arguments of the theorem would establish both $B_{M}^{\omega^{k}}(x^{k},x^{k+1})\to 0$ as well as $B_{M^{*}}^{x^{k+1}}(\omega^{k+1},\omega^{k})=B_{M^{*}}^{x^{k+1}}(\omega^{k}+\tilde{\partial}_{k+1},\omega^{k})\to 0$. A variant of Section˜4.2 could then establish a form of convergence for $H(x^{k+1})$.
4.4 Non-escape, quasi-Féjer monotonicity, linear convergence
The next lemma is essential for all our strong convergence results. The proof is standard; see, e.g., [7, Chapter 15] for the case $\varepsilon_{k}(\bar{x})=0$ and $\Xi=0$. Observe that (35) with the triangle inequality may be used to again prove Section˜2.1Item˜(i) for multilevel methods.
Lemma 4.10.
Suppose Section˜4.2 holds at $\bar{x}\in X$. Then $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, and the sequence is ($p$-strongly) quasi-Féjer, i.e.,
(35) $\frac{p}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}\leq\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}+\varepsilon_{k}(\bar{x}).$
Moreover, $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ if $\eta>0$.
Proof 4.11.
We first treat Section˜4.2 option Item˜(a). Fix $N\in\mathbb{N}$ and suppose $\{x^{j}\}_{j=0}^{N-1}\subset\Omega_{\bar{x}}$. Observe that $\langle\Xi x|x\rangle=0$ for all $x\in X$ by the skew-adjointness of $\Xi$. Since $0\in H(\bar{x})$, using (27) in the implicit algorithm (22), we thus get
$-\langle M(x^{k+1}-x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ $\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x})$
for all $k\in\{0,\ldots,N-1\}$. By $\breve{\Lambda}\leq(1-\eta)M$ and the Pythagoras’ identity (see [35, (2.3)])
(36) $\langle M(x-z)|x-\bar{x}\rangle=\frac{1}{2}\|x-z\|_{M}^{2}+\frac{1}{2}\|x-\bar{x}\|_{M}^{2}-\frac{1}{2}\|z-\bar{x}\|_{M}^{2}\quad(x,z,\bar{x}\in X),$
we obtain
(37) $\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}\geq\frac{\eta}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1+2\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Using $1+2\gamma\geq p$, multiplying by $p^{k}$, and summing over $k=0,\ldots,N-1$ yields
(38) $\frac{1}{2}\|x^{0}-\bar{x}\|_{M}^{2}+\sum_{k=0}^{N-1}p^{k}\varepsilon_{k}(\bar{x})\geq\sum_{k=0}^{N-1}\frac{\eta p^{k}}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{p^{N}}{2}\|x^{N}-\bar{x}\|_{M}^{2}.$
Multiplying by $p^{-N}\leq 1$ and using $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and (29), it follows
(39) $\frac{\delta^{2}}{2}=\frac{\delta^{2}-2r_{p}}{2}+r_{p}>\sum_{k=0}^{N-1}\frac{\eta p^{k-N}}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1}{2}\|x^{N}-\bar{x}\|_{M}^{2}.$
Hence $x^{N}\in\mathbb{O}_{M}(\bar{x},\delta)$. Since $x^{0}\in\Omega_{\bar{x}}$ by Section˜4.2, an inductive argument shows that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\bar{x}}$ for all $k\in\mathbb{N}$, justifying the above steps. Finally, (37) shows (35), while the claim $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$ follows from (39) and $\eta>0$.
Regarding option Section˜4.2Item˜(b), arguing as above with (28) in place of (27), we get in place of (37) the estimate
(40) $\frac{1}{2}\|x^{k}-\bar{x}\|_{M}^{2}\geq\mathcal{G}(x^{k+1};\bar{x})+\frac{\eta}{2}\|x^{k+1}-x^{k}\|_{M}^{2}+\frac{1+\gamma}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Using $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$, we proceed (with $1+\gamma\geq p$) as in option Item˜(a) to establish (39), and from there onwards.
A closer look at (38) immediately yields linear convergence if $p>1$ and we remove $p^{-N}$ from (29).
Corollary 4.12.
Suppose Section˜4.2 holds at $\bar{x}\in X$ with $p>1$ and the inequality in (29) strengthened to
(41) $\frac{1}{2}\delta^{2}>\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}\varepsilon_{k}(\bar{x})<\infty.$
Then $\|x^{N}-\bar{x}\|_{M}^{2}\to 0$ at the rate $O(p^{-N})$.
4.5 Local convergence of function values
We now proceed to function values and duality gaps. The idea of possibly assuming both Section˜4.2Item˜(a) and a relaxed version of Item˜(b), as an alternative to just the latter, is to be able to study descent at non-minimising critical points. For simplicity, we only treat sublinear convergence.
Theorem 4.13.
Suppose Section˜4.2 holds at $\bar{x}\in X$ and, for a non-empty set $\hat{X}\subset X$, Eq.˜28 holds for all $\hat{x}\in\hat{X}$ with $\breve{\Lambda}=\breve{\Lambda}_{\hat{x}}\leq M$, $\gamma=\gamma_{\hat{x}}\geq 0$, and $\Omega_{\hat{x}}\supset\mathbb{O}_{M}(\bar{x},\delta)$. Then
(42) $\sup_{\hat{x}\in\hat{X}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\leq\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}(\hat{x})\right)\quad\text{for all}\quad N\in\mathbb{N}.$
If $\Xi=0$ and Section˜4.2 holds [^1] Since the proof of the present Theorem 4.13 shows that $x^{k}\in\mathbb{O}_{M}(\hat{x},\delta)$ for all $k\in\mathbb{N}$, to prove the required (30), it would be enough to assume that just Section 4.2Item (i) holds with $\Omega\supset\mathbb{O}_{M}(\hat{x},\delta)$., then, for all $N\in\mathbb{N}$,
(43) $F+G](x^{N})\leq\inf_{\hat{x}\in\hat{X}}[F+G](\hat{x})+\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2N}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\left(\frac{1}{N}\varepsilon_{\mathrm{desc},k}(\hat{x})+\frac{k+1}{N}\varepsilon_{\mathrm{desc},k}\right)\right).$
Proof 4.14.
Lemma˜4.10 shows for all $k\in\mathbb{N}$ that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\bigcap_{\hat{x}\in\hat{X}}\Omega_{\hat{x}}$. Hence, for any $\hat{x}\in\hat{X}$, we may follow the proof of the lemma for case Item˜(b) of Section˜4.2 to establish (40) for $\bar{x}=\hat{x}$. To reach this point, the assumption $\inf_{x\in\mathbb{O}_{M}(\delta,\bar{x})}\mathcal{G}(x;\bar{x})\geq 0$ was not yet needed. Now, summing (40) over $k=0,\ldots,N-1$, we obtain
(44) $\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\varepsilon_{\mathrm{desc},k}(\hat{x})\geq\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})+\frac{1}{2}\|x^{N}-\hat{x}\|_{M}^{2}.$
Taking the supremum over $\hat{x}\in\hat{X}$, this establishes (42).
[…]
4.6 Weak convergence
We next prove the weak convergence of the iterates. We call the self-adjoint and positive semi-definite preconditioner $M\in\mathbb{L}(X;X^{*})$ admissible for weak convergence if $\|x^{k}\|_{M}\to 0$ implies $Mx^{k}\to 0$.
Example 4.15.
Suppose $M=A^{*}A$ for some $A\in\mathbb{L}(X;V)$ for a Hilbert space $V$. Then the $M$-semi-norm convergence clearly implies $Ax^{k}\to 0$, and consequently $Mx^{k}\to 0$. Thus $M$ is weakly admissible. In Hilbert spaces, every positive semi-definite self-adjoint operator has such a square root $A$ with $V=X$. For a convolution-based construction in the space of Radon measures, see [35, Theorem 2.4].
Theorem 4.16.
[…]
Proof 4.17.
Lemma˜4.10 proves that $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)$ for all $k\in\mathbb{N}$, as well as that $\sup_{N\in N}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{M}^{2}<\infty$. The latter establishes $\|x^{k+1}-x^{k}\|_{M}\to 0$, and through admissibility for weak convergence, and (22), that $\tilde{\partial}_{k+1}=-M(x^{k+1}-x^{k})\to 0$ strongly in $X^{*}$. Moreover, Section˜4.2 yields $\|\tilde{\partial}_{k+1}-x^{*}_{k+1}\|_{X^{*}}\to 0$ for some $x^{*}_{k+1}\in H(x^{k+1})$. Consequently $x^{*}_{k+1}\to 0$. Since $x^{k}\in\mathbb{O}_{M}(\bar{x},\delta)\subset\Omega_{\hat{x}}$, as in the proof of Lemma˜4.10, we show the quasi-Féjer monotonicity (35) for all $\hat{x}\in\hat{X}$ and $k\in\mathbb{N}$.
Suppose then that $x^{k_{j}+1}\mathrel{\rightharpoonup}\hat{x}$ for a subsequence $\{k_{j}\}_{j\in\mathbb{N}}\subset\mathbb{N}$ and a $\hat{x}\in X$. We want to show that $\hat{x}\in\hat{X}$. We consider two cases:
  1. 1.
[…]
  2. 2.
[…]
Example 4.18.
In the setting of Section˜2 and Theorem˜2.8, the weak-*-to-strong continuity of $F^{\prime}$ can be achieved, for example, when $F(x)=\frac{1}{2}\|S(x)-b\|^{2}$ for a Lipschitz and bounded $S$ with finite-dimensional range.
Remark 4.19.
All of our theory also applies when $X$ is the dual space of a separable normed space $X_{*}$, and we replace in our definitions $X^{*}$ by the predual space $X_{*}$, that is, subdifferentials are subsets of $X_{*}$, and $M,\Lambda\in\mathbb{L}(X;X_{*})$, etc. With this change the theory applies, for example, to $X$ a space of Radon measures, as in [35]. Then Theorem˜4.16 proves the weak-* convergence.
5 Operator-relative regularity
To facilitate treating primal-dual methods as forward-backward methods with respect to suitable operators in the next section, we introduce operator-relative smoothness and monotonicity concepts. That is, we will not use the standard norms of the space, but (semi-)norms of the form $\|x\|_{\Lambda}:=\sqrt{\langle x|\Lambda x\rangle_{X,X^{*}}}$ for a positive semi-definite self-adjoint $\Lambda\in\mathbb{L}(X;X^{*})$. This will be useful for the treatment of primal-dual methods as forward-backward methods with respect to appropriate (semi-)norms. We start with the relevant definitions in Section˜5.1, and then prove the relevant operator-relative descent inequalities and three-point monotonicity in Section˜5.2.
5.1 Definitions
For a self-adjoint positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$ on a normed space $X$, we say that the Gâteaux derivative $DF$ of $F:X\to\mathbb{R}$ is $\Lambda$-firmly Lipschitz if
$\langle DF(z)-DF(x)|h\rangle_{X^{*},X}\leq\|x-z\|_{\Lambda}\|h\|_{\Lambda}\quad(x,z,h\in X).$
This implies for any $h\in X$ with $\|h\|_{X}\leq 1$ that
(45) $\|DF(z)-DF(x)\|_{X^{*}}^{2}\leq\|z-x\|_{\Lambda}^{2}\|h\|_{\Lambda}^{2}\leq\|\Lambda\|_{\mathbb{L}(X;X^{*})}\|z-x\|_{\Lambda}^{2}\leq\|\Lambda\|_{\mathbb{L}(X;X^{*})}^{2}\|z-x\|_{X}^{2}.$
In particular, $DF$ is $\|\Lambda\|$-Lipschitz.
Likewise, we call $DF$ locally $\Gamma$-monotone in $\Omega\ni\bar{x}$ for a self-adjoint $\Gamma\in\mathbb{L}(X;X^{*})$ if
$\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle\geq q_{\Gamma}(z-\bar{x})\quad(z\in\Omega).$
where we write
(46) $q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}$
when $\|\,\boldsymbol{\cdot}\,\|_{\Gamma}$ may not be a well-defined (semi-)norm. We do not at this stage assume $\Gamma$ to be positive semi-definite. We, however, call $\Gamma$ Young if there exists a self-adjoint positive semi-definite $|\Gamma|\in\mathbb{L}(X;X^{*})$ such that
$2\langle\Gamma x|z\rangle_{X^{*},X}\leq\|x\|_{|\Gamma|}^{2}+\|z\|_{|\Gamma|}^{2}\quad(x,z\in X).$
Likewise, we call $G$ $\Gamma$-subdifferentiable and $\partial G$ $\Gamma$-monotone if, respectively,
$G(\tilde{x})-G(x)\geq\langle q|\tilde{x}-x\rangle+\frac{1}{2}q_{\Gamma}(\tilde{x}-x)\quad\text{or}\quad\langle\tilde{q}-q|\tilde{x}-x\rangle\geq q_{\Gamma}(\tilde{x}-x)$
for all $q\in\partial G(x)$; $\tilde{q}\in\partial\tilde{G}(x)$, and $x,\tilde{x}\in X$.
Remark 5.1.
Aside from $\|\,\boldsymbol{\cdot}\,\|_{M}$, introduced in the next section, which presently needs to satisfy the Pythagoras’ identity, our work does not strictly depend on $\|\,\boldsymbol{\cdot}\,\|_{\Lambda}$, $\|\,\boldsymbol{\cdot}\,\|_{|\Gamma|}$ or $q_{\Gamma}$ being produced by operators. They could be arbitrary semi-norms and quadratics, if we introduced the formal calculus $q_{a\Lambda+b\Gamma}:=a\|\,\boldsymbol{\cdot}\,\|_{\Lambda}^{2}+bq_{\Gamma}$, etc., for $a,b\in\mathbb{R}$, and defined the concept of $\Gamma$ being Young appropriately. For simplicity, we have chosen to work with operators.
5.2 Estimates
We first prove a $\Lambda$-firmly Lipschitz descent lemma, as a generalisation of the basic descent inequality (8).
Lemma 5.2.
On a normed space $X$, suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Gâteaux derivative for a self-adjoint positive semi-definite $\Lambda\in\mathbb{L}(X;X^{*})$. Then
(47) $F(x)-F(z)-\langle DF(z)|x-z\rangle_{X^{*},X}\leq\frac{1}{2}\|z-x\|_{\Lambda}^{2}.$
Proof 5.3.
By the mean value theorem and the assumed firm Lipschitz property,
$F(x)-F(z)-\langle DF(z)|x-z\rangle_{X^{*},X}=\int_{0}^{1}\langle DF(z+t(x-z))-DF(z)|x-z\rangle\,\mathrm{d}t\leq\int_{0}^{1}t\|x-z\|_{\Lambda}^{2}\,\mathrm{d}t.$
Integrating, the claim follows.
The next result provides an operator-relative generalisation of the three-point descent inequality (10). Compared to the treatment of the latter in [32, Appendix B], it is important that $x$ ($x^{k+1}$ in the application to forward steps at $x^{k}$) is not a priori restricted to the neighbourhood $\Omega$ of $\Gamma$-monotonicity at $\bar{x}$.
Lemma 5.4.
On a normed space $X$, let $F:X\to\mathbb{R}$ and suppose $DF$ is $\Lambda$-Lipschitz for some $\Lambda\in\mathbb{L}(X;X^{*})$, and $\Gamma$-monotone at $\bar{x}\in X$ in a convex neighbourhood $\Omega\ni\bar{x}$ and a Young $\Gamma\in\mathbb{L}(X;X^{*})$. Then, for any $\beta>0$, for all $z\in\Omega$ and $x\in X$,
$\langle DF(z)|x-\bar{x}\rangle\geq F(x)-F(\bar{x})+\frac{1}{2}q_{\Gamma-\beta|\Gamma|}(x-\bar{x})-\frac{1}{2}q_{\Lambda+\beta^{-1}|\Gamma|-\Gamma}(x-z).$
Proof 5.5.
Similarly to the proof of the descent inequality in Lemma˜5.2, the mean value theorem applied to $\varphi(t):=F(\bar{x}+t(z-\bar{x}))$, followed by the assumed local $\Gamma$-monotonicity of $DF$, and the Young property of $\Gamma$, establishes
$F$ $(\bar{x})-F(z)-\langle DF(z)|\bar{x}-z\rangle_{X^{*},X}$
$=\int_{0}^{1}\langle DF(z+t(\bar{x}-z))-DF(z)|\bar{x}-z\rangle\,\mathrm{d}t\geq\int_{0}^{1}tq_{\Gamma}(\bar{x}-z)\,\mathrm{d}t=\frac{1}{2}q_{\Gamma}(\bar{x}-z)$
$=\frac{1}{2}q_{\Gamma}(x-\bar{x})+\frac{1}{2}q_{\Gamma}(x-z)-\langle\Gamma(x-\bar{x})|x-z\rangle\geq\frac{1}{2}q_{\Gamma-\beta|\Gamma|}(x-\bar{x})+\frac{1}{2}q_{\Gamma-\beta^{-1}|\Gamma|}(x-z).$
Applying Lemma˜5.2 and summing this inequality with the descent inequality it provides, we obtain the claim.
Finally, we provide a three-point monotonicity inequality that can be used instead of Lemma˜5.4 when function values are not desired. The proof we present here slightly improves [7, Lemma 15.1] even in the Hilbert space scalar factor case.
Lemma 5.6.
On a normed space $X$, let $F:X\to\mathbb{R}$ and suppose $DF$ is $\Lambda$-Lipschitz for some $\Lambda\in\mathbb{L}(X;X^{*})$, and $\Gamma$-monotone for a self-adjoint $\Gamma\in\mathbb{L}(X;X^{*})$ in a neighbourhood $\Omega\ni\bar{x}$ for some $\bar{x}\in X$. Then, for any $\beta,\zeta>0$, for all $z\in\Omega$ and $x\in X$, with $\tilde{\Gamma}:=\Gamma-(\zeta/2)\Lambda$, we have
$\langle DF(z)-DF(\bar{x})|x-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}-\beta|\tilde{\Gamma}|}(x-\bar{x})-q_{\Lambda/(2\zeta)+\beta^{-1}|\tilde{\Gamma}|-\tilde{\Gamma}}(x-z).$
Proof 5.7.
Using both the $\Gamma$-monotonicity and the $\Lambda$-firmly Lipschitz property, and finishing with Young’s inequality, we obtain
$\langle DF(z)-DF(\bar{x})|x-\bar{x}\rangle_{X^{*},X}$ $=\langle DF(z)-DF(\bar{x})|z-\bar{x}\rangle_{X^{*},X}+\langle DF(z)-DF(\bar{x})|x-z\rangle_{X^{*},X}$
$\geq q_{\Gamma}(z-\bar{x})-\|z-\bar{x}\|_{\Lambda}\|x-z\|_{\Lambda}\geq q_{\tilde{\Gamma}}(z-\bar{x})-\frac{1}{2\zeta}\|x-z\|_{\Lambda}^{2}.$
Arguing for $q_{\tilde{\Gamma}}$ as in the proof of Lemma˜5.4, we obtain the claim.
6 Outer algorithms
We now explicitly verify Sections˜4.2, 4.2 and 4.2 for both basic forward-backward splitting and the PDPS, as well as their inexact versions based on the estimation of $F^{\prime}(x^{k})$ by $\widetilde{F^{\prime}}(x^{k})$ formed using inner and adjoint algorithms satisfying the tracking theory of Section˜2. We first provide in Section˜6.1 a general result for opeartor-relative forward-backward type methods for (21). This forms the basis of verification of the assumptions both for forward-backward splitting in Section˜6.2, and primal-dual methods in Section˜6.3. In the final Section˜6.4 we also treat adjoint mismatch as another example of our general approach to inexactness.
6.1 A general result
Let $M,\Lambda\in\mathbb{L}(X;X^{*})$ be self-adjoint postive and semi-definite, with $M$ preconditioning operator from Section˜4. To use the tracking theory of Section˜2, we will impose the choices
(48) ${d_{X}(x,\tilde{x})}=\|x-\tilde{x}\|_{M},\enskip{d_{X^{*}}(x^{*},\tilde{x}^{*})}=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*}-\tilde{x}^{*})]^{1/2},\enskip{b_{X}(x^{k+1},x^{k})}=\|x^{k+1}-x^{k}\|_{\Lambda}.$
Then the Fenchel–Young inequality and homogeneity ensure that $d_{X^{*}}$} is Young to $d_{X}$}, as defined in Section˜2.2. If $M$ is invertible, ${d_{X^{*}}(x^{*},\tilde{x}^{*})}=\|x^{*}-\tilde{x}^{*}\|_{M^{-1}}$. More generally, we have the following bound.
Lemma 6.1.
With $d_{X^{*}}$ as in (48), $\|x^{*}-\tilde{x}^{*}\|_{X^{*}}\leq\|M\|_{\mathbb{L}(X;X^{*})}^{1/2}{d_{X^{*}}(x^{*},\tilde{x}^{*})}$ for all $x^{*},\tilde{x}^{*}\in X^{*}$.
Proof 6.2.
We have $f^{*}\leq g^{*}$ for $f\geq g$ from the definition of the Fenchel conjugate. Abbreviating $m:=\|M\|_{\mathbb{L}(X;X^{*})}$, and using [7, Lemma 5.8 and Lemma 5.5], it follows for any $x^{*}\in X^{*}$ that
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}(x^{*})\geq(\tfrac{m}{2}\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(x^{*})=m(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{X}^{2})^{*}(x^{*}/m)=\frac{m}{2}\|x^{*}/m\|_{X^{*}}^{2},=\frac{1}{2m}\|x^{*}\|_{X^{*}}^{2}.$
[…]
In brief, the next theorem says that
  1. 1.
    Sections˜4.2 and 4.2, used for subdifferential convergence by Theorem˜4.6, require that the initialisation $x^{0}$ be in the set $\Omega$ where the tracking assumptions hold, and that the step lengths (encoded in $M$) be small compared to the operator-relative Lipschitz factor $\Lambda$.
  2. 2.
    Section˜4.2, used for stronger convergence results by Corollaries˜4.12, 4.13 and 4.16, also requires sufficient strong subdifferentiability.
Theorem 6.3.
On a normed space $X$, let $M,\Lambda\in\mathbb{L}(X;X^{*})$ be self-adjoint and positive semi-definite. Suppose $F:X\to\mathbb{R}$ has a $\Lambda$-firmly Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Given an initial $x^{0}\in X$, for all $k\in\mathbb{N}$, construct $\widetilde{F^{\prime}}(x^{k})$ obeying Section˜2.1 (or, more generally, Theorems˜2.5 and 2.8 without Section˜2.1) for the distances (48) and an $\Omega\subset X$. Update $x^{k+1}$ by solving
(49) $0\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}+M(x^{k+1}-x^{k}).$
Let $\varsigma_{p}$, $\kappa$ and $e_{p,k}$ be as in Theorem˜2.5. Then,
  1. (i)
[…]
    and, for some $\tilde{\gamma}>0$,
    $0\leq\breve{\Lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+\tilde{\gamma}M\leq 2(1-\eta)M.$
  2. (ii)
    Section˜4.2 holds if $\Lambda\leq cM$ for a $c>0$.
Suppose further that $G$ is $\Gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\Gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}_{M}(\bar{x},\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\Gamma_{F},\Gamma_{G}\in\mathbb{L}(X;X^{*})$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If
$x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})\quad\text{with}\quad r_{p}=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}e_{p,k}/(2\tilde{\gamma})<\delta^{2}/2,$
[…]
  1. 3.
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\Gamma}_{F}:=\Gamma_{F}-(\zeta/2)\Lambda$ and some $\zeta,\beta>0$,
    $(\gamma+\tilde{\gamma})M$ $\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|\quad\text{for a}\quad\gamma\geq(p-1)/2\quad\text{and}$
    $0\leq\breve{\Lambda}$ $:=(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+2(\beta^{-1}|\tilde{\Gamma}_{F}|-\tilde{\Gamma}_{F})\leq(1-\eta)M.$
  2. 4.
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$,
    $(\gamma+\tilde{\gamma})M$ $\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|\quad\text{for a}\quad\gamma\geq p-1\quad\text{and}$
    $0\leq\breve{\Lambda}$ $:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}\leq(1-\eta)M.$
Proof 6.4.
Item˜(i): By Lemma˜5.2, we have
$\langle F^{\prime}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}.$
Combining this with Theorem˜2.5 for $\bar{x}=x^{k}$ establishes (compare (9))
$\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-x^{k}\rangle_{X^{*},X}\geq F(x^{k+1})-F(x^{k})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}$
with $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}<\infty$ whenever $\{x^{n}\}_{n=0}^{k}\subset\Omega$. Further combining with the subdifferentiability of $G$, we verify (26). Since we assume $\breve{\Lambda}\leq 2(1-\eta)M$ and take $\varepsilon_{\mathrm{desc},k}\propto e_{p,k}$, this verifies Section˜4.2Items˜(i) and (ii). Because $\Xi=0$, Item˜(iii) requires $F+G](x^{N})\leq r_{\mathrm{desc}}+[F+G](x^{0})$ to imply $x^{N}\in\Omega$. This holds whenever $\Omega\supset\operatorname{sub}_{r_{\mathrm{desc}}+[F+G](x^{0})}(F+G)$, as we have assumed. Likewise, we prove Item˜(iv) with the lower bound $\inf[F+G]-[F+G](x^{0})>$-$\infty$.
Item˜(ii): We have $\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}\leq\|F^{\prime}(x^{k+1})-\widetilde{F^{\prime}}(x^{k})\|_{X^{*}}$ through the choice
$x_{k+1}^{*}=F^{\prime}(x^{k+1})-\widetilde{F^{\prime}}(x^{k})+\tilde{\partial}_{k+1}\in F^{\prime}(x^{k+1})+\partial G(x^{k+1})+\Xi x^{k+1}=H(x^{k+1}).$
Hence, the triangle inequality, Eqs.˜45 and 6.1 followed by Theorem˜2.8 establish
(50) $\begin{split}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}&\leq\|F^{\prime}(x^{k})-F^{\prime}(x^{k+1})\|_{X^{*}}+\|\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})\|_{X^{*}}\\[-8.61108pt] &\leq\|\Lambda\|^{1/2}\|x^{k+1}-x^{k}\|_{\Lambda}+\|M\|_{\mathbb{L}(X;X^{*})}^{1/2}{d_{X^{*}}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\\ &\leq\|\Lambda\|^{1/2}\|x^{k+1}-x^{k}\|_{\Lambda}+\left(\|M\|_{\mathbb{L}(X;X^{*})}e_{\mathrm{lip},k}\right)^{1/2},\end{split}$
where $e_{\mathrm{lip},k}\geq 0$ satisfies
(51) $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}<\infty\implies\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}e_{\mathrm{lip},k}<\infty.$
The antecedent of (51) implies $\lim_{k\to\infty}\inf_{x_{k+1}^{*}\in H(x^{k+1})}\|x_{k+1}^{*}-\tilde{\partial}_{k+1}\|_{X^{*}}=0$ through Eqs.˜50 and 51. Since $\Lambda\leq cM$ by assumption, this establishes Section˜4.2.
[…]
We have also explicitly assumed the remaining neighbourhood conditions of Section˜4.2, as well as $0\leq\breve{\Lambda}\leq(1-\eta)M$, so only need to verify the respective (27) or (28).
Item˜3: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma˜5.6 and our assumption $\breve{\Lambda}\geq 0$, since $x^{k}\subset\Omega_{\bar{x}}$, we have
$\langle F^{\prime}(x^{k})-F^{\prime}(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\zeta^{-1}\Lambda+2(\beta^{-1}|\tilde{\Gamma}_{F}|-\tilde{\Gamma}_{F})}^{2}.$
Here we recall from (46) the notation $q_{\Gamma}(x):=\langle\Gamma x|x\rangle_{X^{*},X}$, used when $\|\,\boldsymbol{\cdot}\,\|_{\Gamma}$ may not be a well-defined (semi-)norm. Combining this with Theorem˜2.5 and the definition of $\breve{\Lambda}$ gives
$\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
Since $G$ is $\Gamma_{G}$-monotone and $\Xi$ is skew-symmetric, we have
$\langle\partial G(x^{k+1})-\partial G(\bar{x})+\Xi(x^{k+1}-\bar{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq\|x^{k+1}-\bar{x}\|_{\Gamma_{G}}.$
Due to (49), we have $M(x^{k+1}-x^{k})=\tilde{\partial}_{k+1}\in\widetilde{F^{\prime}}(x^{k})+\partial G(x^{k+1})+\Xi x^{k+1}$. By definition, $H(x)=\partial G(x)+F^{\prime}(x)+\Xi x$. Hence, summing the previous two inequalities, it follows that
$\langle\tilde{\partial}_{k+1}-H(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq q_{\Gamma_{G}+\tilde{\Gamma}_{F}-\beta\tilde{\Gamma}_{F}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).$
After an application of the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\tilde{\Gamma}_{F}-\beta|\tilde{\Gamma}_{F}|$, this verifies (27) through the.
Item˜4: Suppose $\{x^{n}\}_{n=0}^{k}\subset\Omega_{\bar{x}}$. By Lemma˜5.4, since $x^{k}\subset\Omega_{\bar{x}}$, we have
$\langle F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}\geq F(x^{k+1})-F(\bar{x})+\frac{1}{2}q_{\Gamma_{F}-\beta{|\Gamma_{F}|}}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}}^{2}.$
Combining this with Theorem˜2.5 gives and the definition of $\breve{\Lambda}$ gives (compare (11))
$\begin{split}\langle\widetilde{F^{\prime}}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}&\geq F(x^{k+1})-F(\bar{x})\\ &+\frac{1}{2}q_{\Gamma_{F}-\beta{|\Gamma_{F}|}-\tilde{\gamma}M}(x^{k+1}-\bar{x})-\frac{1}{2}\|x^{k+1}-x^{k}\|_{\breve{\Lambda}}^{2}-\varepsilon_{k}(\bar{x}).\end{split}$
Similarly to the claim Item˜3, we now verify (28) by combining this with the ($\Gamma_{G}$-strong) subdifferentiability of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$.
Remark 6.5 (error term).
Recalling Remark˜2.7, we can make $r_{\mathrm{desc}}\geq 0$ and $r_{p}\geq 0$ arbitrarily small by taking high-quality first inner and adjoint steps.
Remark 6.6 (linear convergence).
From (52) and Theorem˜2.5, we see that through good-quality first inner and adjoint steps, (41) can be made to hold. Then Corollary˜4.12 is applicable to results satisfying Section˜4.2 for $p>1$ through Theorem˜6.3.
6.2 Forward-backward splitting
We now interpret Theorem˜6.3 for both standard exact forward-backward splitting in a Hilbert space, as well as outer forward-backward splitting when we construct $\widetilde{\nabla F}$ with inner and adjoint methods that satisfy the tracking theory of Section˜2; in particular, the methods of Section˜3.
For clarity of the statement of the next corollary, we omit any mention of parameters such as $\eta$ and $\tilde{\eta}$ that are not important for the algorithm itself, and that can be deduced from the other choices.
Corollary 6.7 (Inexact outer forward-backward splitting on a Hilbert space).
On a Hilbert space $X$, suppose $F:X\to\mathbb{R}$ has an $L$-Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Pick a step length parameter $\tau>0$, and for all $k\in\mathbb{N}$, construct $\widetilde{\nabla F}(x^{k})$ obeying Section˜2.1 for an $\Omega\subset X$ and the distances [^3] The factors in these distances, which arise from the convenience of the operator-relative approach of Theorem 6.3, can always be compensated for by the factors $\pi_{u},\pi_{w},\alpha_{u}$, and $\alpha_{w}$ of Section 2.1.
${d_{X}(x,\tilde{x})}=\tau^{-1}\|x-\tilde{x}\|_{X},\enskip{d_{X^{*}}(x^{*},\tilde{x}^{*})}=\tau\|x^{*}-\tilde{x}^{*}\|_{X},\enskip{b_{X}(x^{k+1},x^{k})}=L\|x^{k+1}-x^{k}\|_{X}.$
Update
(53) $x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\widetilde{\nabla F}(x^{k})).$
Let $\varsigma_{p}$, $\kappa$ and $e_{p,k}$ be as in Theorem˜2.5. Then:
  1. (i)
    Section˜4.2 holds provided $\inf[F+G]>$-$\infty$, $0\leq L\tau+2\varsigma_{p}\sqrt{L\tau}<2$, and
[…]
  2. (ii)
    Section˜4.2 holds.
Suppose further that $G$ is $\gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}(\bar{x},\tau\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\gamma_{F},\gamma_{G}\in\mathbb{R}$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If
$x^{0}\in\mathbb{O}\left(\bar{x},\tau\sqrt{\delta^{2}-2r_{p}}\right)\quad\text{with}\quad r_{p}:=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}e_{p,k}/(2\tilde{\gamma})<\delta^{2}/2,$
then:
  1. 3.
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{F}:=\gamma_{F}-(\zeta/2)L$ and some $\zeta,\beta>0$,
    $(p-1)/2$ $\leq\gamma:=\tau[\gamma_{G}+\tilde{\gamma}_{F}-\beta|\tilde{\gamma}_{F}|]-\tilde{\gamma}\quad\text{and}$
    $0$ $\leq\tau[(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})]<1$
  2. 4.
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$,
    $p-1$ $\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|]-\tilde{\gamma}\quad\text{and}$
    $0$ $\leq\tau[(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]<1.$
Proof 6.8.
We apply Theorem˜6.3, whose conditions we need to verify. In its operator-relative framework, we take $\Xi=0$, $M=\tau^{-1}\operatorname{Id}$, and $\Lambda=L\operatorname{Id}$. Then the implicit step (49) holds by (53), and the condition $\Lambda\leq cM$ for a $c>0$ of Theorem˜6.3Item˜(ii) reduces to $\tau L\leq c$, which automatically holds. Further, when
  1. 1.
    in Theorem˜6.3Item˜(i), for (the minimising choice) $\tilde{\gamma}:=(\varsigma_{p}^{2}L\tau)^{-1/2}$ and $0\leq\breve{\lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\tilde{\gamma}\tau^{-1}<\tau^{-1}$, we take $\eta=1-\tau\breve{\lambda}/2>0$;
  2. 2.
    in Theorem˜6.3Item˜3 for $0\leq\breve{\lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\beta^{-1}|\gamma_{F}|-\gamma_{F}<\tau^{-1}$ we take $\eta=1-\tau\breve{\lambda}>0$;
  3. 3.
    in Theorem˜6.3Item˜4 for $0\leq\breve{\lambda}:=(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})<\tau^{-1}$ we take $\eta=1-\tau\breve{\lambda}>0$;
and in each case we take $\breve{\Lambda}=\breve{\lambda}\operatorname{Id}$ and $\tilde{\eta}=0$, then the conditions of Theorem˜6.3Items˜(i), 4 and 3 readily translate to the present respective conditions.
Remark 6.9.
According to the corollary, we can expect the subdifferential convergence Section˜4.2 to hold for small enough $\tau>0$, but stronger forms of convergence via Section˜4.2 will require the differential estimate to be good enough that $\varsigma_{p}^{2}L$ is small with respect to the available strong monotonicity. Studying (13), the main constants that control $\varsigma_{p}$ are $\pi_{u}$ and $\pi_{w}$. They can always be reduced if we repeatedly apply the inner and adjoint solvers that satisfy Section˜2.1Items˜(i) and (ii).
For exact forward-backward splitting with $\widetilde{F^{\prime}}(x^{k})=F^{\prime}(x^{k})$, by taking $p=1$, $\varsigma_{p}=0$, and $e_{p,k}=0$, and then letting γ~
0
[…]
in the previous result, we immediately obtain the following corollary. In Item˜(i), we even take $\Omega=X$, since the tracking inequalities now trivially work with that choice, and we do not assume any local properties form $F$ and $G$; in Items˜3 and 4 we do.
Corollary 6.10 (Exact outer forward-backward splitting on a Hilbert space).
On a Hilbert space $X$, suppose $F:X\to\mathbb{R}$ has an $L$-Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Pick a step length parameter $\tau>0$. Update
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\nabla F(x^{k})).$
Then:
  1. (i)
    Section˜4.2 holds provided $\inf[F+G]>$-$\infty$, and $0\leq L\tau<2.$
  2. (ii)
    Section˜4.2 holds.
Suppose further that $G$ is $\gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}(\bar{x},\tau\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\gamma_{F},\gamma_{G}\in\mathbb{R}$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If $x^{0}\in\mathbb{O}(\bar{x},\tau\delta)$, then:
  1. 3.
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{F}:=\gamma_{F}-(\zeta/2)L$ and some $\zeta,\beta>0$, we have $0\leq\gamma_{G}+\tilde{\gamma}_{F}-\beta|\tilde{\gamma}_{F}|$ and $0\leq\tau[\zeta^{-1}L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})]<1.$
  2. 4.
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$, we have $0\leq\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|$ and $0\leq\tau[L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]<1.$
Now that we have provided step length and growth conditions that prove LABEL:, 4.2, 4.2 and 4.2 for both exact and single-loop forward-backward splitting for bilevel problems, we can use Theorems˜4.6, 4.13 and 4.16 to prove convergence.
6.3 Primal-dual proximal splitting
We consider now the problem (23), i.e.,
(54) $\min_{z\in Z}\penalty 10000\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\penalty 10000\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y),$
where $Z$ and $Y$ are normed spaces equipped with self-adjoint and positive semi-definite $M_{z}\in\mathbb{L}(Z;Z^{*})$ and $M_{y}\in\mathbb{L}(Z;Z^{*})$ that work to generate the fundamental (semi-)norms in these spaces, respecting the Pythagoras’ identity (36). In Hilbert spaces, we would simply take $M_{z}=\operatorname{Id}$ and $M_{y}=\operatorname{Id}$. The functions $g:Z\to\overline{\mathbb{R}}$ and $h:Y^{*}\to\overline{\mathbb{R}}$ are convex, proper, and lower semicontinuous, and $f:Z\to\mathbb{R}$ possibly non-convex but Fréchet differentiable with $LM_{z}$-firmly Lipschitz Fréchet derivative for an $L\geq 0$. Moreover, $K\in\mathbb{L}(Z;Y^{*})$ and $h=(h_{*})^{*}$ for some $h_{*}:Y\to\overline{\mathbb{R}}$.
We will translate Theorem˜6.3 to the outer PDPS of Example˜4.1 when we estimate $f^{\prime}(z^{k})$ through the tracking theory of Section˜2. In Hilbert spaces, we would thus be interested in the method (24), i.e.,
$\left\{\begin{array}[]{l}z^{k+1}:=\operatorname{prox}_{\tau g}(z^{k}-\tau\widetilde{\nabla f}(z^{k})-\tau K^{*}y^{k}),\\ y^{k+1}:=\operatorname{prox}_{\sigma h_{*}}(y^{k}+\sigma K(2z^{k+1}-z^{k})).\end{array}\right$.
In normed spaces, we only represent the problem and method in the implicit forms (23) and (25) with
(55) $F(z,y):=f(z),\quad G(z,y):=g(z)+h_{*}(y),\quad\text{and}\quad\Xi:=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$
while the inexact PDPS becomes an instance of (49) with
(56) $\widetilde{F^{\prime}}(z^{k},y^{k}):=\begin{pmatrix}\widetilde{f^{\prime}}(z^{k})\\ 0\end{pmatrix}\quad\text{and}\quad M:=\begin{pmatrix}\tau^{-1}M_{z},&-K^{*}\\ -K&\sigma^{-1}M_{y}\end{pmatrix}$
We start by extending the standard step length assumptions of the PDPS into normed spaces. In the next lemma, in the Hilbert space setting with $M_{y}=\operatorname{Id}$ and $M_{z}=\operatorname{Id}$, we can take $K_{z}=K$ and $K_{y}=\operatorname{Id}$.
Assumption 6.10 (PDPS step length condition).
Suppose $K=K_{y}K_{z}$ for some $K_{z}\in\mathbb{L}(Z;V)$, $K_{y}\in\mathbb{L}(V;Y^{*})$, and a normed space $V$. Given $\lambda\geq 0$, the step length parameters $\tau,\sigma>0$ satisfy
$K_{y}K_{y}^{*}\leq M_{y}\quad\text{and}\quad\tau\lambda M_{z}+\tau\sigma K_{z}^{*}K_{z}\leq M_{z}.$
Lemma 6.11 (PDPS preconditioning operator).
If Section˜6.3 holds, then $M$ is positive semi-definite and for any $\gamma_{z},\gamma_{y}\geq 0$ and $\gamma:=\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}/2$, we have
$\lambda\operatorname{diag}(M_{z},\;0)\leq M\quad\text{and}\quad\gamma M\leq\operatorname{diag}(\gamma_{z}M_{z},\;\gamma_{y}M_{y}).$
Proof 6.12.
By a simple application of Young’s inequality and Section˜6.3, we have
$\|(z,y)\|_{M}^{2}=\tau^{-1}\|z\|_{M_{z}}+\sigma^{-1}\|y\|_{M_{y}}-2\langle K_{z}z|K_{y}^{*}y\rangle_{Y^{*},Y}\geq\|z\|^{2}_{\tau^{-1}M_{z}-\sigma K_{z}^{*}K_{z}}\geq\lambda\|z\|^{2}_{M_{z}}$
for any $x=(z,y)\in Z\times Y$. This establishes the first claimed inequality. The second follows by using Young’s inequality and Section˜6.3 to establish
$\gamma\|(z,y)\|_{M}^{2}\leq\gamma\|z\|^{2}_{\tau^{-1}M_{z}+\sigma K_{z}^{*}K_{z}}+\gamma\|y\|^{2}_{\sigma^{-1}M_{y}+\sigma^{-1}K_{y}^{*}K_{y}}\leq\frac{2\gamma}{\tau}\|z\|^{2}_{M_{z}}+\frac{2\gamma}{\sigma}\|y\|^{2}_{M_{y}}.$
We can now translate Theorem˜6.3 to the outer PDPS of Example˜4.1. It is missing the verification of Section˜4.2Items˜(iii) and (iv) for subdifferential convergence. Because $\Xi$ is not cyclically monotone (see [27, Chapter 24]), we see no way in general for the PDPS to satisfy that property. [^4] However, we could try to enforce the conditions, monitoring for convergence failure by setting expected bounds on $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})=[F+G](x^{N})-[F+G](x^{0})-\sum_{k=0}^{N-1}\langle\Xi x^{k+1}|x^{k}\rangle.$ In fact, if $\inf F+G>$-$\infty$, we only need to ensure that the latter sum term sum stays within chosen bounds, without having to calculate potentially costly function values.
Theorem 6.13 (PDPS with inexact $\widetilde{f^{\prime}}$; everything else exact).
Assume the setup of Eqs.˜54 and 6.3 for some $\tau,\sigma,\lambda>0$. Suppose that Section˜2.1 holds for $f$ in $\Omega\subset Z$ with [^5] We again recall that $d_{Z^{*}}(z^{*},\tilde{z}^{*})=\|z^{*}-\tilde{z}^{*}\|_{M^{-1}_{z}}$ if $M_{z}$ is invertible, in particular with $M_{z}=\operatorname{Id}$ for $Z$ a Hilbert space.
$d_{Z}(z,\tilde{z})=\|z-\tilde{z}\|_{\lambda M_{z}},\ d_{Z^{*}}(z^{*},\tilde{z}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{\lambda M_{z}}^{2})^{*}(z^{*}-\tilde{z}^{*})]^{1/2},\ \text{and}\ {b_{Z}(z^{k+1},z^{k})}=\|z^{k+1}-z^{k}\|_{LM_{z}}.$
Then
  1. (i)
    Section˜4.2 holds.
[…]
then
  1. 2.
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{f}:=\gamma_{f}-(\zeta/2)L$ and some $\beta,\zeta>0$ and $\eta\geq 0$,
    (58a) $(p-1)/2\leq\gamma$ $:=\min\{(\gamma_{g}+\tilde{\gamma}_{f}-\beta|\tilde{\gamma}_{f}|)\tau,\gamma_{h_{*}}\sigma\}/2-\tilde{\gamma}\quad\text{and}$
    (58b) $0\leq\breve{\lambda}$ $:=\zeta^{-1}L+2(\beta^{-1}|\tilde{\gamma}_{f}|-\tilde{\gamma}_{f})+2\varsigma_{p}^{2}\tilde{\gamma}^{-1}L\leq(1-\eta)\lambda.$
  2. 3.
    Section˜4.2 option Item˜(b) holds if $\Omega_{\bar{z}}$ is convex, and, for some $\beta>0$ and $\eta\geq 0$,
    (59a) $p-1\leq\gamma$ $:=\min\{(\gamma_{g}+\gamma_{f}-\beta|\gamma_{f}|)\tau,\gamma_{h_{*}}\sigma\}/2-\tilde{\gamma}\quad\text{and}$
    (59b) $0\leq\breve{\lambda}$ $:=L+\beta^{-1}|\gamma_{f}|-\gamma_{f}+\varsigma_{p}^{2}\tilde{\gamma}^{-1}L\leq(1-\eta)\lambda.$
Proof 6.14.
Recall the definitions Eqs.˜55 and 56. Observe that $F^{\prime}$ is $\Lambda$-firmly Lipschitz and $\Gamma_{F}$-monotone, and $G$ is $\Gamma_{G}$-strongly convex for
$\Lambda:=\operatorname{diag}(LM_{z},\;0),\quad\Gamma_{F}:=\operatorname{diag}(\gamma_{F}M_{z},\;0),\quad\text{and}\quad\Gamma_{G}:=\operatorname{diag}(\gamma_{G}M_{z},\;\gamma_{h_{*}}M_{y}).$
To use Theorem˜6.3, we directly verify Theorems˜2.5 and 2.8 for $x^{*}=F^{\prime}(x^{k+1})$, instead of proving Section˜2.1 for the extended functions. By Theorem˜2.5 applied to $f$ and $\widetilde{f^{\prime}}$ and Lemma˜6.11, if $\{x^{n}=(z^{n},y^{n})\}_{n=0}^{k}\subset\Omega_{\bar{x}}$, we have for $x^{k+1}=(z^{k+1},y^{k+1})$ and any $\bar{x}=(\bar{z},\bar{y})$ that
$\begin{split}\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}&=\langle\widetilde{f^{\prime}}(z^{k})-f^{\prime}(z^{k})|z^{k+1}-\bar{z}\rangle_{Z^{*},Z}\\ &\geq-\frac{\tilde{\gamma}}{2}\|z^{k+1}-\bar{z}\|_{\lambda M_{z}}^{2}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}\|z^{k+1}-z^{k}\|_{LM_{z}}-\frac{1}{2\tilde{\gamma}}e_{p,k}\\ &\geq-\frac{\tilde{\gamma}}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}.\end{split}$
Recalling the choice of distances (48), this verifies Theorem˜2.5 for $F$ and $\widetilde{F^{\prime}}$ in $\Omega_{\bar{x}}$.
[…]
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}((z^{*},0))=\sup_{(z,y)}\langle z^{*}|z\rangle-\frac{1}{2}\|(z,y)\|_{M}^{2}\leq\sup_{z}\langle z^{*}|z\rangle-\frac{\lambda}{2}\|z\|_{M_{z}}^{2}=(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{\lambda M_{z}}^{2})^{*}(z^{*}).$
If $\{x^{n}=(z^{n},y^{n})\}_{n=0}^{k}\subset\Omega_{\bar{x}}$, Theorem˜2.8 applied to $f$ and $\widetilde{f^{\prime}}$ now establishes
${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq d_{Z^{*}}^{2}(\widetilde{f^{\prime}}(z^{k}),f^{\prime}(z^{k}))\leq e_{\mathrm{lip},k},$
where $e_{\mathrm{lip},k}$ satisfies (15). The “whenever” of (15) is equivalent to $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}<\infty$. This verifies Theorem˜2.8 for $F$ and $\widetilde{F^{\prime}}$.
We proceed with proving our specific claims.
Item˜(i): Lemma˜6.11 proves $\Lambda\leq(L/\lambda)M$. Clearly $\Lambda\geq 0$. Now we use Theorem˜6.3Item˜(ii).
Item˜3: Observe that $|\Gamma_{F}|=\operatorname{diag}(|\gamma_{f}|M_{z},\;0)$. Taking $\gamma_{z}:=\gamma_{g}+\gamma_{f}-\beta|\gamma_{f}|$ and $\gamma_{y}:=\gamma_{h_{*}}$ in Lemma˜6.11, and using Eq.˜59a and (59b), we obtain the required bounds
(60) $(1-\eta)M$ $\geq(1-\eta)\lambda\operatorname{diag}(M_{z},\;0)\geq\operatorname{diag}(\breve{\lambda}M_{z},0)$
$=\breve{\Lambda}:=\Lambda+\beta^{-1}|\Gamma_{F}|-\Gamma_{F}+\varsigma_{p}^{2}\tilde{\gamma}^{-1}\Lambda\quad\text{and}$
$\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$ $=\operatorname{diag}(\gamma_{z}M_{z},\;\gamma_{y}M_{y})\geq(1/2)\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}M\geq(\gamma+\tilde{\gamma})M.$
Taking $\delta:=\lambda\delta_{z}$, (57) implies, as required, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $2r_{p}<\delta^{2}$. By (60), we have $\mathbb{O}_{M}(\bar{x},\delta)\subset\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\times\operatorname{dom}h_{*}\subset\Omega_{\bar{z}}\times\operatorname{dom}h_{*}=\Omega_{\bar{x}}$. By construction and assumption, we have $\breve{\Lambda}\geq 0$. The claim now follows from Theorem˜6.3Item˜4.
Item˜2: completely analogous to Item˜3, observing that $\tilde{\Gamma}_{F}=\operatorname{diag}(\tilde{\gamma}_{f}M_{z},0)$.
Remark 6.15 (Choice of the parameters).
If $f^{\prime}$ is locally $\gamma_{f}$-strongly monotone, the conditions Eqs.˜58 and 59 can be ensured by taking the free parameters $\tilde{\gamma},\zeta,\beta>0$ sufficiently small (to ensure Eqs.˜58a and 59a), and then $\lambda>0$ sufficiently large (to ensure Eqs.˜58b and 59b). Through the basic step length conditions Section˜6.3, a large $\lambda$ will force the primal step length $\tau>0$ to small, and allow the dual step length $\sigma>0$ to be correspondingly large. The initial iterate should also satisfy the locality bound (57), where both $\lambda$ and $\tau$ appear, the latter through $M$ defined in (25b). For Section˜4.2 only the basic step length conditions are required besides the tracking estimates on $f$.
Now that we have provided step length and growth conditions that prove LABEL:, 4.2, 4.2 and 4.2 for single-loop PDPS for bilevel problems, we can use Theorems˜4.6, 4.13 and 4.16 to prove different forms of convergence. In fact, further specialising Theorem˜4.13 to the PDPS, besides inexactness, as a novelty compared to [5, 6, 20, 12], subject to $h_{*}$ having a bounded domain, we get an estimate on the convex envelope of the objective, i.e., the Fenchel biconjugate. In non-reflexive spaces, we define the latter as a function in $X$ instead of $X^{**}$ by taking first the conjugate and then the equivalently defined preconjugate: $h^{**}:=(h_{*})_{*}$.
Corollary 6.16.
Assume the setup of Eqs.˜54 and 6.3 for some $\tau,\sigma,\lambda>0$, as well as that Theorem˜6.13Items˜2 and 3 hold for $p=1$ at some $\bar{z}\in Z$ with $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Also suppose that that $\operatorname{dom}h_{*}$ is bounded, $0\in H(\bar{x})$ for some $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$, and that the condition (57) on the initial iterate holds. Then, for the ergodic iterates $\tilde{z}^{N}:=\frac{1}{N}\sum_{k=0}^{N-1}z^{k}$, for all $N\in\mathbb{N}$, we have
$f+g+h\circ K]^{**}(\tilde{z}^{N})\leq[f+g+h\circ K](\bar{z})+\sup_{\hat{y}\in\operatorname{dom}h_{*}}\frac{1}{2N}\|(z^{0},y^{0})-(\bar{z},\hat{y})\|_{M}^{2}+\frac{\sum_{k=0}^{N-1}e_{1,k}}{2\tilde{\gamma}N}.$
Here $f+g+h\circ K](\bar{z})=[f+g+h\circ K]^{**}(\bar{z})$ if $\bar{z}$ is a global minimiser of $f+g+h\circ K$.
Proof 6.17.
[…]
(61) $\sup_{\hat{x}\in\hat{X}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\leq\sup_{\hat{x}\in\hat{X}}\left(\frac{1}{2}\|x^{0}-\hat{x}\|_{M}^{2}+\sum_{k=0}^{N-1}\frac{e_{1,k}}{2\tilde{\gamma}}\right)\quad\text{for all}\quad N\in\mathbb{N}.$
Let $\hat{x}=(\bar{z},\hat{y})\in\hat{X}$. With the expression of Example˜4.3 for the gap, we expand and estimate using the definition of the Fenchel (bi)conjugate and $h^{**}=h$ as well as $f+g]^{**}\leq f+g$ that
[…]
$\geq\bigl([f+g]^{**}(z^{k+1})+\langle Kz^{k+1}|\hat{y}\rangle-h_{*}(\hat{y})\bigr)-[f+g+h\circ K](\bar{z}).$
Summing over $k\in\{0,\ldots,N-1\}$, taking the supremum over $\hat{y}\in\operatorname{dom}h_{*}$, and using Jensen’s inequality, therefore
$\sup_{\hat{y}\in\operatorname{dom}h_{*}}\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};\hat{x})\geq N[(f+g)^{**}+h\circ K](\tilde{z}^{N})-N[f+g+h\circ K](\bar{z}).$
[…]
$f+g+h\circ K\geq[f+g+h\circ K]^{**}=((f+g)^{*}\mathop{\Box}[h\circ K]^{*})^{*}=(f+g)^{**}+h\circ K.$
Moreover, the inequality is an equality at a global minimiser (or if $f$ is convex). Now the claim follows from (61).
Remark 6.18.
Taking $p>1$ in the proof of Corollary˜6.16, linear convergence rates could be obtained as in Corollary˜4.12 for the iterates.
6.4 Further topics
We finally consider adjoint mismatch as in [19], keeping everything else exact.
Theorem 6.19 (PDPS with adjoint mismatch).
Assume the setup of Example˜4.1 with $\tau\sigma\|K\|^{2}\leq 1$ and, for simplicity, $f=0$ and Hilbert $Z$ and $Y$. Suppose $\operatorname{dom}h_{*}$ is bounded, and that $g$ and $h_{*}$ are, respectively, $\gamma_{g}$- and $\gamma_{h_{*}}$-strongly convex for some $\gamma_{g}>0$ and $\gamma_{h_{*}}\geq 0$. Let $\gamma:=\min\{\gamma_{g}\tau/4,\gamma_{h_{*}}\sigma/2\}$. In the PDPS (24), not able to compute $K^{*}$, replace it with a “mismatched” adjoint $K^{*\approx}$. Then, for any $\bar{x}\in Z\times Y$ and $p\in(1,1+2\gamma]$, Section˜4.2Item˜(a) holds with $\breve{\Lambda}=0$, $\Omega_{\bar{x}}=Z\times Y$, $\delta=\infty$, $r_{p}\leq\varepsilon/(1-p)$, and
$\varepsilon_{k}(\bar{x})=\frac{1}{2\gamma_{g}}\|(K^{*\approx}-K^{*})y^{k}\|_{Z}^{2}\leq\varepsilon:=\frac{1}{2\gamma_{g}}(\|K^{*\approx}-K^{*}\|\operatorname{diam}\operatorname{dom}h_{*})^{2}.$
Proof 6.20.
With $M$, $G$, and $F$ given by Example˜4.1, the abstract algorithm (22) reads
$M(x^{k+1}-x^{k})=:\tilde{\partial}_{k+1}=x_{k+1}^{*}+((K^{*\approx}-K^{*})y^{k},0)\quad\text{for a}\quad x_{k+1}^{*}\in H(x^{k+1}).$
Here $H$ is defined in (21). Using Lemma˜6.11 in the final step, we estimate
$\langle\tilde{\partial}_{k+1}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$ $=\langle\tilde{\partial}_{k+1}-x_{k+1}^{*}|x^{k+1}-\bar{x}\rangle_{X^{*},X}+\langle x_{k+1}^{*}-H(\hat{x})|x^{k+1}-\bar{x}\rangle_{X^{*},X}$
$\geq\langle(K^{*\approx}-K^{*})y^{k},z^{k+1}-\bar{z}\rangle+\gamma_{g}\|z^{k+1}-\bar{z}\|_{Z}^{2}+\gamma_{h_{*}}\|y^{k+1}-\bar{y}\|_{Y}^{2}$
$\geq\frac{\gamma_{g}}{2}\|z^{k+1}-\bar{z}\|_{Z}^{2}+\gamma_{h_{*}}\|y^{k+1}-\bar{y}\|_{Y}^{2}-\frac{1}{2\gamma_{g}}\|(K^{*\approx}-K^{*})y^{k}\|_{Z}^{2}$
$\geq\gamma\|x^{k+1}-\bar{x}\|_{M}^{2}-\varepsilon_{k}(\bar{x}).$
Therefore, (27) holds with the stated choices. Moreover, we have $\sum_{k=0}^{N-1}p^{k-N}\leq 1/(p-1)$ for any $p\in(1,1+2\gamma]$, verifying (29) and consequently Section˜4.2Item˜(a).
Remark 6.21 (Stochastic optimisation methods).
Also stochastic optimisation methods can be approached in our framework through lifting: we take $X$ as a space of random variables on a space $\mathscr{X}$, set $F(x)=\mathbb{E}[\mathscr{F}\circ x]$ for a function $\mathscr{F}$ on $\mathscr{X}$, $Mx:=\mathbb{E}^{*}\mathscr{M}\mathbb{E}[x]$ for $\mathscr{M}\in\mathbb{L}(\mathscr{X};\mathscr{X}^{*})$, etc., where $\mathbb{E}$ is the expectation. Without inexactness, this produces parallel copies of the very same optimisation method for every random event $\omega$. We then model the random choices made on every step through inexactness, subject to Sections˜4.2, 4.2 and 4.2.
Appendix A Scalar tracking results
We prove a a simple scalar tracking result, which will be used to establish the results of Section˜2. The following is a scalar version of Section˜2.1.
Assumption A.0.
For a given $k\geq 1$, for all $j=1,\ldots,k$, the scalars ${d^{u}_{j}},{d^{w}_{j}},{\varrho_{j}}\geq 0$ and $\widetilde{d}_{j}$$\in\mathbb{R}$ satisfy for some $\pi_{u},\pi_{w},\mu_{u}>0$, and $\kappa_{u},\kappa_{w}>1$, as well as $\alpha_{u},\alpha_{w}>0$, the inequalities
(i) $\kappa_{u}{d^{u}_{j+1}}$ $\leq{d^{u}_{j}}+\pi_{u}{\varrho_{j}},$
(ii) $\kappa_{w}{d^{w}_{j+1}}$ $\leq{d^{w}_{j}}+{\mu_{u}}{d^{u}_{k+1}}+\pi_{w}{\varrho_{j}},\quad\text{and}$
(iii) ${\widetilde{d}_{j}}$ $\leq\alpha_{u}{d^{u}_{j+1}}+\alpha_{w}{d^{w}_{j+1}}.$
We wish to develop simple bounds for $\widetilde{d}_{k}$} that can be readily used in convergence proofs when Appendix˜A is instantiated as Section˜2.1. These core estimates then allow us to isolate the contributions of initialisation and update errors, and thereby quantify the impact of inexact inner and adjoint solutions over multiple iterations on the differential approximations. We start with a result that unrolls the recursion in Appendix˜A.
Lemma A.1.
Let Appendix˜AEqs.˜i and ii hold for a $k\geq 1$. Then, letting $\iota_{k}:=\sum_{m=1}^{k}\kappa_{u}^{-m}\kappa_{w}^{-(k+1-m)}$ (understanding that $\iota_{0}=0$), for all $k\geq 0$ and $\alpha_{u},\alpha_{w}\geq 0$, we have
(62) $R^{k+1}(\alpha_{u},\alpha_{w})=\alpha_{u}{d^{u}_{k+1}}+\alpha_{w}{d^{w}_{k+1}}$ $\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-k}{d^{w}_{1}}$
$+\sum_{j=0}^{k-1}\bigl(\alpha_{u}\kappa_{u}^{-(k-j)}\pi_{u}+\alpha_{w}[\iota_{k-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(k-j)}\pi_{w}]\bigr){\varrho_{j+1}}.$
Proof A.2.
For $k=1$, we have $\iota_{1}=\kappa^{-1}_{u}\kappa^{-1}_{w}$ and, by assumption, ${d^{u}_{2}}\leq\kappa^{-1}_{u}{d^{u}_{1}}+\kappa^{-1}_{u}\pi_{u}{\varrho_{1}}\text{ and }{d^{w}_{2}}\leq\kappa^{-1}_{w}{d^{w}_{1}}+\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{2}}+\kappa^{-1}_{w}\pi_{w}{\varrho_{1}}.$ Multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, then summing up, observing to cancel the two instances of $\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{2}}$, establishes (62).
We then take $k=n+1$, and proceed by induction, assuming (62) to hold for $k=n$. Again, ${d^{u}_{n+2}}\leq\kappa^{-1}_{u}{d^{u}_{n+1}}+\kappa^{-1}_{u}\pi_{u}{\varrho_{n+1}}\text{ and }{d^{w}_{n+2}}\leq\kappa^{-1}_{w}{d^{w}_{n+1}}+\kappa^{-1}_{w}{\mu_{u}}{d^{u}_{n+1}}+\kappa^{-1}_{w}\pi_{w}{\varrho_{n+1}}$ by assumption. As in the case $k=1$, multiplying the former by $\alpha_{u}+\alpha_{w}\kappa^{-1}_{w}{\mu_{u}}$ and the latter by $\alpha_{w}$, and then summing up, yields
$R^{n+2}(\alpha_{u},\alpha_{w})=\alpha_{u}{d^{u}_{n+2}}+\alpha_{w}\kappa^{-1}_{w}{d^{w}_{n+2}}$ $\leq(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}}){d^{u}_{n+1}}+\alpha_{w}\kappa^{-1}_{w}{d^{w}_{n+1}}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}]){\varrho_{n+1}}.$
The first two terms on the right-hand side equal $R^{n+1}(\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}},\alpha_{w}\kappa^{-1}_{w})$, so using (62) for $k=n$, we continue
$R^{n+2}(\alpha_{u},\alpha_{w})$ $\leq((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-n}+\alpha_{w}\kappa^{-1}_{w}\iota_{n}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa^{-1}_{w}\kappa_{w}^{-n}{d^{w}_{1}}$
$+\sum_{j=0}^{n-1}\bigl((\alpha_{u}\kappa^{-1}_{u}+\alpha_{w}\kappa^{-1}_{w}\kappa^{-1}_{u}{\mu_{u}})\kappa_{u}^{-(n-j)}\pi_{u}+\alpha_{w}\kappa^{-1}_{w}[\iota_{n-j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-(n-j)}\pi_{w}]\bigr){\varrho_{j+1}}$
$+(\alpha_{u}\kappa^{-1}_{u}\pi_{u}+\alpha_{w}[\kappa^{-1}_{w}\kappa^{-1}_{u}\pi_{u}{\mu_{u}}+\kappa^{-1}_{w}\pi_{w}]){\varrho_{n+1}}$
$=(\alpha_{u}\kappa_{u}^{-(n+1)}+\alpha_{w}\mu_{u}(\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n})){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-(n+1)}{d^{w}_{1}}$
$+\sum_{j=0}^{n}\bigl(\alpha_{u}\kappa_{u}^{-(n+1-j)}\pi_{u}+\alpha_{w}[(\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}){\mu_{u}}\pi_{u}+\kappa_{w}^{-(n+1-j)}\pi_{w}]\bigr){\varrho_{j+1}}.$
Here $\kappa^{-1}_{w}\kappa_{u}^{-(n+1-j)}+\kappa^{-1}_{w}\iota_{n-j}=\iota_{n+1-j}$, as by the definition of $\iota_{n+1}$, for any $n\geq 0$,
(63) $\iota_{n+1}=\sum_{m=1}^{n+1}\kappa_{u}^{-m}\kappa_{w}^{-(n+2-m)}=\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\sum_{m=1}^{n}\kappa_{u}^{-m}\kappa_{w}^{-(n+2-m)}=\kappa^{-1}_{w}\kappa_{u}^{-(n+1)}+\kappa^{-1}_{w}\iota_{n},$
Thus we obtain (62) for $k=n+1$.
The next two lemmas form our core estimates. To simplify the estimates, recalling that $\kappa_{u},\kappa_{w}>1$, we observe that
(64) $p^{k}\iota_{k}\leq p^{-1}k(\kappa/p)^{-(k+1)}\quad\text{for}\quad\kappa:=\min(\kappa_{u},\kappa_{w})>1\text{ and any }p\in(0,\kappa).$
Thus, by sum formulae for arithmetic-geometric progressions [14, formula 0.113],
(65) $\sum_{k=0}^{n-1}p^{k}\iota_{k}\leq\sum_{k=0}^{\infty}p^{k}\iota_{k}\leq p^{-1}(\kappa/p-1)^{-2}=p(\kappa-p)^{-2}\quad\text{for all }n\in\mathbb{N}.$
The next lemma bounds the distance between the true differential and estimate at iteration $k$ through an error term. In the two lemmas that follow, we further estimate the error term. We use the first lemma directly in Theorem˜2.8.
Lemma A.3.
Suppose Appendix˜A holds for a $k\geq 1$. Then for any $p\in(0,\kappa)$, we have
(66) ${\widetilde{d}_{k}^{2}}\leq\breve{e}_{p,k},$
where, for $\psi_{j}:=\alpha_{u}\kappa_{u}^{-j}\pi_{u}+\alpha_{w}[\iota_{j}{\mu_{u}}\pi_{u}+\kappa_{w}^{-j}\pi_{w}]$ and $\overline{\kappa}:=\max\{\kappa_{u},\kappa_{w}\}$, we set
(67) $\varsigma_{p}$ $:=\frac{\overline{\kappa}}{p}\sum_{j=0}^{\infty}p^{j}\psi_{j}\leq\frac{(\alpha_{u}\pi_{u}+\alpha_{w}\pi_{w})\kappa\overline{\kappa}}{p(\kappa-p)}+\frac{\alpha_{w}{\mu_{u}}\pi_{u}\overline{\kappa}}{p^{2}(\kappa-p)^{2}}\quad\text{and}$
(68) $\breve{e}_{p,k}$ $:=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}p^{k}}{(d^{u}_{1})^{2}}+\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}p^{k}}{(d^{w}_{1})^{2}}+\sum_{j=0}^{k-1}\frac{\varsigma_{p}\psi_{k-j}}{p^{k-j}}{\varrho_{j+1}}.$
Proof A.4.
We prove for all $s\in\mathbb{R}$ that
(69) $4s{\widetilde{d}_{k}}-4s^{2}\leq\breve{e}_{p,k}.$
Maximising the left hand side with respect to $s$ then establishes (66).
By Appendix˜AEq.˜iii, we have
${\widetilde{d}_{k}}\leq\alpha_{u}{d^{u}_{k+1}}+\alpha_{w}{d^{w}_{k+1}}=:R^{k+1}.$
Since $\{x^{n}\}_{n=0}^{k}\subset\Omega$, invoking the inner and adjoint tracking Appendix˜AEqs.˜i and ii and Lemma˜A.1, we obtain
$R^{k+1}\leq(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}}){d^{u}_{1}}+\alpha_{w}\kappa_{w}^{-k}{d^{w}_{1}}+\sum_{j=0}^{k-1}\psi_{k-j}{\varrho_{j+1}}.$
Using Young’s inequality several times here, we deduce for any $\theta_{k}^{u},\theta_{k}^{w},\theta_{k,j}>0$ that
(70) $\begin{split}4sR^{k+1}&\leq\frac{(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})^{2}}{\theta_{k}^{u}}{(d^{u}_{1})^{2}}+\frac{(\alpha_{w}\kappa_{w}^{-k})^{2}}{\theta_{k}^{w}}{(d^{w}_{1})^{2}}\\ &+\sum_{j=0}^{k-1}\frac{\psi_{k-j}^{2}}{\theta_{k,j}}{\varrho_{j+1}^{2}}+4\left(\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}\right)s^{2}.\end{split}$
Take $\theta_{k}^{u}=p^{k}\varsigma_{p}^{-1}\pi_{u}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})$, $\theta_{k}^{w}=p^{k}\varsigma_{p}^{-1}\pi_{w}\alpha_{w}\kappa_{w}^{-k}$, and $\theta_{k,j}=\varsigma_{p}^{-1}p^{k-j}\psi_{k-j}$. Observe from (63) that $\iota_{k}\leq\kappa_{w}\iota_{k+1}$. Hence $p^{k}\iota_{k}\leq(\kappa_{w}/p)p^{k+1}\iota_{k+1}$, and further, $p^{k}\psi_{k}\leq(\overline{\kappa}/p)p^{k+1}\psi_{k+1}$, where $\overline{\kappa}/p>1$. Now
$\theta_{k}^{u}+\theta_{k}^{w}+\sum_{j=0}^{k-1}\theta_{k,j}=\frac{1}{\varsigma_{p}}\left(p^{k}\psi_{k}+\sum_{j=1}^{k}p^{j}\psi_{j}\right)\leq\frac{\overline{\kappa}}{\varsigma_{p}p}\sum_{j=0}^{k+1}p^{j}\psi_{j}\leq 1.$
Thus rearranging (70) establishes (69). Finally, the bound in (67) on $\varsigma_{p}$ follows from Eq.˜65 and $\sum_{j=0}^{\infty}(p/\kappa)^{j}=1/(1-p/\kappa)=\kappa/(\kappa-p)$.
We use the next result in Theorem˜2.5.
Lemma A.5.
Suppose Appendix˜A holds for a $k\geq 1$. Then for any $p\in[1,\kappa)$, we have
(71) ${\widetilde{d}_{k}^{2}}\leq\varsigma_{p}^{2}{\varrho_{k+1}}+e_{p,k},$
where, for $\breve{e}_{p,k}$ defined (68),
(72) $e_{p,k}:=\breve{e}_{p,k}-\varsigma_{p}^{2}{\varrho_{k+1}}.$
satisfies
(73) $\begin{split}\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k}e_{p,k}&\leq\frac{{(d^{u}_{1})^{2}}}{\pi_{u}}\bigg(\frac{\varsigma_{p}\alpha_{u}\kappa}{\kappa-1}+\frac{\varsigma_{p}\alpha_{w}{\mu_{u}}}{(\kappa-1)^{2}}\bigg)+\frac{{(d^{w}_{1})^{2}}}{\pi_{w}}\bigg(\frac{\varsigma_{p}\alpha_{w}\kappa}{\kappa-1}\bigg).\end{split}$
Proof A.6.
We only need to prove (73); the rest is immediate from Lemma˜A.3. Let
$A_{k}$ $=\frac{\varsigma_{p}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}{\mu_{u}})}{\pi_{u}}{(d^{u}_{1})^{2}},$ $B_{k}$ $=\frac{\varsigma_{p}\alpha_{w}\kappa_{w}^{-k}}{\pi_{w}}{(d^{w}_{1})^{2}},$
$C_{k}$ $=\sum_{j=0}^{k-1}p^{j}\varsigma_{p}\psi_{k-j}{\varrho_{j+1}},\quad\text{and}$ $D_{k}$ $=p^{j}\varsigma_{p}^{2}{\varrho_{k+1}}.$
Then $p^{k}e_{p,k}=:A_{k}+B_{k}+C_{k}-D_{k}$ from Eqs.˜72 and 68. By the assumption $p\in[1,\kappa)$, we have $\frac{\overline{\kappa}}{p}p^{j}\geq 1$. Hence from (67) we have that $\varsigma_{p}=\frac{\overline{\kappa}}{p}\sum_{j=0}^{\infty}p^{j}\psi_{j}\geq\sum_{j=0}^{\infty}\psi_{j}.$ Now
$\begin{split}\sum_{k=0}^{N-1}C_{k}&=\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}p^{j}\varsigma_{p}\psi_{k-j}{\varrho_{j+1}^{2}}=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{k=j+1}^{N-1}\psi_{k-j}{\varrho_{j+1}^{2}}\\ &=\varsigma_{p}\sum_{j=0}^{N-2}p^{j}\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}{\varrho_{j+1}^{2}}\leq\sum_{j=0}^{N-2}p^{j}\varsigma_{p}^{2}{\varrho_{j+1}^{2}}\leq\sum_{k=0}^{N-1}D_{k}.\end{split}$
Moreover, using (65) and the sum formula for geometric series, we estimate that $\sum_{k=0}^{N-1}(A_{k}+B_{k})$ is less than the right-hand side of (73).
We use the next lemma in Theorem˜2.8.
Lemma A.7.
Suppose Appendix˜A holds for all $k\geq 1$. Then, for $\breve{e}_{1,k}$ given in (67) we have
$\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\breve{e}_{1,k}<\infty\quad\text{whenever}\quad\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{\varrho_{j+1}^{2}}<\infty.$
Proof A.8.
Observe that
(74) $\breve{e}_{1,k}=\frac{\varsigma_{1}}{\pi_{u}}(\alpha_{u}\kappa_{u}^{-k}+\alpha_{w}\iota_{k}\mu_{u}){(d^{u}_{1})^{2}}+\frac{\varsigma_{1}}{\pi_{w}}\alpha_{w}\kappa_{w}^{-k}{(d^{w}_{1})^{2}}+\sum_{j=0}^{k-1}\varsigma_{1}\psi_{k-j}{\varrho_{j+1}^{2}}.$
Using (65) and the sum formula for geometric series, we bound the sum of the first two terms over $k\in\{0,\ldots,N-1\}$ by
${(d^{u}_{1})^{2}}\frac{\varsigma_{1}}{\pi_{u}}\bigg(\frac{\alpha_{u}\kappa}{\kappa-1}+\frac{\alpha_{w}\mu_{u}}{(\kappa-1)^{2}}\bigg)+{(d^{w}_{1})^{2}}\frac{\varsigma_{1}}{\pi_{w}}\bigg(\frac{\alpha_{w}\kappa}{\kappa-1}\bigg).$
Lemma˜A.3 establishes $\varsigma_{1}=\overline{\kappa}\sum_{j=0}^{\infty}\psi_{j}<\infty.$ With this, we sum the third term of (74) over $k\in\{0,\ldots,N-1\}$, and change the order of summation to obtain
$\varsigma_{1}\sum_{k=0}^{N-1}\sum_{j=0}^{k-1}\psi_{k-j}{\varrho_{j+1}^{2}}=\varsigma_{1}\sum_{j=0}^{N-2}\left(\sum_{\ell=0}^{N-2-j}\psi_{\ell+1}\right){\varrho_{j+1}^{2}}\leq\frac{\varsigma_{1}^{2}}{\overline{\kappa}}\sum_{j=0}^{N-2}{\varrho_{j+1}^{2}}.$
Thus $\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}\breve{e}_{1,k}<\infty$ whenever $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}{\varrho_{j+1}^{2}}<\infty$.
Appendix B Opial’s lemma for quasi-Féjer monotonicity
Here we prove a generalisation of Opial’s lemma [25] for quasi-Féjer monotonicity, i.e, Féjer monotonicity with an additive error term. We prove it in normed spaces for Bregman divergences (34), as they add no extra difficulties. In an even more general variable-metric framework, a similar result is also proved in [23, Proposition 2.7]. Our simplified proof follows the outline of that in [7], and is nearly identical to the one in [35], where the errors took a more specific form.
For the proof, we recall the following deterministic version of the results of [26]:
Lemma B.1.
Let $\{a_{k}\}_{k\in\mathbb{N}}$, $\{b_{k}\}_{k\in\mathbb{N}}$, $\{c_{k}\}_{k\in\mathbb{N}}$, and $\{d_{k}\}_{k\in\mathbb{N}}$ be non-negative and $a_{k+1}\leq a_{k}(1+b_{k})+c_{k}-d_{k}$ for all $k\in\mathbb{N}$. If $\sum_{k=0}^{\infty}b_{k}<\infty$ and $\sum_{k=0}^{\infty}c_{k}<\infty$, then (i) $\lim_{k\to\infty}a_{k}$exists and is finite; and (ii) $\sum_{k=0}^{\infty}d_{k}<\infty$.
Lemma B.2.
Let either $X$ be the dual space of a corresponding separable normed space $X_{*}$, or, alternatively, let $X$ be reflexive. Also let $M:X\to\mathbb{R}$ be convex, proper, and Gâteaux differentiable with $M^{\prime}:X\to X_{*}$ weak-*-to-weak continuous. Finally, let $\hat{X}\subset X$ be non-empty and $\{e_{k}(\bar{x})\}_{k\in\mathbb{N}}\in\mathbb{R}$ for all $\bar{x}\in\hat{X}$. If
  1. (i)
    all weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$ belong $\hat{X}$;
  2. (ii)
    $B_{M}(x^{k+1},\bar{x})\leq B_{M}(x^{k},\bar{x})+e_{k}(\bar{x})$ for some $e_{k}(\bar{x})\geq 0$ for all $\bar{x}\in\hat{X}$ and $k\in\mathbb{N}$; and
  3. (iii)
    $\sum_{k=0}^{\infty}e_{k}(\bar{x})<\infty$ for all $\bar{x}\in\hat{X}$;
then all weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$ satisfy $\hat{x},\bar{x}\in\hat{X}$ and
(75) $\langle M^{\prime}(\hat{x})-M^{\prime}(\bar{x})|\hat{x}-\bar{x}\rangle=0.$
[…]
Proof B.3.
Let $\bar{x}$ and $\hat{x}$ be weak-* limit points of $\{x^{k}\}_{k\in\mathbb{N}}$. Since Bregman divergences $B_{M}\geq 0$ for convex $M$, the conditions Items˜(iii) and (ii) establish the assumptions of Lemma˜B.1 for $a_{k}=B_{M}(x^{k};\bar{x})$, $b_{k}=0$, $c_{k}=e_{k}(\bar{x})$, and $d_{k}=0$. It follows that $\{B_{M}(x^{k};\bar{x})\}_{k\in\mathbb{N}}$ is convergent. Likewise we establish that $\{B_{M}(x^{k};\hat{x})\}_{k\in\mathbb{N}}$ is convergent. Therefore, by the obvious three-point identity for Bregman divergences (see, e.g., [33]),
$\langle M^{\prime}(x^{k})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=B_{M}(x^{k};\hat{x})-B_{M}(x^{k};\bar{x})+B_{M}(\hat{x};\bar{x})\to c\in\mathbb{R}.$
Since $\bar{x}$ and $\hat{x}$ are a weak-* limit point, there exist subsequences $\{x^{k_{n}}\}_{n\in\mathbb{N}}$ and $\{x^{k_{m}}\}_{m\in\mathbb{N}}$ with $x^{k_{n}}\mathrel{\rightharpoonup}\bar{x}$ and $x^{k_{m}}\mathrel{\rightharpoonup}\hat{x}$. By the weak-*-to-weak continuity of $M^{\prime}:X\to X_{*}$, (75) follows from
$\langle M^{\prime}(\bar{x})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=\lim_{n\to\infty}\langle M^{\prime}(x^{k_{n}})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=c=\lim_{m\to\infty}\langle M^{\prime}(x^{k_{m}})-M^{\prime}(\hat{x})|\bar{x}-\hat{x}\rangle=0.$
If $\{x^{k}\}_{k\in\mathbb{N}}$ is bounded, and $X$ is the dual space of some separable normed space $X_{*}$, it contains a weakly-* convergent subsequence by the Banach–Alaoglu theorem, so a limit point exists as claimed. If $X$ is reflexive, the Eberlein–S̆mulyan theorem establishes the same result. Hence, if (75) implies $\bar{x}=\hat{x}$, then every convergent subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ has the same weak limit. It lies in $\hat{X}$ by Item˜(i). The final claim now follows from a standard subsequence–subsequence argument: Assume to the contrary that there exists a subsequence of $\{x^{k}\}_{k\in\mathbb{N}}$ not convergent to $\hat{x}$. Then the above argument provides a further subsequence converging to $\hat{x}$. This contradicts the fact that any subsequence of a convergent sequence converges to the same limit.